### v1

# Experimental targets for dark photon dark matter

Abstract
Ultralight dark photon dark matter features distinctive cosmological and astrophysical signatures and is also supported by a burgeoning direct-detection program searching for its kinetic mixing with the ordinary photon over a wide mass range. Dark photons, however, cannot necessarily constitute the dark matter in all of this parameter space. In minimal models where the dark photon mass arises from a dark Higgs mechanism, early Universe dynamics can easily breach the regime of validity of the low-energy effective theory for a massive vector field. In the process, the dark sector can collapse into a cosmic string network, precluding dark photons as viable dark matter. We establish the general conditions under which dark photon production avoids significant backreaction on the dark Higgs and identify regions of parameter space that naturally circumvent these constraints. After surveying implications for known dark photon production mechanisms, we propose novel models that set well-motivated experimental targets across much of the accessible parameter space. We also discuss complementary cosmological and astrophysical signatures that can probe the dark sector physics responsible for dark photon production.
I Introduction
The cosmological and astrophysical evidence for cold, nonbaryonic dark matter provides some of the most compelling motivation for physics beyond the Standard Model [1, 2, 3]. These observations, however, only evince dark matter through its gravitational effects, allowing a broad range of new particles to effectively reproduce the cold dark matter paradigm in its observed regimes. From a bottom-up perspective, dark matter candidates span dozens of orders of magnitude in mass and all possible spins; the only general requirement of their nongravitational interactions with Standard Model (SM) particles is that they are sufficiently weak to have evaded detection thus far.
Confronted with so vast a landscape of possibilities, top-down theoretical approaches can provide essential guidance by motivating concrete models and identifying consistent scenarios. The canonical example, a weakly interacting massive particle (WIMP) [1], sets specific experimental targets by determining the dark matter relic abundance via its interaction strength with SM particles through thermal freeze-out [4, 5, 6]. Dark matter as a thermal relic would be too warm at late times if its mass were below $\sim\mathrm{keV}$ [7, 8, 9, 10, 11, 12, 13, 14]; in this mass range, the dark matter must also be bosonic [15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27]. Axions and dark photons dominate theoretical and experimental efforts in this ultralight regime, since symmetries can protect their masses from large quantum corrections. Notably, the quantum chromodynamics (QCD) axion, initially proposed to resolve the strong CP problem [28, 29, 30], is an excellent dark matter candidate [31, 32, 33, 34] and has inspired a broad experimental program targeting its couplings to SM particles in the KSVZ-DFSZ range [35, 36, 37, 38].
New, massive gauge bosons may not solve specific problems in the SM like the axion does, but they are well motivated (if not expected) in extensions of the SM [39, 40, 41, 42, 43, 44, 45, 46] as the force carriers of new interactions. Indeed, massive gauge bosons are a feature of the SM itself. Moreover, the dark photon inherits the axion’s broad experimental reach [47]: while detecting the axion’s mixing with the SM photon requires catalysis by a background magnetic field, the dark photon’s kinetic mixing with the SM photon takes place in vacuum. Every axion experiment is therefore also a dark photon experiment.
In this work, we investigate how theoretical considerations motivate the extensive dark photon parameter space accessible to upcoming searches. We focus on a fundamental requirement for any dark matter candidate: a consistent mechanism for generating its relic abundance. In particular, the low-energy effective theory of a massive, kinetically mixed vector appropriate to describe its phenomenology for direct detection may well be inadequate to describe its cosmological production at early times (i.e., at higher energy). In minimal models where the ultraviolet (UV) completion of the dark photon’s mass is a Higgs mechanism, the cutoff scale is directly tied to the kinetic mixing strength. Reference [48] demonstrated that breaching this cutoff has severe consequences: topological defects (Nielsen-Olesen strings [49]) form, converting any coherent dark photon background into string kinetic and potential energy. The resulting string network is neither a viable dark matter candidate nor detectable by haloscopes. In the minimal scenario of vector production from inflationary fluctuations [50], the entire experimentally accessible parameter space is excluded on these grounds [48]. Here we extend this analysis to constrain additional dark photon production mechanisms, propose extensions thereof and novel scenarios with enhanced experimental prospects—including an in-depth discussion and generalization of the models of scalar-mediated dark photon production introduced in Refs. [51, 52]—and identify theoretically motivated targets within the dark photon parameter space.
In Sec. II, we derive model-independent bounds on kinetic mixing as a function of wave number and redshift of production, applying these bounds to various models proposed in prior literature. We then consider the clockwork mechanisms as a production-agnostic means to relax these bounds in Sec. III, demonstrating that they introduce no additional fine-tuning issues. In Sec. IV, we revisit dark photon production from oscillating axions and describe mechanisms that unify enhanced production efficiency with dynamics that avoid string formation. Finally, in Sec. V we consider a class of models coupling the dark photon to a new singlet scalar, discussing the scalar’s potential to alleviate backreaction constraints by modulating the fundamental parameters of the Abelian-Higgs theory. We show that the Damour-Polyakov mechanism [53] can weaken bounds on inflationary production, allowing accessible kinetic mixing for dark photons heavier than an ${\mathrm{eV}}$. We further discuss two qualitatively distinct postinflationary production mechanisms driven by the scalar [51, 52] that, with some fine-tuning, may open all kinetic mixing parameter space accessible to upcoming haloscopes. In Sec. VI, we summarize and discuss the implications of our findings and conclude in Sec. VII. A number of appendices provide additional details on scalar–Abelian-Higgs models (Appendix A), quantum corrections to the scalar effective potential in these models (Appendix B), the strong-coupling (Appendix C) and narrow-resonance (Appendix D) regimes of dark photon production from axions, and the effect of the Standard Model plasma on the narrow scalar resonance discussed in Sec. V.4 (Appendix E).
Throughout this paper, we use natural units in which $\hbar=c=1$, define the reduced Planck mass $M_{\mathrm{pl}}=1/\sqrt{8\pi G}$, fix a cosmic-time Friedmann-Lemaître-Robertson-Walker (FLRW) metric $\mathrm{d}s^{2}=\mathrm{d}t^{2}-a(t)^{2}\delta_{ij}\mathrm{d}x^{i}\mathrm{d}x^{j}$ with $a(t)$ the scale factor, and employ the Einstein summation convention for spacetime indices. Dots denote derivatives with respect to cosmic time $t$, and the Hubble rate is $H\equiv\dot{a}/a$. We use boldface for spatial three vectors.
II Backreaction, defect formation, and a survey of production scenarios
We begin by reviewing Abelian-Higgs theory and the thresholds associated with backreaction onto the dark Higgs and defect formation [48]. We then identify generic conditions for postinflationary cosmological production of dark photons that evade these thresholds. We discuss defect bounds on inflationary production of dark photons in Sec. II.1, considering the effects of modified postinflationary expansion histories (Sec. II.1.1) and possible nonminimal couplings to gravity and the inflaton itself (Sec. II.1.2). In Sec. II.2 we derive model-independent bounds from a dark photon’s kinetic mixing with the hot Standard Model plasma, and in Sec. II.3 we discuss dark photon production via an oscillating axion. We briefly consider vector misalignment in Sec. II.4.
In this work, we consider a dark sector described by various extensions of the Abelian-Higgs model. We denote the dark photon with $A$, not to be confused with the Standard Model photon ${A_{\mathrm{SM}}}$. The dark Higgs field $\Phi$ that provides the dark photon its mass is a complex scalar charged under $A$ with gauge coupling $g_{D}$ and a vacuum expectation value (VEV) $v$. We refer to the dark Higgs simply as the Higgs, since the SM Higgs is not relevant to our discussion. The action of Abelian-Higgs model is $S=\int\mathrm{d}^{4}x\,\sqrt{-g}\mathcal{L}$ with Lagrangian
$\mathcal{L}$ $=-\frac{1}{4}F_{\mu\nu}{F}^{\mu\nu}+\frac{1}{2}D_{\mu}\Phi\left(D^{\mu}\Phi\right)^{\ast}-\frac{\lambda}{4}\left({\left|\Phi\right|}^{2}-v^{2}\right)^{2},$ (1)
where $F_{\mu\nu}=\partial_{\mu}A_{\nu}-\partial_{\nu}A_{\mu}$ is the dark photon field strength, $D_{\mu}=\partial_{\mu}-\mathrm{i}g_{D}A_{\mu}$ is the gauge-covariant derivative, and $\lambda$ is the Higgs self coupling.
In general, we seek to identify the conditions under which a cosmological abundance of dark photons in the early Universe—typically, their maximum density which they reach when first produced—invalidates the low energy effective theory given by the Proca Lagrangian,
$\mathcal{L}$ $=-\frac{1}{4}F_{\mu\nu}{F}^{\mu\nu}+\frac{1}{2}m_{A}^{2}A_{\mu}A^{\mu},$ (2)
and whatever description of dark photon production that relies upon it. The Proca Lagrangian derives from the low-energy limit of Eq. 1 when the dark photon’s U(1) symmetry is broken. In the broken phase, the dark photon and Higgs have masses $m_{A}=g_{D}v$ and $m_{h}=\sqrt{2\lambda}v$, where the Higgs is decomposed as $\Phi\equiv(v+h)e^{\mathrm{i}\Pi/v}$. In unitary gauge, which sets $\Pi=0$, the dark photon acquires a longitudinal mode as it eats the Higgs phase, i.e., $A_{\mu}-\partial_{\mu}\Pi/m_{A}\to A_{\mu}$. The theory is then described by the action
$\mathcal{L}$ $=-\frac{1}{4}F_{\mu\nu}F^{\mu\nu}+\frac{1}{2}\partial_{\mu}h\partial^{\mu}h-V_{\mathrm{eff}}(h,A),$ (3)
where the effective potential for the Higgs and dark photon is
$V_{\mathrm{eff}}(h,A)$ $=\frac{\lambda v^{4}}{4}\left[1+\left(\frac{h}{v}+1\right)^{4}\right]-\frac{1}{2}\left(\frac{h}{v}+1\right)^{2}\left(\lambda v^{4}+m_{A}^{2}A_{\mu}A^{\mu}\right).$ (4)
The first term in the effective potential is a pure quartic with a unique minimum at the symmetric point $h=-v$. The latter term generates the symmetry-breaking minimum at $h=0$ provided that the dark photon amplitude is not too large—that is, the dark photon backreacts strongly onto the Higgs if
$\left|m_{A}^{2}A_{\mu}A^{\mu}\right|$ $\gtrsim\lambda v^{4}.$ (5)
As pointed out in Ref. [54], beyond this threshold the dark photon is no longer well described by the Proca Lagrangian Eq. 2, and whatever dynamics are responsible for producing the dark photon are (presumably) substantially modified.
More recently, Ref. [48] showed that the consequences of violating Eq. 5 can be severe. The vacuum state of the Abelian-Higgs theory is characterized by a ring of degenerate vacua, allowing for topological line defects—namely, Nielsen-Olisen strings [49]. Each string has a quantized magnetic flux $\Phi_{0}=2\pi/g_{D}$; the strings therefore interact with the background dark electromagnetic field, converting the would-be dark matter into string energy.111 Though strings only have an intrinsic magnetic moment per unit length, moving strings also possess an electric dipole moment; they therefore interact with both electric and magnetic fields of the dark photon. For the purpose of understanding string formation, the dark photon, which otherwise would have constituted the cold dark matter, can often be approximated as a constant, homogeneous magnetic field (see Ref. [48] for a more comprehensive discussion). Magnetic flux in excess of one unit ($2\pi/g_{D}$) prefers to be confined to a topological defect by a phenomenon known as flux trapping [55]. That is, strings are the energetically preferable field configuration when the magnetic field exceeds the first critical field $B_{c1}\sim g_{D}v^{2}$, corresponding to a magnetic flux $2\pi/g_{D}$ inside an area $\sim 1/m_{A}^{2}$. However, there is an energy barrier to string formation (due to the differing spacetime symmetry and topology before and after string formation), which only vanishes when the magnetic field exceeds the parametrically larger “superheating” field $B_{\mathrm{sh}}=\sqrt{\lambda}v^{2}$.222 The phase transition is entirely nonthermal, making “superheating” a misnomer. In this case, the magnetic flux confines to strings with interstring spacing much larger than the string core, $1/\sqrt{g_{D}B_{\mathrm{sh}}}\gg m_{h}^{-1}$, so $\mathrm{U}(1)$ symmetry is restored locally. Only once the external magnetic field exceeds the second critical field $B_{c2}\sim(\lambda/g_{D}^{2})^{1/2}m_{h}^{2}$ do the string cores overlap, restoring symmetry globally.
The idealized limit of a purely magnetic, homogeneous background field is not precisely applicable to all scenarios for dark photon production, some of which produce semirelativistic or nonrelativistic modes (with wave numbers $k\lesssim m_{A}$) that are coherent on scales $1/m_{A}$ (or larger) but also contain electric fields. Reference [48] demonstrated with numerical simulations that defects nucleate nonetheless for tachyonic dark photon production (see Sec. II.3) at energy densities above the superheating threshold. The superheating threshold also turns out to match that for strong backreaction onto the Higgs identified in Eq. 5. We therefore take Eq. 5 as generic upper limit on the dark photon field amplitude: as Proca theory is an invalid description of the Abelian-Higgs model in this regime either way, and the dynamics intended to produce massive dark photons are in all likelihood disrupted. In the rest of this section we discuss exceptions to this argument as needed.
To study the implications of Eq. 5 as a consistency limit on dark photon cosmologies, we first note that in almost all the cases we consider, dark photons are produced with a peaked spectrum. Namely, the relic abundance is dominated by dark photons over a narrow range of wave numbers centered about the peak $k_{\star}$, which enables rephrasing the backreaction bound in terms of the dark photon energy density $\rho_{A^{\prime}}$. Heuristically (see Appendix A for more formal expressions),
$\lambda v^{4}\gg-m_{A}^{2}g^{\mu\nu}A_{\mu}A_{\nu}$ $\sim\frac{m_{A}^{2}}{m_{A}^{2}+k_{\star}^{2}/a^{2}}\rho_{A,\perp}+\rho_{A,\parallel},$ (6)
where $\rho_{A,\perp}$ and $\rho_{A,\parallel}$ denote the contributions from transverse and longitudinal modes, respectively. Equation 6 shows that for equal energy densities, relativistic, transverse dark photons couple more weakly to the Higgs than nonrelativistic ones by a factor of $(am_{A}/k_{\star})^{2}$. But the energy density in relativistic modes redshifts as $a^{-4}$, one factor of $a$ faster than nonrelativistic modes, and must therefore be produced with a larger initial energy density to match the observed dark matter density. The backreaction constraint is therefore relaxed, albeit modestly, if transverse dark photons dominate $\rho_{A}$ and become nonrelativistic as late as possible; Eq. 6 is relatively more constraining if relativistic longitudinal modes dominate. Figure 1 depicts the scaling of Eq. 6 for freely evolving dark photons in all possible regimes of wave number $k/am_{A}$ and Hubble rate $H/m_{A}$.
Figure 1: Redshifting of the dark photon’s contribution to the Higgs’s effective potential, $\propto m_{A}^{2}A_{\mu}A^{\mu}$, in various regimes of free evolution. The horizontal and vertical axes are the physical Hubble scale and the physical momentum of a given mode. Gray lines depict the trajectory of modes of a fixed comoving wave number in time (from the top right to the bottom left). The relationship between $H$ and $k/a$ fixes a radiation-dominated Universe with $H\propto a^{-2}$; modifications to the expansion history would merely alter the line separating the sub- and superhorizon regimes.
We derive a concrete bound on $g_{D}$ from Eq. 6 by taking dark photons to be produced during radiation domination (when the Hubble rate is $H_{\star}=1/2t_{\star}$) with typical wave number $k_{\star}$ and energy density $\rho_{A}(t_{\star})$. The energy density of the dominant modes scales like $a^{-4}$ until it becomes nonrelativistic when $\sqrt{H_{\mathrm{NR}}/H_{\star}}k_{\star}/a_{\star}=m_{A}$ (where $H\propto a^{-2}$ during radiation domination). Afterward, the dark photon’s energy density dilutes like matter ($\propto a^{-3}$). For a present-day abundance of dark photons $\Omega_{A}=\rho_{A}(t_{0})/\rho(t_{0})$ with $\rho(t_{0})$ the critical density, and $\Omega_{m}$ defined likewise for the total matter density,
$\frac{\Omega_{A}}{\Omega_{m}}\frac{3M_{\mathrm{pl}}^{2}H_{\mathrm{eq}}^{2}}{2}$ $\approx\left(\frac{H_{\mathrm{eq}}}{H_{\mathrm{NR}}}\right)^{3/2}\left(\frac{H_{\mathrm{NR}}}{H_{\star}}\right)^{2}\rho_{A}(t_{\star}),$ (7)
where $H_{\mathrm{eq}}$ is the Hubble rate at matter-radiation equality, equal to $2.26\times 10^{-28}\leavevmode\nobreak\ {\mathrm{eV}}$ in Planck’s best-fit cosmology [56]. Solving for $\rho_{A}(t_{\star})$ and plugging into Eq. 6 gives
$g_{D}$ $\lesssim 10^{-14}\,\lambda^{1/4}\left(\frac{m_{A}}{\mu{\mathrm{eV}}}\right)^{5/8}\left(\frac{m_{A}}{H_{\star}}\right)^{3/8}\left(\frac{\Omega_{A}}{\Omega_{m}}\right)^{-1/4}\begin{dcases}\left(\frac{H_{\mathrm{NR}}}{H_{\star}}\right)^{-1/8},&\rho_{A,\perp}\gg\rho_{A,\parallel},\\ \left(\frac{H_{\mathrm{NR}}}{H_{\star}}\right)^{1/8},&\rho_{A,\perp}\ll\rho_{A,\parallel}.\end{dcases}$ (8)
Independent of polarization, the bounds on $g_{D}$ are least severe when $H_{\star}$ is as small as possible, as illustrated by Fig. 2. That is, the later dark photons are produced, the lower their peak energy density.
Figure 2: Illustration of dark photon production that avoids backreaction onto the Higgs by delaying the time of production (see also Ref. [52]). After it is produced and becomes nonrelativistic, the dark matter has a known energy density at any scale factor $a$ (indicated by the green line) extrapolated from its present-day value. For any choice of model parameters, its energy density would exceed the threshold for backreaction (in the red shaded region) at some early time. By sufficiently delaying the production of dark photons, i.e., until some critical Hubble rate $H_{\star}$ (indicated by the blue line), they never backreact onto the Higgs (per the solid black curve). Dark photons produced too early (as in the dashed black curve) backreact onto the Higgs, possibly cutting off production before the total dark matter energy density is produced, and possibly collapsing into a string network that is not viable cold dark matter.
The case with purely transverse dark photons benefits from remaining relativistic until as late as is allowed, while longitudinal-mode production is least constrained if nonrelativistic at the start; the dependence on $H_{\mathrm{NR}}/H_{\star}$ is modest in either case. As we show in the following, models of nonthermal production typically take place when $H_{\star}\gtrsim m_{A}$, setting a useful benchmark.
Crucially, reducing the gauge coupling $g_{D}$ to avoid backreaction comes at the cost of also decreasing the strength of kinetic mixing between the dark and SM photons. In the simplest scenarios, kinetic mixing is generated by loops of fermions charged under both the visible and dark photons and is therefore proportional to their charge under both groups [57]:
$\varepsilon$ $\sim\frac{eg_{D}}{16\pi^{2}},$ (9)
with loop factor $16\pi^{2}$ chosen to match Ref. [48]. In Fig. 3, we depict bounds deriving from Eq. 8 for several benchmark scenarios in the $\varepsilon$-$m_{A}$ parameter space, superimposed with current limits [58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 94, 95, 96, 97, 98, 99, 100, 101, 102, 103, 104, 105, 106] and the prospective reach of DM-Radio [107], Dark E-field [108], ALPHA [109], MADMAX [110], BREAD [111], LAMPOST [112], SuperCDMS [113], and LZ [114].
Figure 3: Parameter space for kinetically mixed dark photons. Current exclusions from astrophysical [58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68] and cosmological [69, 70, 71, 72, 73] probes are depicted in dark gray and those from haloscope and other laboratory searches [76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 94, 95, 96, 97, 98, 99, 100, 101, 102, 103, 104, 105, 106] in light gray. The projected reaches [47] of DM-Radio [107], Dark E-field [108], ALPHA [109], MADMAX [110], BREAD [111], LAMPOST [112], SuperCDMS [113], and LZ [114] are depicted in yellow. The red regions depict various benchmarks thresholds for which the dark photon substantially backreacts onto the dark Higgs’s effective potential [Eq. 8] and can collapse into a string network, taking the fiducial kinetic mixing of Eq. 9. These benchmark scenarios include inflationary production [Eq. 12]; nonrelativistic, postinflationary production taking place when the Hubble $H_{\star}<m_{A}$; and a maximally conservative bound that only requires avoiding backreaction only around the time when scales relevant to CMB anisotropies enter the horizon.
Much of the open, experimentally accessible parameter space lies in the range $10^{-16}\lesssim\varepsilon\lesssim 10^{-10}$; only the lower end of this range, however, is within the reach of future searches for scenarios with $H_{\star}\gtrsim m_{A}$.
A maximally conservative bound might assume that the dark matter existed no earlier than directly evidenced by data—namely, no earlier than required by observations of the cosmic microwave background (CMB) anisotropies. The visible CMB is sensitive to scales that enter the horizon in the few decades of expansion around matter-radiation equality; we take $H_{\star}\approx 10^{-22}\leavevmode\nobreak\ {\mathrm{eV}}$ as a reasonable threshold. Figure 3 shows that a hypothetical scenario delaying production to so late a time is viable in nearly all the parameter space of future searches. Interestingly, bounds on dark photon dark matter from cosmological observations [69, 70, 71, 72, 73] (labeled “dark matter from cosmology/astrophysics” in Fig. 3) reside within this conservative exclusion region—that is, in the parameter space excluded by Refs. [69, 70, 71, 72, 73], dark photons with kinetic mixing Eq. 9 and a standard Higgs mass could not be the dark matter anyway.
II.1 Inflationary production
The most minimal mechanism for cosmological particle production is gravitational [115]. During inflation, minimally coupled degrees of freedom are excited as their wave number exits the horizon, so long as their action breaks conformal symmetry. Though the transverse polarizations of a vector are conformally symmetric up to their small mass $m_{A}\ll H_{I}$, the longitudinal mode of a massive vector is equivalent to a minimally coupled scalar field; its action thus violates conformal symmetry no matter how small its mass. Reference [50] showed that longitudinal modes exit the horizon during inflation with amplitude $kH_{I}/2\pi m_{A}$; per Fig. 1, they remain frozen until they either become massive (i.e., when $H\leq m_{A}$) or enter the horizon. A mode with wave number $k$ reenters the horizon at $a_{\times}=k/H(a_{\times})$ and becomes nonrelativistic at $a_{\mathrm{NR}}=k/m_{A}$. Assuming that the radiation era began immediately after inflation (so that $H\propto a^{2}$) and using Eq. 148, the energy density per logarithmic wave number thus scales as
$\frac{\mathrm{d}{\rho_{\parallel}}}{\mathrm{d}{\ln k}}\propto\frac{m_{A}^{2}}{a^{2}}\frac{\mathrm{d}{\left|A_{\parallel}\right|^{2}}}{\mathrm{d}{\ln k}}$ $\sim m_{A}^{2}\left(\frac{H_{I}}{2\pi}\right)^{2}\left(\frac{H(a)}{m_{A}}\right)^{3/2}\cdot\begin{dcases}\left(\frac{k}{a_{\star}m_{A}}\right)^{2},&k<a_{\star}m_{A},\\ \left(\frac{k}{a_{\star}m_{A}}\right)^{-1},&k>a_{\star}m_{A}\end{dcases}$ (10)
at late times $a>a_{\star}$ and $a>a_{\times}$. The present abundance of longitudinal dark photons produced gravitationally is thus dominated by the mode that enters the horizon when $H=m_{A}$ and takes the form [50]
$\Omega_{A}$ $=\Omega_{\mathrm{DM}}\left(\frac{m_{A}}{0.1\leavevmode\nobreak\ \mathrm{meV}}\right)^{1/2}\left(\frac{H_{I}}{5\times 10^{13}\leavevmode\nobreak\ {\mathrm{GeV}}}\right)^{2}.$ (11)
Dark photons with mass $m_{A}$ below $0.1\leavevmode\nobreak\ \mathrm{meV}$ cannot be sufficiently produced during inflation due to bounds on the tensor-to-scalar ratio from recent Planck and BICEP/Keck observations [116].333 Namely, taking the amplitude of the scalar power spectrum $A_{s}=2.1\times 10^{-9}$ [56] and the tensor-to-scalar ratio $r<0.036$ [116] bounds $H_{I}=\sqrt{\pi^{2}rA_{s}/2}M_{\mathrm{pl}}<4.7\times 10^{13}\leavevmode\nobreak\ \mathrm{GeV}$.
During inflation, unsuppressed formation of topological defects takes place if $H_{I}^{2}\gtrsim\mu=\pi v^{2}\ln(\min[\lambda/g_{D}^{2},\sqrt{\lambda}v/H_{I}])$ [117, 48], where $\mu$ is the string tension.444 References [118, 119] extended the study of inflationary dark photon production to account for a Higgs mass, though without considering the implications for direct detection. After inflation, these strings approach a scaling solution and absorb any remaining coherent dark photon field. Thus, if inflation is to account for vector dark matter, then the scale of inflation $H_{I}$ must be below the Higgs vev $v$, in turn bounding the dark gauge coupling by [48]
$g_{D}\lesssim\frac{m_{A}}{H_{I}}$ $\lesssim 2\times 10^{-22}\left(\frac{m_{A}}{{\mathrm{eV}}}\right)^{5/4}\left(\frac{\Omega_{A}}{\Omega_{\mathrm{DM}}}\right)^{-1/2},$ (12)
well out of range of any foreseeable experiment (Fig. 3).555 Ref. [48] points out that this bound may in fact be conservative, as even if string production is exponentially suppressed ($H_{I}\lesssim v$) strings may continue to grow in the background of the dark electromagnetic field to reach a scaling solution. In fact, this concern is also relevant for Eq. 8: large fluctuations, though exponentially unlikely, can produce strings. The precise threshold value of $H_{I}$ or $\rho_{A}(t_{\star})$ for which strings would be cosmologically relevant today requires further study.
The abundance in Eq. 11, as computed in Ref. [50], assumes instantaneous reheating, i.e., that the radiation era begins immediately following inflation. There is no guarantee that a prolonged era dominated by matter (or something more exotic) did not occur first after inflation, which modifies the relationship between the inflationary scale $H_{I}$ and the relic abundance when the dark photon’s mass $m_{A}$ exceeds the Hubble rate at reheating $H_{\mathrm{RH}}$. (Here by reheating we mean the moment the radiation era began, not necessarily that when the SM itself is populated and thermalized.) Modified expansion histories therefore shift the bound on $g_{D}$ in Eq. 12 (as well as the minimal allowed dark photon mass $m_{A}$ given an upper bound on $H_{I}$ from CMB observations).
II.1.1 Postinflationary expansion history
The results of Ref. [50] were extended to an arbitrary reheating temperature in Ref. [120, 121] and to arbitrary expansion histories before reheating in Ref. [122]. In the simplest inflationary scenarios, once slow roll ends, the inflaton retains the entirety of the Universe’s energy as it oscillates about the minimum of its potential (which is typically taken to be quadratic, such that the inflaton is matterlike in this period). The Universe only reheats once the inflaton’s dominant decay channel becomes efficient relative to the expansion rate. Since the inflaton’s energy density redshifts more slowly than the SM radiation it eventually decays into, the relative abundance of dark photons is lower for later reheating. Achieving the full relic abundance of the dark matter at a fixed dark photon mass then requires a larger inflationary scale $H_{I}$, such that avoiding defect formation imposes an even stronger limit on $g_{D}$. In fact, the relic abundance is independent of $m_{A}$ for $m_{A}>H_{\mathrm{RH}}$ [120, 121, 122], instead scaling with $\sqrt{H_{\mathrm{RH}}}H_{I}^{2}$. Requiring the dark photon to make up all of the dark matter then places a lower bound on the Hubble rate when the radiation era began as a function of $H_{I}$. The bound on the dark gauge coupling Eq. 12 is then penalized by a factor of $(m_{A}/H_{\mathrm{RH}})^{-1/4}$ for dark photons heavier than $H_{\mathrm{RH}}$ (Fig. 4).
More generally, any early epoch with equation of state $w<1/3$ further reduces the maximal gauge coupling that avoids inflationary defect formation. For stiffer equations of state $w>1/3$, the Universe’s energy density redshifts faster than radiation, instead easing constraints on $g_{D}$. The net enhancement in abundance in the $w\to 1$ limit is $\sim\sqrt{m_{A}/H_{\mathrm{RH}}}$ [122], requiring a lower inflationary scale to match the dark matter relic abundance and alleviating bounds on $g_{D}$ by a factor $4]{m_{A}/H_{\mathrm{RH}}}$. The minimum reheat temperature allowed by big bang nucleosynthesis (BBN) is of order $10\leavevmode\nobreak\ \mathrm{MeV}$ [123, 124, 125, 126, 127, 128, 129], corresponding to $H_{\mathrm{RH}}\approx 10^{-14}\leavevmode\nobreak\ {\mathrm{eV}}$ and an enhancement in allowed $g_{D}$ of about three orders of magnitude for $m_{A}\sim{\mathrm{eV}}$. Figure 4 shows that, while no viable parameter space is accessible to current or future searches assuming an instantaneous onset of radiation domination, a stiff era indeed marginally allows for kinetic mixing within reach of LZ [114] and only just below current exclusions from XENON [83, 90, 94, 95].
Figure 4: Kinetic mixing parameter space [Eq. 9] of dark photon dark matter produced during inflation. Red regions depict parameter space that are constrained by defect formation under various assumptions on the expansion history between inflation and nucleosynthesis, as labeled on the figure: the standard case of instantaneous reheating, an extreme case in which the Universe is kination-dominated ($w=1$) until BBN, and various extended epochs of matter domination with reheating temperatures indicated on the figure. Current limits from LAMPOST [91], FUNK [87], SENSEI [88], DAMIC [84], DarkSide [98], and XENON [83, 90, 94, 95] are depicted in gray and the projected reach [47] of BREAD [111], LAMPOST [112], SuperCDMS [113], and LZ [114] in yellow. Joint bounds from stellar probes are outlined in white.
On the other hand, if the Universe were instead matter-dominated for some or all of the epoch between inflation and BBN (a possibility one has little reason to exclude a priori), viable kinetic mixings are up to $\sim 2.5$ decades further out of reach of these experiments, depending on how late reheating occurs.666 Note that $T_{\mathrm{RH}}=444\leavevmode\nobreak\ \mathrm{GeV}$ is the lowest possible reheating temperature for which dark photons (of any mass) produced gravitationally during inflation may make up the entire dark matter abundance when taking $H_{I}<4.7\times 10^{13}\leavevmode\nobreak\ \mathrm{GeV}$ (see Footnote 3).
II.1.2 Nonminimal couplings
Nonminimal couplings to gravity may affect the relic abundance of dark photons generated during inflation [130, 131]. However, nonminimally coupled Proca theories of the form
$\mathcal{L}$ $=-\frac{1}{4}F_{\mu\nu}{F}^{\mu\nu}+\frac{1}{2}m_{A}^{2}A_{\mu}A^{\mu}-\frac{1}{2}\xi_{1}RA_{\mu}A^{\mu}-\frac{1}{2}\xi_{2}R^{\mu\nu}A_{\mu}A_{\nu}$ (13)
exhibit ghost instabilities [132, 133, 134]. References [135, 136] more recently discuss these runaway instabilities during inflation, arguing that they are present if the nonminimal couplings $\xi_{1}$ and $\xi_{2}$ are larger than $\sim(m_{A}/H_{I})^{2}$ in magnitude; the dynamics have only been solved numerically, preventing a simple analysis of the effect on defect formation bounds. References [135, 136] claim that radiative corrections in Abelian-Higgs theories generate nonminimal couplings of order $(m_{A}/m_{h})^{6}=g_{D}^{6}/(2\lambda)^{3}$. The runaway instability is then absent if $g_{D}/\sqrt{2\lambda}<\sqrt[3]{m_{A}/H_{I}}$, which is a dramatically weaker requirement than that to evade inflationary defect formation [Eq. 12].
Moreover, the nonminimal couplings in Eq. 13 must ultimately descend from gauge invariant operators. (Otherwise, technical naturalness would not protect the dark photon’s mass from large radiative corrections.) The only gauge-invariant operator linear in $A_{\mu}$ is $D_{\mu}\Phi$; the nonminimally coupled Abelian-Higgs theory therefore must take the form
$\mathcal{L}$ $=-\frac{1}{4}F_{\mu\nu}{F}^{\mu\nu}+\frac{1}{2}\left(g^{\mu\nu}-\frac{\xi_{1}}{m_{A}^{2}}Rg^{\mu\nu}-\frac{\xi_{2}}{m_{A}^{2}}R^{\mu\nu}\right)D_{\mu}\Phi\left(D_{\nu}\Phi\right)^{\ast}-\frac{\lambda}{4}\left({\left|\Phi\right|}^{2}-v^{2}\right)^{2}.$ (14)
As mentioned by Ref. [48], such a nonminimal coupling for the dark Higgs drives its VEV to larger values during inflation, which would mitigate defect formation if an increase in the inflationary scale were not required to compensate for a suppression incurred in the dark photon abundance (as derived in Sec. V.2.1). Furthermore, the effective Higgs mass squared is also suppressed by a factor $\sim m_{A}^{2}/\xi_{i}R$, not to mention the additional derivative couplings induced by the time-dependence of the coefficient of the Higgs’s kinetic term. Unless $m_{h}/H_{I}$ is even larger than $\sqrt{|\xi_{i}|}H_{I}/m_{A}$, the Higgs is thus light during inflation and would be produced gravitationally to a comparable degree; since the Higgs is much heavier at late times, it would dominate the dark sector density. When requiring the Higgs to remain heavy during inflation (i.e., $m_{A}^{2}m_{h}^{2}/|\xi_{i}|H_{I}^{4}>1$), if $|\xi_{i}|\sim(m_{A}/m_{h})^{6}$ as expected from radiative corrections then $|\xi_{i}|<(m_{A}/H_{I})^{3}$ and the runaway instability is absent.
Though Eq. 14 is written in terms of $\xi_{i}/m_{A}^{2}$ to match the form of Eq. 13, in reality we expect the coefficient of these dimension-6 operators to be $1/\Lambda^{2}$ where $\Lambda$ is the energy scale of the physics that generates the nonminimal couplings. (Written in terms of $\Lambda$, the $m_{A}\to 0$ limit explicitly restores gauge symmetry and the dark photon mass’s technical naturalness remains manifest.) In order that the low-energy theory remain a valid description of gravity during inflation, $H_{I}$ (if not the energy scale $\sqrt{H_{I}M_{\mathrm{pl}}}\gg H_{I}$) should be below $\Lambda$; one would thus expect that $\xi_{i}\sim(m_{A}/\Lambda)^{2}<(m_{A}/H_{I})^{2}$, even discounting powers of the dark gauge coupling expected from vertex factors. It is therefore not even clear whether the radiatively generated $\xi_{i}$ should be large enough to trigger the runaway instability in the first place without breaching the validity of the EFT.
II.1.3 Direct inflaton couplings
The dark photon (or one of the particles it interacts with) could also directly couple to the inflaton, enabling more efficient production of dark photon fluctuations and therefore a lower scale of inflation to achieve the dark matter relic abundance [137, 138, 139, 140, 141, 142, 143]. These mechanisms are nonetheless constrained at least to the same extent as any late time production mechanism, i.e., Eq. 8 with $H_{\star}=m_{A}$; however, the maximal energy density is necessarily higher than that assumed in Eq. 8 given that the dark photons were produced before the end of inflation. Referring to Fig. 1, $A_{\mu}A^{\mu}$ redshifts no more slowly than $a^{-2}$, and while $H>m_{A}$ all modes redshift as such except for subhorizon, longitudinal ones for which $A_{\mu}A^{\mu}\propto a^{-4}$. Taking a general equation of state $w$ in this epoch such that $H\propto a^{-3(1+w)/2}$, Eq. 8 is penalized by an additional factor $(m_{A}/H_{I})^{1/3(1+w)}$. If longitudinal modes enter the horizon at a scale factor $a_{\times}$ before $H=m_{A}$, a further penalty of $(a_{\times}/a_{\star})^{1/2}$ is incurred. These factors only account for the dark photons existing at the end of inflation; production earlier in inflation yields yet stronger bounds (but with model dependence that is difficult to capture on general grounds).
Similar scenarios where the dark photon instead couples to a rolling spectator field during inflation [142] are subject to the same bounds. However, Refs. [138, 139, 140, 142] considered only phenomenological parametrizations of the time dependence of the coupling function to the inflaton or spectator, neglecting the production that occurs (in concrete models) as the coupled scalar eventually oscillates about the minimum of its potential—i.e., preheating. This contribution most likely substantially modifies the results in these models, given that preheating would produce dark photons on scales as small as $k\sim a_{e}H_{I}$ where $a_{e}$ is the scale factor at the end of inflation. On the other hand, the results of Refs. [137, 141, 143] are limited to the linear regime in which backreaction effects are negligible; full nonlinear dynamics of such models were treated in Refs. [144, 145, 146, 147, 148, 149, 150, 151, 152, 153], albeit not specialized to dark photons specifically as the dark matter.
II.2 Irreducible thermal fluctuations
Dark photons that kinetically mix with the SM photon are inevitably thermally excited to some degree; depending on the thermal history of the Universe, these fluctuations can be large enough to restore symmetry to the dark Higgs such that topological defects form via the Kibble-Zurek mechanism [154]. The dynamics of the photon–dark-photon system are described by the Lagrangian
$\mathcal{L}$ $=-\frac{1}{4}F_{\mu\nu}F^{\mu\nu}-\frac{1}{4}{F_{\mathrm{SM}}}_{\mu\nu}{F_{\mathrm{SM}}}^{\mu\nu}+\frac{1}{2}m_{A}^{2}A_{\mu}A^{\mu}-\frac{\varepsilon}{2}F_{\mu\nu}{F_{\mathrm{SM}}}^{\mu\nu}-eJ_{\mathrm{SM}}^{\mu}{A_{\mathrm{SM}}}_{\mu},$ (15)
where $J_{\mathrm{SM}}$ is the current of electrically charged SM particles. Charged SM particles are sterile with respect to the combination $S=A+\varepsilon{A_{\mathrm{SM}}}$, making the $\{{A_{\mathrm{SM}}},S\}$ basis a convenient choice to discuss the propagation of (dark) photons in the SM plasma [58]. The dispersion relation of the active state $A_{\mathrm{SM}}$ receives corrections in the background of the Standard Model plasma, resulting in the well-known effective mixing angle between $A_{\mathrm{SM}}$ and $S$ [58],
$\epsilon_{\mathrm{eff},\lambda}^{2}$ $=\frac{\varepsilon^{2}m_{A}^{4}}{(m_{\lambda}^{2}-m_{A}^{2})^{2}+\omega^{2}\Gamma_{\lambda}^{2}}+\mathcal{O}(\varepsilon^{4}),$ (16)
where $m_{\lambda}$ and $\Gamma_{\lambda}$ are the plasma frequency and damping rate of SM photons with polarization state $\lambda$ and $\omega$ is the frequency of the photons.
At sufficiently early times, the plasma frequency of transverse modes far exceeds the bare mass of the dark photon, i.e., $m_{\perp}^{2}\propto e^{2}T^{2}\gg m_{A}^{2}$; the effective mixing angle is then suppressed relative to its vacuum value as $\varepsilon_{\mathrm{eff},\perp}\sim\varepsilon m_{A}^{2}/m_{\perp}^{2}$. In other words, the transverse photon $A_{\mathrm{SM,\perp}}$ is not only an interaction eigenstate but also nearly a propagation eigenstate. The sterile state $S$ is therefore suppressed relative to the SM photon by a factor $\varepsilon m_{A}^{2}/m_{\perp}^{2}$ and hence
$A_{\perp}$ $\approx-\varepsilon\left[1+\mathcal{O}(m_{A}^{2}/m_{\perp}^{2})\right]A_{\mathrm{SM,\perp}}.$ (17)
The dark photon fluctuations of order $\varepsilon A_{\mathrm{SM,\perp}}$ do not set the late-time relic dark matter abundance—they are strongly coupled to the plasma and continually reabsorbed—rather, the sterile state itself adiabatically transitions into the dark photon as the Universe cools and the plasma frequency drops. These $A_{\perp}$ fluctuations do, however, backreact onto the dark Higgs, and at sufficiently large temperature and dark gauge coupling the small, thermal dark photon fluctuations are enough to restore symmetry globally to the Higgs. That is, the active state $A_{\mathrm{SM}}$ receives a small effective mass from the Higgs field $\varepsilon g_{D}|\Phi|$, leading to a thermal effective potential for the Higgs [155] $V_{\mathrm{eff}}(|\Phi|^{2})\supset\varepsilon^{2}g_{D}^{2}|\Phi|^{2}T^{2}/24$. This effective potential restores symmetry to the Higgs when Standard Model temperature is sufficiently large, and consequently defects form by the Kibble-Zurek mechanism, precluding the formation of dark photon dark matter at some later epoch. Requiring that symmetry remains broken when SM temperature is its largest ($T_{\mathrm{max}}$) sets a bound
$\varepsilon$ $\lesssim 1.1\times 10^{-6}\lambda^{1/6}\left(\frac{m_{A}}{{\mathrm{\mu eV}}}\right)^{1/3}\left(\frac{T_{\mathrm{max}}}{10\leavevmode\nobreak\ {\mathrm{MeV}}}\right)^{-1/3}\left(\frac{16\pi^{2}\varepsilon}{eg_{D}}\right)^{2/3}.$ (18)
This limit is depicted in Fig. 10 and discussed in Sec. VI.
II.3 Axion oscillations and tachyonic resonance
While inflationary and thermal dark photon production are both minimal scenarios in their physical content and assumptions, neither can produce dark photon dark matter at small masses and large couplings—nor can either produce viable dark photon dark matter below an meV at all. Detectably large kinetic mixing in the sub-meV mass range requires a nonthermal, postinflationary, and nonminimal production mechanism (or an ad hoc mechanism to boost $\varepsilon$ relative to $g_{D}$; see Sec. III). One possibility is that the dark photon inherits the relic abundance of a misaligned scalar field via resonant particle production. A canonical example is a coupling to an axion $\phi_{a}=f_{a}\theta$ with mass $m_{a}$ and decay constant $f_{a}$ through the axial term $\tilde{F}_{\mu\nu}F^{\mu\nu}/4=\mathbf{E}\cdot\mathbf{B}$ [54, 156, 157, 158],
$\mathcal{L}\supset\frac{1}{2}f_{a}^{2}\partial_{\mu}\theta\partial^{\mu}\theta-V(\theta)+\frac{\beta}{4}\theta\tilde{F}_{\mu\nu}F^{\mu\nu},$ (19)
where $\tilde{F}_{\mu\nu}\equiv\epsilon_{\mu\nu\rho\sigma}F^{\rho\sigma}$ is the dual field strength tensor. The axion lives in a compact field space, and its potential often takes the simple periodic form
$V$ $=m_{a}^{2}f_{a}^{2}\left(1-\cos\theta\right).$ (20)
After a period of inflation with $m_{a}\ll H_{I}\ll f_{a}$, the axion field assumes a nearly homogeneous initial condition $\theta(0,\mathbf{x})\equiv\theta_{0}$ selected from the interval $[0,2\pi)$ with approximately uniform probability. To linear order in spatial fluctuations, the equations of motion for the circular polarizations of the dark photon $A_{\pm}$ [Eq. 139c] and the spatially averaged axion $\bar{\theta}(t)$ are
00 0 $=\ddot{\theta}+3H\dot{\theta}+m_{a}^{2}\theta-\frac{\beta}{f_{a}^{2}}\langle\mathbf{E}\cdot\mathbf{B}\rangle$ (21a)
00 0 $=\ddot{A}_{\pm}+H\dot{A}_{\pm}+\left(\frac{k^{2}}{a^{2}}+m_{A}^{2}\pm\beta\dot{\bar{\theta}}\frac{k}{a}\right)A_{\pm},$ (21b)
neglecting nonlinearities in the axion potential.
Provided $\beta$ is large enough that $\beta\dot{\bar{\theta}}\geq 2m_{A}$, the effective frequency of one of the two transverse polarization states is negative, leading to the exponential growth of $A_{\pm}$ due to a tachyonic resonance [54, 156]. In the simplest scenarios, the axion’s damped and (nearly) harmonic oscillations begin around the time when $H=m_{a}$. The maximum velocity of an axion oscillating in a quadratic potential is set by its mass, i.e. $|\dot{\bar{\theta}}|\leq m_{a}\theta_{0}$; tachyonic resonance then occurs if $\beta\theta_{0}>2m_{A}/m_{a}$. A more precise determination of the conditions for efficient resonance requires solving the equations of motion (see Appendix D); the so-called “broad resonance” regime requires
$\beta\theta_{0}$ $\gg\max\left\{1,4m_{A}/m_{a}\right\}.$ (22)
In the broad resonance regime, the dark photon mode that grows the most has wave number $\mathcal{O}(\beta m_{a})$.777 The dynamics of broad parametric resonance via an oscillating, parity-even scalar [51] (or the dark Higgs itself [159]) are qualitatively similar. Production via scalars does allow for a number of qualitatively distinct regimes, which we discuss in detail in Sec. V.
Assuming $\theta_{0}\sim 1$, efficient dark photon production requires $\beta$ larger than unity to some degree. If $\beta$ is just large enough for production to be efficient [Eq. 22], then by kinematics the dark photon cannot be heavier than the axion and production must occur at $H_{\star}\gtrsim m_{A}$. This regime is well studied [54, 156, 157, 158]. The resulting limit on the kinetic mixing that avoids backreaction onto the Higgs is displayed in Fig. 3 [assuming the fiducial relationship between $g_{D}$ and $\varepsilon$, Eq. 9] and coincides with the constraint taken in Ref. [54] to ensure backreaction onto the Higgs’s mass is negligible. Production at $H_{\star}\gtrsim m_{A}$ does offer substantially improved detection prospects compared to minimal inflationary production. Notably, however, most of the experimental prospects are only for masses above an ${\mathrm{meV}}$—a range already accessible to inflationary production (even if ad hoc enhancements of the kinetic mixing would be required for direct detection; see Sec. III). Thus, the sub-${\mathrm{meV}}$ mass range (as probed by axion haloscopes) remains out of reach without additional machinery.
When the dark photon’s energy density becomes comparable to that of the axion, it backreacts onto the axion via nonlinear processes, altering the rate and direction of energy transfer. For resonant production mechanisms from massive, oscillating (pseudo)scalars in general, the final energy partition between the dark photon and the scalar ultimately must be quantified by numerical simulations. Those presented in Ref. [54] show that whether the majority of the axion’s energy density ends up in the dark photon depends on the mass ratio and becomes less efficient as $m_{a}/m_{A}$ becomes large, though the range of ratios that have been simulated is limited. The nonlinear dynamics of backreaction do not generically yield a large hierarchy in abundance; Ref. [156] discusses a number of model-dependent remedies to dilute any remaining axion energy density.
In standard constructions where the axion-photon coupling is generated via Peccei-Quinn (PQ) fermions, order-unity couplings $\beta$ are not the natural expectation. The anomaly coefficient to the PQ charge of the $\mathrm{U}(1)_{D}$-charged fermions is [160, 54]
$C_{\theta AA}$ $\sim\sum_{i=1}^{N_{f}}Q_{i}^{2}\sim N_{f}Q_{f}^{2},$ (23)
where $Q_{i}$ is the PQ charge of the $i$th fermion, and $N_{f}$ is the total number of fermions. In terms of $C_{\theta AA}$ and the dark fine-structure constant $\alpha_{D}=g_{D}^{2}/4\pi$, the axial coupling is
$\beta$ $=C_{\theta AA}\frac{\alpha_{D}}{2\pi}.$ (24)
For modest charges and number of PQ fermions, $\beta\sim g_{D}^{2}\ll 1$, in extreme tension with the requirement for tachyonic resonance [Eq. 22]. Namely, for gauge couplings that evade backreaction onto the Higgs [Eq. 8],
$C_{\theta AA}$ $\gtrsim 10^{31}\frac{\beta}{10}\sqrt{\frac{\Omega_{A}}{\Omega_{\mathrm{DM}}}}\left(\frac{m_{A}}{\mu{\mathrm{eV}}}\right)^{-5/4}\left(\frac{m_{A}}{m_{a}}\right)^{-1/4}.$ (25)
References [161, 160, 54] discuss several model-building avenues to enhance the axial coupling relative to the nominal expectation, including taking $N_{f}$ and $Q_{f}$ large as well as the clockwork mechanism [162]. We discuss clockwork and its consequences for backreaction onto the Higgs in more detail in Sec. III. Abelian clockwork may also simultaneously account for large hierarchies between $\varepsilon$ and $g_{D}$ and between $\beta$ and $\alpha_{D}$ [54].
II.4 Vector misalignment
Much like scalar fields, the energy density of a homogeneous vector field dilutes as $a^{-3}$ when $H\lesssim m_{A}$, and the resulting condensate behaves as cold dark matter [163]. But in contrast to a scalar, a vector’s energy density dilutes as $a^{-2}$ at early times ($H\gtrsim m_{A}$) when it is not oscillating, as can be inferred from the $k/a\ll m_{A}$ limit in Fig. 1. Successful vector misalignment after inflation therefore requires a nonminimal coupling to gravity of the form Eq. 13 with $\xi_{1}=-1/6$ and $\xi_{2}=0$ to avoid this additional $a^{-2}$ dilution [69]. Setting aside the theoretical issues with nonminimal couplings discussed in Sec. II.1.2, vector misalignment provides a useful benchmark for nonthermal production of dark photons—in particular, one whose initial conditions lack magnetic fields.
In the absence of spatial gradients, the temporal component $A_{0}$ vanishes and the dark photon is purely electric. Defining the rescaled field $\tilde{A}_{i}=\bar{A}_{i}/a$, each component of $\tilde{A}_{i}$ satisfies the equation for the evolution of a homogeneous scalar with additional terms proportional to $1+6\xi_{1}\equiv 1-\kappa$:
00 0 $=\ddot{\tilde{A}}_{i}+3H\dot{\tilde{A}}_{i}+\left[m_{A}^{2}+(1-\kappa)\left(\dot{H}+2H^{2}\right)\right]\tilde{A}_{i}.$ (26)
During inflation, the solutions to this equation all decay exponentially if $\kappa<1$, while one instead grows exponentially if $\kappa>1$; the vector field $\tilde{A}_{i}$ is only constant if $\kappa$ is tuned very close to one. Fixing $\kappa=1$, vector misalignment and scalar misalignment are mathematically identical in vacuum.
A homogeneous dark photon field contains no magnetic flux, offering misalignment a possible loophole to avoid the defect bound Eq. 5. On the other hand, if $m_{A}^{2}\tilde{A}_{i}(0)^{2}\gtrsim\lambda v^{4}$, the dark photon condensate causes the Higgs field to oscillate about the symmetric point $\Phi=0$; the dark photon itself then slowly rolls due to its oscillating (and decreasing) mass. Perturbations in one of $\Phi$ or $\Phi^{\ast}$ experience a tachyonic instability that ultimately fragments the Higgs, although whether misalignment would seed topological defects is a question that can only be addressed by numerical simulations (see Appendix B of [48]).
III Clockworked couplings and backreaction
The fiducial kinetic mixing Eq. 9 is comparable to the dark gauge coupling, as prescribed by the simple, canonical model that kinetic mixing is generated by $N_{f}$ heavy fermions charged under both the SM and dark gauge groups [57, 164]:
$\varepsilon$ $=\frac{eg_{D}}{12\pi^{2}}\sum_{i=1}^{N_{f}}Q_{e,i}Q_{D,i}\ln\frac{m_{i}^{2}}{\mu^{2}},$ (27)
where $m_{i}$ are the masses of the so-called “portal matter” fermions, each with charges $Q_{e,i}$ and $Q_{D,i}$ under electromagnetism and the dark $\mathrm{U}(1)$, respectively. The additional constraint
$\sum_{i=1}^{N_{f}}Q_{e,i}Q_{D,i}$ $=0$ (28)
is often imposed so that $\varepsilon$ is independent of the renormalization scale $\mu$. Assuming that the portal matter masses are comparable and their charges are order unity leads to the parametric estimate of Eq. 9. Clearly, taking a large number of fermions and/or their charges large enhances $\varepsilon/g_{D}$. A less contrived possibility, perhaps, is the clockwork mechanism [162, 48]. In this section we consider the viability of clockwork as a means to evade backreaction onto the Higgs and possible defect formation when invoked both to enhance the kinetic mixing itself (Sec. III.1) and to generate the hierarchically large axion couplings required for successful tachyonic production (Sec. III.2), as discussed in Sec. II.3.
The Abelian clockwork mechanism employs a chain of $N+1$ $\mathrm{U}(1)$ gauge symmetries that are broken to a single $\mathrm{U}(1)$ when $N$ charged scalars (Higgses) acquire VEVs. Each gauge field $A_{i}$ is labeled by an integer $i\in\{1,\ldots,N+1\}$, and each Higgs $\Phi_{i,i+1}$ is charged under two neighboring $\mathrm{U}(1)$s on the chain with charges $1$ and $-Q$. We denote the additional dark Higgs that couples only to the dark photon $A$ (and generates its mass) by $\Phi$ without subscripts, and for clarity we never refer to the clockwork gauge fields without an index subscript (leaving $A$ to denote only “the” dark photon). The clockwork Lagrangian is explicitly
$\begin{split}\mathcal{L}&=\frac{1}{2}\left|\left(\partial^{\mu}-\mathrm{i}gA_{N+1}^{\mu}\right)\Phi\right|^{2}+\frac{\lambda}{4}\left(|\Phi|^{2}-v^{2}\right)^{2}+gJ_{\mu}A_{1}^{\mu}\\ &\hphantom{{}={}}-\frac{1}{4}\sum_{i=1}^{N+1}F^{\mu\nu}_{i}F^{i}_{\mu\nu}+\sum_{i=1}^{N}\left[\frac{1}{2}\left|D_{i,i+1}\Phi_{i,i+1}\right|^{2}-\frac{\lambda_{i,i+1}}{4}\left(\left|\Phi_{i,i+1}\right|^{2}-v_{i,i+1}^{2}\right)^{2}\right],\end{split}$ (29)
where the gauge covariant derivative is
$D_{i,i+1}^{\mu}$ $=\partial^{\mu}+\mathrm{i}gA_{i}^{\mu}-\mathrm{i}QgA_{i+1}^{\mu}$ (30)
and $J_{\mu}$ is the current of particles charged under $A_{1}$ (e.g. the portal fermions). We also take all gauge couplings $g$ and charges $Q$ independent of $i$ for simplicity. We take the $N$ Higgses $\Phi_{i,i+1}$ to have VEVs $v_{i,i+1}$ larger than the VEV $v$ of the Higgs $\Phi$ that ultimately breaks the last $\mathrm{U}(1)$ and gives the dark photon its mass.
The effective potential for the gauge fields is a sum of independent quadratics $g^{2}\sum_{i}v_{i,i+1}^{2}(A_{i}-QA_{i+1})^{2}$ that is minimized when each term in the sum is independently zero. At low enough energies, integrating out the heavy Higgses and heavy gauge fields therefore sets $|\Phi_{i,i+1}|=v_{i,i+1}$ and $A_{i}=QA_{i+1}$ for all $i<N+1$. The resulting low-energy effective theory is
$\mathcal{L}$ $=\frac{1}{2}\left|\left(\partial_{\mu}-\mathrm{i}g_{D}A_{\mu}\right)\Phi\right|^{2}+\frac{\lambda}{4}\left(|\Phi|^{2}-v^{2}\right)^{2}-\frac{1}{4}F^{\mu\nu}F_{\mu\nu}+g_{D}Q^{N}J_{\mu}A^{\mu},$ (31)
where the canonically normalized dark photon field is identified as
$A^{\mu}$ $=\sqrt{\sum_{i=0}^{N}Q^{2i}}A_{N+1}^{\mu}=\sqrt{\frac{Q^{2(N+1)}-1}{Q^{2}-1}}A_{N+1}^{\mu}$ (32)
with a dark gauge coupling
$g_{D}$ $=g\sqrt{\frac{Q^{2}-1}{Q^{2(N+1)}-1}}.$ (33)
Particles charged under the first clockwork site have an effective charge under the dark photon $Q^{N}$ times larger than their charge under $A_{1}$ itself, exponentially enhancing the kinetic mixing relative to the Higgs coupling $g_{D}$:
$\varepsilon$ $\sim Q^{N}\frac{eg_{D}}{16\pi^{2}}.$ (34)
Implementing clockwork to explain large kinetic mixing with a small effective charge for the dark Higgs $\Phi$ is the reverse of the traditional invocation—to account for a small kinetic mixing with order-unity gauge couplings.
While Eq. 31 takes the form of the standard Abelian-Higgs model Eq. 1 (plus a current), it is only strictly correct while the final gauge symmetry remains unbroken. After $\Phi$ acquires its VEV, the ground state configuration of the clockwork gauge fields is modified from that above ($A_{i}=QA_{i+1}$ for all $i<N+1$) because the effective potential has an additional contribution $g^{2}v^{2}A_{N+1}^{2}/2$. Whereas in Eq. 31 the massless mode couples to no clockwork Higgses in the low-energy theory, the lightest mass eigenstate [after the final $\mathrm{U}(1)$ is broken] couples to them all—and, with a large abundance, can in principle backreact strongly on any of the clockwork Higgses. In Sec. III.1 we assess whether the dark photon backreacts on any of the clockwork Higgses at a lower threshold than it does onto the dark Higgs $\Phi$, i.e., whether clockwork modifies the backreaction condition Eq. 5. In Sec. III.2 we then consider whether the heavier clockwork gauge fields themselves may be produced (as a byproduct of the dynamics that produce the lightest mode, the dark photon) and themselves trigger backreaction onto any of the clockwork scalar fields.
III.1 Large kinetic mixing
In the absence of the additional dark Higgs $\Phi$, the massless mode of the clockwork fields is precisely that which couples to none of the clockwork Higgses $\Phi_{i,i+1}$. Breaking the last gauge symmetry (with $\Phi$) only generates effective couplings to the clockwork Higgses by perturbing the lightest mode in the clockwork spectrum away from the would-be massless mode identified in Eq. 31. The size of the dark photon’s couplings to the clockwork Higgses must therefore be controlled by that of the dark Higgs, $g_{D}$. Since $v_{i,i+1}\gtrsim v$, the threshold dark photon density at which any of the other clockwork gauge symmetries are restored is manifestly higher than that of the dark photon and dark Higgs, i.e., the dark Higgs still provides the strongest condition to evade backreaction and defect formation.
To verify this expectation, we explicitly compute the dark photon’s couplings to the clockwork Higgses by expanding in small $v/v_{i,i+1}$. The gauge fields have an unperturbed mass matrix
$M^{2}$ $=g^{2}\begin{pmatrix}v_{1,2}^{2}&-Qv_{1,2}^{2}\\ -Qv_{1,2}^{2}&Q^{2}v_{1,2}^{2}+v_{2,3}^{2}&-Qv_{2,3}^{2}\\ &\ddots&\ddots&\ddots\\ &&-Qv_{N-1,N}^{2}&Q^{2}v_{N-1,N}^{2}+v_{N,N+1}^{2}&-Qv_{N,N+1}^{2}\\ &&&-Qv_{N,N+1}^{2}&Q^{2}v_{N,N+1}^{2}\end{pmatrix}.$ (35)
Independent of the VEVs of the $\Phi_{i,i+1}$, by inspection this matrix always has a null eigenvector
$\mathbf{e}_{0}$ $=\sqrt{\frac{1-Q^{-2}}{1-Q^{-2(N+1)}}}\left[1,Q^{-1},\ldots,Q^{-N}\right],$ (36)
which, for $v=0$, is the (massless) dark photon field $A$. However, specifying the other eigenvectors requires specifying the VEVs of the clockwork Higgses. To simplify the calculation, we assume all the VEVs are the same, $v_{i,i+1}=\mathrm{v}$, in which case the remaining eigenvectors have the form [162]
$[\mathbf{e}_{k}]_{j}$ $=\sqrt{\frac{2g^{2}\mathrm{v}^{2}}{(N+1)m_{k}^{2}}}\left(Q\sin\frac{\pi jk}{N+1}-\sin{\frac{\pi(j+1)k}{N+1}}\right),$ (37)
where the mass-squared eigenvalues are
$m_{k}^{2}$ $=g^{2}\mathrm{v}^{2}\left[Q^{2}+1-2Q\cos\frac{\pi k}{N+1}\right].$ (38)
Note that all the masses of the broken $\mathrm{U}(1)$s are parametrically $g^{2}Q^{2}\mathrm{v}^{2}$.
When $v\ll\mathrm{v}$, breaking the final gauge symmetry modifies the mass matrix Eq. 35 by a small perturbation
$\delta M^{2}$ $=g^{2}\operatorname{diag}\left(0,\ldots,0,v^{2}\right).$ (39)
The perturbation to the eigenvectors can be computed at first order in $v^{2}/\mathrm{v}^{2}$ as
$\delta{\mathbf{e}}_{k}$ $=\sum_{j\neq k}\frac{\mathbf{e}_{k}^{T}\delta M^{2}\mathbf{e}_{j}}{m_{k}^{2}-m_{j}^{2}}\mathbf{e}_{j}.$ (40)
The lightest of these states, which we identify as the dark photon, is then
$\left[\delta\mathbf{e}_{0}\right]_{i}$ $=-\left(\frac{v}{\mathrm{v}}\right)^{2}\frac{g_{D}}{g}\frac{2Q}{N+1}\sum_{j=1}^{N}\sin\frac{\pi Nj}{N+1}\frac{Q\sin\frac{\pi ij}{N+1}-\sin{\frac{\pi(i+1)j}{N+1}}}{\left[Q^{2}+1-2Q\cos\frac{\pi j}{N+1}\right]^{2}}.$ (41)
We may straightforwardly infer the dark photon’s couplings to all the clockwork Higgses via their kinetic terms, projecting each clockwork photon onto the dark photon (i.e., the mass eigenstate $\mathbf{e}_{0}+\delta\mathbf{e}_{0}$):
$\mathcal{L}$ $\supset\frac{1}{2}g^{2}A_{\mu}A^{\mu}\sum_{i=1}^{N}\left(\left[\delta\mathbf{e}_{0}\right]_{i}-Q\left[\delta\mathbf{e}_{0}\right]_{i+1}\right)^{2}|\Phi_{i,i+1}|^{2},$ (42)
which we empirically observe to evaluate to
$\mathcal{L}$ $\supset\frac{1}{2}g_{D}^{2}\left(\frac{v}{\mathrm{v}}\right)^{4}A_{\mu}A^{\mu}\sum_{i=1}^{N-1}\left(Q^{N-i}\frac{Q^{2(i+1)}-1}{Q^{2(N+1)}-1}\right)^{2}|\Phi_{i,i+1}|^{2}.$ (43)
In the limit of large $Q^{2}$,
$\mathcal{L}$ $\supset\frac{1}{2}g_{D}^{2}\left(\frac{v}{\mathrm{v}}\right)^{4}A_{\mu}A^{\mu}\sum_{i=1}^{N-1}\frac{|\Phi_{i,i+1}|^{2}}{Q^{2(N-i)}}=\frac{1}{2}m_{A}^{2}A_{\mu}A^{\mu}\left(\frac{v}{\mathrm{v}}\right)^{2}\sum_{i=1}^{N-1}\frac{|\Phi_{i,i+1}/\mathrm{v}|^{2}}{Q^{2(N-i)}}.$ (44)
The dark photon $A$ backreacts significantly onto to $\Phi_{i,i+1}$ when the corresponding terms in Eq. 44 exceed $\lambda\mathrm{v}^{4}$. Not only is the dark photon’s contribution to the clockwork Higgses’ effective potentials in Eq. 44 suppressed relative to that for the dark Higgs $\Phi$ by $(v/\mathrm{v})^{2}$ and at least two powers of $Q$, their backreaction thresholds are parametrically higher by a factor of $(\mathrm{v}/v)^{4}$. If only the $A$ field is produced and no other eigenstate, clockwork therefore remains a viable means to boost kinetic mixing without requiring a large hierarchy $\mathrm{v}/v\sim g/g_{D}$, contrary to the claim made in Appendix C.1 of Ref. [48].
III.2 Large axion coupling
While the preceding section establishes that the dark photon itself does not backreact strongly onto any of the clockwork Higgses before it does so to the dark Higgs, any production channel via an interaction-basis gauge field could in principle produce mass eigenstates other than the dark photon. (Production could simply occur via a coupling directly to the linear combination of clockwork fields that make up the lightest mode, but this possibility seems rather contrived.) So long as the clockwork gauge fields’ masses lie above a kinematic threshold for production, these heavier states won’t be directly produced. In principle, however, the interactions directly responsible for production may also generate interactions between the dark photon and other mass eigenstates.
We now show that such effects are negligible for production of dark photons from an oscillating axion (Sec. II.3) as a concrete example. Efficient dark photon production from an oscillating axion typically requires couplings $\beta$ that are parametrically large compared to that expected to arise by integrating out heavy, PQ-charged fermions charged under the $\mathrm{U}(1)$ gauge group of $A$, for which $\beta\propto g_{D}^{2}\ll 1$. If these fermions couple to the more strongly coupled end of the clockwork chain [as in Eq. 31], then $\beta$ is enhanced to $\beta\propto Q^{2N}g_{D}^{2}\sim g^{2}$, after which only a moderate enhancement to $\beta$ is necessary to reach $\beta\gtrsim 40$ (see the discussion in Sec. IV and Ref. [54]).
Following the notation of the previous section, the Chern-Simons term for the clockwork vector $A_{1}^{\mu}$ expands to
$F_{1}^{\mu\nu}\tilde{F}^{1}_{\mu\nu}$ $=\sum_{i,j=0}^{N}[\mathbf{e}_{i}]_{1}[\mathbf{e}_{j}]_{1}\mathsf{F}_{i}^{\mu\nu}\tilde{\mathsf{F}}^{j}_{\mu\nu},$ (45)
where we use $\mathsf{A}_{i}^{\mu}$ to denote the $i$th mass-basis vector and $\mathsf{F}_{i}^{\mu\nu}$ its field strength. The equations of motion for the mass-basis fields are then
$\ddot{\mathsf{A}}_{i,\pm}+H\dot{\mathsf{A}}_{i,\pm}+\left[\frac{k^{2}}{a^{2}}+m_{i}^{2}\pm\beta\dot{\theta}\frac{k}{a}\left([\mathbf{e}_{i}]_{1}\right)^{2}\right]\mathsf{A}_{i,\pm}$ $=\mp\beta\dot{\theta}\frac{k}{a}\sum_{j=0,\,j\neq i}^{N}[\mathbf{e}_{i}]_{1}[\mathbf{e}_{j}]_{1}\mathsf{A}_{j,\pm}.$ (46)
Following Sec. II.3, if the massive modes are sufficiently heavy ($m_{i}\sim g\mathrm{v}>\beta\dot{\theta}\sim\beta m_{a}\theta_{0}$), only the lightest mode (the dark photon) is produced. The mixing term on the right-hand side of Eq. 46 is therefore dominated by the $j=0$ term in the sum. This source generates heavy modes with amplitude of order
$\mathsf{A}_{i,\pm}$ $\sim\frac{\beta\dot{\theta}k/a}{k^{2}/a^{2}+m_{i}^{2}}[\mathbf{e}_{i}]_{1}[\mathbf{e}_{0}]_{1}\mathsf{A}_{0,\pm}\propto\left(m_{a}/m_{i}\right)^{2}[\mathbf{e}_{i}]_{1}[\mathbf{e}_{0}]_{1}\mathsf{A}_{0,\pm},$ (47)
noting that $\dot{\theta}\sim m_{a}\theta_{0}$ and $k/a\sim m_{a}$ in the broad resonance regime. The eigenvector elements are each no larger in magnitude than $\sim 1/Q$, so the backreaction thresholds for the clockwork Higgses are order $\lambda\mathrm{v}^{4}\gtrsim m_{i}^{2}A_{i}^{\mu}A^{i}_{\mu}\sim m_{a}^{4}\mathsf{A}_{0,\pm}^{2}/m_{i}^{2}$. Taking the dark photon’s own kinematic threshold from Eq. 22 for order-unity axion misalignments, $m_{a}\gtrsim m_{A}/\beta$, shows that avoiding backreaction onto the clockwork Higgses requires $\lambda\mathrm{v}^{4}\gtrsim(m_{A}/m_{i})^{2}/\beta^{4}\cdot m_{A}^{2}\mathsf{A}_{0,\pm}^{2}$. Since $\beta>1$ and $m_{A}<m_{i}$ (and $v<\mathrm{v}$) by construction, avoiding backreaction onto the dark Higgs for the dark photon itself is a parametrically stronger condition than for the other clockwork gauge symmetries; the backreaction bound Eq. 5 is therefore not modified by clockwork.
IV Detectable dark photons from axions
As discussed in Sec. II.3, the dimensionless axion-photon coupling is typically $\beta\sim\alpha_{D}/2\pi\ll 1$, and without additional model building, tachyonic resonance from axion oscillations does not occur. Suitable model building can generate a hierarchy in $\beta/\alpha_{D}$, as achieved by Abelian clockwork mechanism described in Sec. III. The enhancement from clockwork alone, however, is limited to $\beta\lesssim 1$ by perturbative unitarity (i.e., of scatterings with the fermions with largest charge); an additional mechanism must further enhance $\beta$ to the values $\gg 10$ required for efficient resonance [161, 160, 54]. Even taking a viable origin of arbitrarily large $\beta$ for granted, backreaction constraints still limit production in much of the experimentally accessible parameter space (Fig. 3), since production typically occurs no later than when $H\approx m_{A}$. The same mechanisms that boost $\beta$ can also enhance $\varepsilon$ relative to $g_{D}$, and in principle axions can produce dark photons across all experimentally accessible parameter space [54, see also Sec. III.2]. However, a $10^{30}$ enhancement to the axion-photon coupling is rather implausible a priori [see Eq. 25].
In this section, we explore scenarios that simultaneously facilitate efficient dark photon production and weaken backreaction and defect formation limits. We demonstrate that models that delay production (thereby relaxing these bounds, as described in Sec. II) can also enable efficient tachyonic resonance at smaller values of $\beta$. In this regime, both objectives work in tandem: delayed production allows for larger $g_{D}$, raising the expected axion coupling $\alpha_{D}/2\pi\propto g_{D}^{2}$, and also reduces the degree to which $\beta$ must be enhanced relative to this expectation. We also consider how the opposite regime—strong coupling—can itself delay production and may require less model-building machinery than the moderately large $\beta$ values considered in prior literature. Figure 5 illustrates these various dynamical regimes in parameter space.
Figure 5: Illustration of the parameter space in which an axion may transfer its energy efficiently to a dark photon. The green region represents the usual range of moderate axion couplings considered, for example, in Ref. [54, 156, 159]. The blue region corresponds to the narrow resonance discussed in Sec. IV.1 and Ref. [158], in which delayed axion oscillations lead to a more efficient resonance and thus allow for $\beta\ll 1$. In the yellow region, the axion–dark-photon coupling is so large that the axion’s oscillations are overdamped and energy transfer takes a parametrically long time. A rigorous understanding of this regime, however, requires numerical simulations (see Sec. IV.2). Such large couplings could possibly be realized if the axion is coupled to the dark photon in the magnetic basis [165, 166], but this possibility also remains speculative and requires further model-building study.
IV.1 Narrow resonance
In the standard scenario described in Sec. II.3, only dark photon modes with $m_{A}<k/a<\beta|\dot{\bar{\theta}}|\propto a^{-3/2}$ undergo tachyonic growth. As the Universe expands, modes continuously exit the resonance band since their wave numbers redshift more slowly than does the width of the resonance band itself. Because the expansion rate $H$ is only just below the axion’s frequency $m_{a}$, resonance is efficient only during the axion’s first few oscillations before unstable modes redshift out of the band and are no longer tachyonic. However, if axion oscillations begin when $H$ is parametrically lower than $m_{a}$, modes grow over a proportionally larger number of oscillations, making resonance efficient at smaller $\beta\theta_{0}$.
As derived in Sec. D.1 and Ref. [158], when axion oscillations begin at an arbitrary Hubble rate $H_{\mathrm{osc}}<m_{a}$, the resonance condition relaxes to
$\beta\theta_{0}\sqrt{\frac{m_{a}}{2H_{\mathrm{osc}}}}\gg 1.$ (48)
Delaying axion oscillations until $H\ll m_{a}$ thus not only reduces the required enhancement of the axion coupling $\beta$, but also weakens constraints on the dark gauge coupling from backreaction onto the Higgs [per Eq. 8]. Because resonance is narrow in wave number about $a_{\mathrm{osc}}m_{a}/2$, resonantly produced dark photons become nonrelativistic when $H_{\mathrm{NR}}/H_{\star}=(a_{\star}/a_{\mathrm{NR}})^{2}=(2m_{A}/m_{a})^{2}$. Figure 6 shows that a sufficiently long delay allows for viable kinetic mixing in reach of all future searches.
To quantify the joint impact on the required hierarchy in $\beta$ and $\alpha_{D}$, combine the efficient resonance condition Eq. 48 with the backreaction bound Eq. 8:
$\frac{\beta}{\alpha_{D}/2\pi}$ $\gg 10^{10}\theta_{0}^{-1}\lambda^{-1/2}\sqrt{\frac{\Omega_{A}}{\Omega_{\mathrm{DM}}}}\frac{2m_{A}}{m_{a}}\left(\frac{m_{A}}{{\mathrm{\mu eV}}}\right)^{-5/2}\left(\frac{H_{\mathrm{osc}}}{10^{-22}\leavevmode\nobreak\ {\mathrm{eV}}}\right)^{5/4}.$ (49)
Equation 49 represents a significant reduction in the required hierarchy relative to the standard scenario, traded instead for a large hierarchy in $m_{a}/H_{\mathrm{osc}}$. The requirement is further reduced from the fiducial values in Eq. 49 by taking a heavier axion or dark photon, but at the expense of exacerbating the hierarchy in $m_{a}/H_{\mathrm{osc}}$. For instance, the fiducial hierarchy is order unity for dark photons with mass $10\leavevmode\nobreak\ \mathrm{meV}$, for which $m_{a}/H_{\mathrm{osc}}>10^{20}$. Delayed oscillations require a dynamical hierarchy rather than a parameter hierarchy, but candidate models may well require parameter hierarchies or tunings of their own.
There are numerous concrete models that delay the onset of axion oscillations [167, 168, 169, 170, 158], although mechanisms relying on attractive self-interactions [167, 168] result in axion fragmentation, spoiling the dark photon’s tachyonic resonance (see Sec. D.2). Repulsive self-interactions suppress axion perturbations, but constructing an axion potential with repulsive self-interactions (while possible [171]) is its own challenge because the periodic axion field always has attractive self interactions at large enough field excursions. Equation 49 tacitly assumes axion self-interactions do not disrupt dark photon production.
Example models whose dynamics can plausibly delay axion oscillations while also avoiding axion fragmentation may arise in axiverse scenarios [43]. Given a large number of axions logarithmically distributed in mass, there is a reasonable chance that a pair of axions have masses within an order-unity factor of each other [169]; candidate pairs indeed appear in concrete string theory compactifications [172]. Such “friendly” pairs can undergo nonlinear autoresonance, in which the axion with the smaller decay constant $f_{a}$ oscillates with fixed amplitude until a parametrically late time, after which it dilutes like matter [169, 170]. Importantly, autoresonance may occur whether axion self-interactions are attractive or repulsive [169]. The time between $H=m_{a}$ and when the axion starts diluting is controlled by the ratio of decay constants between the friendly axions, $\mathcal{F}\gg 1$, as $H_{\mathrm{osc}}\sim m_{a}\mathcal{F}^{-4/3}$. Quantified in decades, a hierarchy in $m_{a}/H_{\mathrm{osc}}$ translates to one $3/4$ths as severe in decay constants.
IV.2 Slow roll
Given that dark photon production from axion oscillations already requires a significant enhancement of $\beta$ beyond its natural expectation, one may as well consider enhancements yet larger than the bare minimum required for efficient resonance. We now point out the possibility that such large $\beta$ (much greater than, say, $50$) may yield a new dynamical regime in which dark photon production is delayed. Moreover, because such delayed production allows for larger $g_{D}$ without backreaction or defect formation, requiring larger $\beta$ itself does not necessarily exacerbate the hierarchy in $\beta/\alpha_{D}$, a claim we quantify below.
The cosmological evolution of massive, noninteracting scalars is relatively simple: the (homogeneous) field is frozen when Hubble friction dominates over the curvature of its potential ($H\gg m_{a}$) and subsequently undergoes underdamped oscillations once $H$ drops below $m_{a}$. The axion’s interaction with dark photons, however, yields a source term $\propto\beta\langle\mathbf{E}\cdot\mathbf{B}\rangle$ in its equation of motion Eq. 21a. The relative importance of dark photon backreaction may be parametrically estimated by comparing energy densities, i.e., comparing $\beta\rho_{A}$ and $\rho_{a}$. For the moderate values of $\beta$ considered in the standard scenario (Sec. II.3), tachyonic resonance of dark photons only becomes efficient in the oscillatory regime, and their backreaction onto the homogeneous axion only becomes important as energy transfer completes. However, for sufficiently large $\beta$, even the axion’s parametrically small velocity at early times,
$\dot{\bar{\theta}}$ $=-m_{a}\theta_{0}\left[m_{a}/5H+\mathcal{O}([m_{a}/H]^{3})\right],$ (50)
is enough to produce dark photons. As derived in Appendix C, if
$\theta_{0}\beta$ $\gg 48.5+3.2\ln\frac{f_{a}^{4}}{m_{a}^{4}\theta_{0}^{2}\beta^{6}},$ (51)
then not only does the slowly rolling axion efficiently produce dark photons, but the dark photons that are produced induce significant backreaction—not just because larger $\beta$ increases the magnitude of their backreaction at fixed energy density, but also because dark photons are produced at earlier times $m_{a}t_{\mathrm{BR}}\sim(\beta\theta_{0})^{-1/2}$. Rather than proceeding to oscillate when $H\leq m_{a}$, the axion may instead undergo an extended phase of slow roll.
Neglecting spatial axion perturbations (a strong assumption we comment on shortly), the structure of Eq. 21—where the dark photon’s backreaction grows with a rate proportional to the axion velocity—suggests the possibility of a quasi-equilibrium state where the rolling axion produces a dark photon background that exerts friction on the axion, in turn regulating the axion’s velocity. In particular, when $\dot{\bar{\theta}}=2m_{A}/\beta$, the axion rolls just fast enough that one dark photon polarization is massless. (The other is massive and may be ignored.) If $\dot{\bar{\theta}}$ were to roll any faster, the massless mode would become tachyonic, the energy in the rolling axion would be converted into dark photons, and $\theta$ would slow down. Under these assumptions, the totality of the axion energy density would gradually drain into the dark photon by the time $H_{\star}\approx m_{A}/\beta$, realizing both efficient and delayed dark photon production at once. Because $\dot{\bar{\theta}}$ is nearly constant, the axion rolls to its minimum almost entirely during the last $e$-fold of production; the majority of the dark photon energy density is then concentrated at wavelengths $k=m_{A}a_{\star}$.888 At extremely large $\beta$ and $m_{a}/m_{A}$, the axion velocity at the onset of slow roll is parametrically larger than its quasi-equilibrium value by an amount proportional to $\sqrt{\beta\theta_{0}}m_{a}/m_{A}$. It is therefore likely that much of the axion’s initial energy converts to highly relativistic dark photons before it slows to its quasi-equilibrium velocity; a large axion decay constant may thus be necessary to achieve the correct dark photon relic abundance.
In Appendix C, we show that the axion enters the overdamped regime just as its perturbations become comparable to its homogeneous component. If the axion indeed fully fragments, the tachyonic resonance is disrupted, meaning the slow-roll solution is not realized. On the other hand, this effective slow-roll regime is indeed realized in numerical simulations of preheating into massless gauge bosons after axion inflation [145, 146, 173, 148, 149, 151, 152, 153], indicating that axion perturbations may be regulated at large axion-photon coupling—in which case the axion would remain approximately homogeneous and overdamped as it rolls to the bottom of its potential. Suggestive as these results are, it remains to be shown whether they extend to the radiation dominated epoch relevant for dark photon dark matter production, a possibility only testable with nonlinear simulations. Given the magnitude of $\beta$ required, the initial resonance is quite broad and nonlinear interactions between modes should be strong (even if they don’t disrupt the homogeneous mode); robust simulations may require significant computational resources. We leave this investigation to future work and reiterate the uncertainty that any conclusions are subject to.
Slow roll, should it occur, requires larger couplings, but not necessarily in units of $\alpha_{D}$. That is, the increase in the maximum $g_{D}$ that evades backreaction onto the Higgs and defect formation can offset the required increase in $\beta$ itself. Taking $H_{\star}=m_{A}/\beta$ and substituting Eq. 8,
$\frac{\beta}{\alpha_{D}/2\pi}$ $\gtrsim 10^{34}\lambda^{-1/2}\sqrt{\frac{\Omega_{A}}{\Omega_{\mathrm{DM}}}}\left(\frac{m_{A}}{{\mathrm{\mu eV}}}\right)^{-1}\left(\frac{H_{\star}}{10^{-22}\leavevmode\nobreak\ {\mathrm{eV}}}\right)^{-1/4},$ (52)
which is in the same ballpark as the standard case [Eq. 25], though with slightly different mass dependence. However, axion–dark-photon couplings $\beta\gg 1$ may in fact require less model building than moderate $\beta\sim\mathcal{O}(10)\gg\alpha_{D}$. If the axion couples to PQ-charged fermions that have both dark electric and magnetic charges (dyons), then $\beta$ may be proportional to (the square of) the dark magnetic charge $4\pi/g_{D}$ rather than the electric charge $g_{D}$ [174]. As a result, $\beta$ would naturally be of order $1/\alpha_{D}$, realizing the slow-roll regime. Because $\beta$ is determined by the dark gauge coupling and itself sets the time of dark photon production $H_{\star}\sim m_{A}/\beta$, the magnetically coupled axion is potentially viable only in a narrow range of kinetic mixing bounded from above and below. Substituting $H_{\star}=\alpha_{D}m_{A}/2\pi$ into Eq. 8 yields
$g_{D}$ $\lesssim 2\times 10^{-8}\lambda^{1/7}\left(\frac{m_{A}}{{\mathrm{\mu eV}}}\right)^{5/14}\left(\frac{\Omega_{A}}{\Omega_{\mathrm{DM}}}\right)^{-1/7}.$ (53)
On the other hand, if the dark gauge coupling is too small, then (at a given $m_{A}$) $\beta$ would be so large that dark photon production is delayed until after we know the dark matter must exist. This places a lower bound on the gauge coupling of
$g_{D}$ $\gtrsim 9\times 10^{-8}\left(\frac{m_{A}}{{\mathrm{\mu eV}}}\right)^{-1/2}\left(\frac{H_{\star,\mathrm{min}}}{10^{-22}\leavevmode\nobreak\ {\mathrm{eV}}}\right)^{1/2}.$ (54)
We plot the resulting parameter space in Fig. 6.
Figure 6: Kinetic mixing parameter space of dark photon dark matter produced from rolling axions. Red lines depict the maximum kinetic mixing [Eq. 9] that evades backreaction onto the Higgs when either the axial coupling $\beta$ is large enough to realize slow roll (Sec. IV.2) or when axion oscillations are delayed (relative to $m_{A}\lesssim m_{a}$) sufficiently long by other means (Sec. IV.1). The blue region depicts the parameter space available when the axion coupling is $\beta\propto 1/\alpha_{D}$, as may be for an axion coupled in the magnetic basis. We note that the slow-roll scenario is subject to theoretical uncertainty regarding nonlinear dynamics and, for the magnetic axion, the consistency of large magnetic couplings (both discussed in Sec. IV.2). The red region depicts parameter space for which backreaction onto the Higgs occurs when scales relevant to CMB anisotropies enter the horizon, i.e., when the dark photon must behave as dark matter; current bounds and projected search space are depicted in gray and yellow as labelled in Fig. 3.
Intriguingly, production from a magnetically coupled axion may only both evade defect formation and occur early enough for masses $m_{A}\gtrsim\mu{\mathrm{eV}}$. Moreover, the available parameter space is already largely excluded, but numerous future axion haloscopes, such as ALPHA [109], MADMAX [110], Dark E-field [106], BREAD [111], LAMPOST [112], and SuperCDMS [82, 113] could easily detect or exclude the parameter space that remains below an ${\mathrm{eV}}$. LZ [114] could also potentially detect viable parameter space around $10\leavevmode\nobreak\ \mathrm{keV}$.
The viability of a magnetic-basis coupling as a means to boost the axion-photon coupling remains unsettled at present [165, 166]. Although most of the objections raised by Ref. [165] are specific to the case of the QCD axion coupled to the Standard Model photon, questions remain even for the case of axionlike particles coupled to a dark photon. Because of the dual Witten effect [175], as the axion rolls from 00 to $2\pi$, the magnetic charge of every dyon changes by one unit, implying the existence of an infinite dyonic tower. Not only must the magnetically charged dyons be very heavy to avoid magnetic Schwinger pair production [176, 177], but Ref. [165] also suggests that integrating out such a tower generically drives the theory to weak coupling. The coupling to heavy dyons would also generate a large mass for the axion, though because $H_{\star}$ is independent of $m_{a}$ once the axion has achieved its equilibrium, slow-roll velocity, production may still proceed as described even if the axion were quite heavy. The dark Higgs is one such electrically charged state that would form the basis of such a dyonic tower [178]. The possibility of axions strongly coupled to dark photons—perhaps with a Stückelberg rather than a Higgs mass—remains an intriguing possibility, and further investigation is required to establish its viability. Finally, note that even if one ignores possible backreaction onto the radial mode associated with the Stückelberg mass, the lower bound [Eq. 54] on $\varepsilon$ remains.
V Dark photon dark matter in scalar–Abelian-Higgs theories
Unlike axions, whose couplings are restricted to those that perturbatively preserve their continuous shift symmetry, scalars can couple to any operator and therefore offer more possible means to produce dark photons and evade strong backreaction. In particular, a scalar field effectively modulates the fundamental parameters of a theory throughout spacetime [179, 180]. A homogeneous and dynamical scalar field coupled to the Abelian-Higgs model may therefore modulate the backreaction threshold [Eq. 5] and the string production threshold [Eq. 12] over time. For example, the dark Higgs VEV could be larger when dark photons are first produced than it is today, raising the backreaction and defect formation thresholds at early times when the dark photon energy density is largest.
We begin this section by introducing a useful parametrization of a scalar’s interactions with the Abelian-Higgs Lagrangian in Sec. V.1. In Sec. V.2, we illustrate how “universal” scalar couplings can alleviate defect bounds in existing scenarios, focusing on gravitational production during inflation; in Sec. V.2.2, we show that this scenario can be realized in well-motivated physics models such as the string dilaton [53]. Scalar couplings also offer unique mechanisms for particle production: in Secs. V.3 and V.4, we explore two models that avoid backreaction and defect formation through the same physics that drives dark photon production [52, 51].
V.1 Equations of motion
The classical Lagrangian describing an Abelian-Higgs theory coupled to a scalar $\phi$ may be parametrized as999 Equation 55 neglects a possible fourth coupling that modulates the quadratic term of the dark Higgs potential independently of the quartic. However, our choice of parametrization is sufficient to modulate all the Abelian-Higgs parameters independently and also includes all possible (two-)derivative interactions.
$\mathcal{L}_{\mathrm{SAH}}$ $=\frac{1}{2}\partial_{\mu}\phi\partial^{\mu}\phi-V(\phi)-\frac{W(\phi)}{4}F_{\mu\nu}F^{\mu\nu}+\frac{X(\phi)}{2}D_{\mu}\Phi\left(D^{\mu}\Phi\right)^{\ast}-\frac{\lambda Y(\phi)}{4}\left({\left|\Phi\right|}^{2}-v^{2}\right)^{2},$ (55)
generalizing Eq. 1. As discussed in Appendix A, written in terms of radial fluctuations $h$ of the Higgs about its VEV in unitary gauge ($\Pi=0$), Eq. 55 expands to
$\begin{split}\mathcal{L}_{\mathrm{SAH}}&=\frac{1}{2}\partial_{\mu}\phi\partial^{\mu}\phi-V(\phi)-\frac{W(\phi)}{4}F_{\mu\nu}F^{\mu\nu}+\frac{X(\phi)}{2}\left(1+\frac{h}{v}\right)^{2}m_{A}^{2}A_{\mu}A^{\mu}\\ &\hphantom{{}={}}+\frac{X(\phi)}{2}\partial_{\mu}h\partial^{\mu}h-Y(\phi)\left(\frac{\lambda}{4}h^{4}+\lambda vh^{3}+\frac{1}{2}m_{h}^{2}h^{2}\right).\end{split}$ (56)
The coupling functions $W$, $X$, and $Y$ could in general depend on multiple scalar fields; for simplicity we consider only a single scalar field, but the extension to multiple is straightforward.
To rewrite the theory with canonically normalized kinetic terms, we rescale the vector and Higgs fields as $A_{\mu}=\mathcal{A}_{\mu}/\sqrt{W(\phi)}$ and $h=\mathfrak{h}/\sqrt{X(\phi)}$. Then, suppressing the explicit $\phi$ dependence of $W$, $X$, and $Y$ for brevity,
$\begin{split}\mathcal{L}_{\mathrm{SAH}}&=\frac{1}{2}\partial_{\mu}\phi\partial^{\mu}\phi-V(\phi)-\frac{1}{4}\left(\mathcal{F}_{\mu\nu}+\mathcal{G}_{\mu\nu}\right)\left(\mathcal{F}^{\mu\nu}+\mathcal{G}^{\mu\nu}\right)+\frac{1}{2}\frac{Xm_{A}^{2}}{W}\left(1+\frac{\mathfrak{h}}{\sqrt{X}v}\right)^{2}\mathcal{A}_{\mu}\mathcal{A}^{\mu}\\ &\hphantom{{}={}}+\frac{1}{2}\partial_{\mu}\mathfrak{h}\partial^{\mu}\mathfrak{h}-\frac{1}{2}\mathfrak{h}\partial^{\mu}\mathfrak{h}\frac{\partial_{\mu}X}{X}+\frac{1}{8}\mathfrak{h}^{2}\frac{\partial_{\mu}X\partial^{\mu}X}{X^{2}}-\frac{\lambda Y}{X^{2}}\left(\frac{\mathfrak{h}^{4}}{4}+v\sqrt{X}\mathfrak{h}^{3}\right)-\frac{1}{2}\frac{Y}{X}m_{h}^{2}\mathfrak{h}^{2},\end{split}$ (57)
where
$\mathcal{F}_{\mu\nu}$ $\equiv\partial_{\mu}\mathcal{A}_{\nu}-\partial_{\nu}\mathcal{A}_{\mu}$ (58a)
and
$\mathcal{G}_{\mu\nu}$ $\equiv\frac{1}{2W}\left(\mathcal{A}_{\mu}\partial_{\nu}W-\mathcal{A}_{\nu}\partial_{\mu}W\right).$ (58b)
To treat scenarios where a homogeneous, classical background scalar field effectively modulates the fundamental parameters of the Abelian-Higgs model through its cosmological evolution, we expand $\phi(t,\mathbf{x})=\bar{\phi}(t)+\delta\phi(t,\mathbf{x})$ where $\delta\phi$ is a small perturbation. The form of Eq. 57 therefore motivates absorbing the $\bar{\phi}$ dependence of the theory into its fundamental parameters as
$m_{A}(\bar{\phi})$ $\equiv m_{A}\sqrt{X(\bar{\phi})/W(\bar{\phi})},$ (59a)
$v(\bar{\phi})$ $\equiv v\sqrt{X(\bar{\phi})},$ (59b)
$\lambda(\bar{\phi})$ $\equiv\lambda Y(\bar{\phi})/X(\bar{\phi})^{2},$ (59c)
and, as a consequence,
$m_{h}(\bar{\phi})$ $=\sqrt{2\lambda(\bar{\phi})}v(\bar{\phi})=m_{h}\sqrt{Y(\bar{\phi})/X(\bar{\phi})}$ (59d)
and
$g_{D}(\bar{\phi})$ $=m_{A}(\bar{\phi})/v(\bar{\phi})=g_{D}/\sqrt{W(\bar{\phi})}.$ (59e)
From this point on, we take the dark photon to have no homogeneous component and set the Higgs to zero for simplicity (with results including the Higgs presented in Appendix A). In an FLRW spacetime [Eq. 138], the linearized equations of motion in Fourier space read
00 0 $=\ddot{\bar{\phi}}+3H\dot{\bar{\phi}}+V^{\prime}(\bar{\phi})$ (60a)
00 0 $=\ddot{\mathcal{A}}_{\pm}+H\dot{\mathcal{A}}_{\pm}+\left[\frac{k^{2}}{a^{2}}+m_{A}(\bar{\phi})^{2}-\frac{H}{2}\frac{\partial_{t}W(\bar{\phi})}{W(\bar{\phi})}-\frac{\partial_{t}^{2}\sqrt{W(\bar{\phi})}}{\sqrt{W(\bar{\phi})}}\right]\mathcal{A}_{\pm}$ (60b)
00 0 $=\ddot{A}_{\parallel}+\frac{\left[3H+\partial_{t}\ln X(\bar{\phi})\right]k^{2}+\left[H+\partial_{t}\ln W(\bar{\phi})\right]a^{2}m_{A}(\bar{\phi})^{2}}{k^{2}+a^{2}m_{A}(\bar{\phi})^{2}}\dot{A}_{\parallel}+\left[\frac{k^{2}}{a^{2}}+m_{A}(\bar{\phi})^{2}\right]A_{\parallel}.$ (60c)
Equation 60 also decomposed the vector onto a basis of transverse ($\mathcal{A}_{\pm}$) and longitudinal ($A_{\parallel}$) polarization; see Appendix A for full details. Note that the equation of motion for the longitudinal component Eq. 60c is written in terms of the noncanonical field, as the rescaling applied for transverse modes does not fully remove scalar couplings from the longitudinal mode’s friction term. Equation 60c also more transparently demonstrates that the relativistic and nonrelativistic limits of the longitudinal dynamics coincide with those of scalars and of transverse vectors, respectively, in accordance with the Goldstone equivalence theorem.
The structure of Eq. 60 is quite general, allowing for a number of qualitatively distinct dark photon production mechanisms. Large field excursions (relative to whatever relevant energy scale), for instance, allow for large variation in the Abelian-Higgs parameters [Eq. 59], in which case both the backreaction threshold and the kinematic conditions for efficient particle production become time dependent. We specialize to two examples that illustrate each of these possibilities. In Sec. V.2 we show that a universally coupled dilaton broadens the parameter space available to inflationary production by modulating the Higgs VEV, simultaneously weakening the defect bound and making production more efficient. In Sec. V.3 we consider the model presented in Ref. [52] of a scalar coupled to the dark photon kinetic term. Such a coupling can both exponentially suppress the dark photon mass at early times and induce a tachyonic instability that achieves extremely delayed production. Finally, in Sec. V.4 we discuss an alternative particle production mechanism that was pointed out by Ref. [51]: for a particular (and finely tuned) mass ratio between the dark photon and an oscillating scalar, parametric resonance can efficiently produce dark photons even with arbitrarily slow resonance, providing another realization of delayed production. This scenario, in contrast to the aforementioned, only depends on the leading-order behavior of the scalar’s couplings about the minimum of its potential. Throughout, we also discuss various unique signatures predicted by each model and their relation to the dark photon’s direct detection parameter space.
V.2 Universal coupling
In this section, we show that defect constraints on inflationary production are weakened if particle masses (possibly including SM particles) are $\phi$ dependent. A scalar’s interactions with matter, be it the dark matter or SM matter, generate for it an effective potential that naturally drives the field to values that minimize the free energy of the system. Absent a bare potential, the scalar $\phi$ rolls to values that either minimize particle masses (for $\phi$-dependent masses) or decouple the theory (for $\phi$-dependent couplings). For the dark Abelian-Higgs theories we consider, the latter possibility would make $g_{D}$, which already must be small to avoid defect formation, decrease yet further as the Universe evolves towards its ground state. On the other hand, a larger dark Higgs VEV during inflation works in favor of weaker defect bounds: defect formation is suppressed when all degrees of freedom are heavier, while cosmological evolution makes them lighter at late times.
The particular choice of coupling functions that endows only the theory’s mass scales with $\phi$ dependence is $W(\phi)=1$ and
$Y(\phi)$ $=X(\phi)^{2}\equiv Z(\phi)^{2},$ (61)
for which, using Eq. 59, the Abelian-Higgs parameters all scale with $Z^{d/2}$ where $d$ is the mass dimension of the particular parameter. In Sec. V.2.1 we outline the model-independent features that weaken defect bounds from inflationary production, and in Sec. V.2.2 we discuss one particularly well-motivated realization from string theory, in which all parameters in the theory depend on the so-called dilaton [53]. While we focus on inflationary production because of its minimalism, the mechanism of Sec. V.2.2 could suppress backreaction onto the Higgs for a number of other production mechanisms discussed in this paper.
V.2.1 Inflationary production and varying masses
To study the impact of evolving mass scales on defect formation and the dark photon relic abundance on a model-agnostic basis, we first work in terms of a time-dependent mass $m_{A}(a)$. The energy density of the longitudinal mode [Eq. 148] scales as
$\frac{\mathrm{d}{\rho_{\parallel}}}{\mathrm{d}{\ln k}}$ $\sim\frac{1}{2a^{2}}m_{A}(a)^{2}\frac{\mathrm{d}{|A_{\parallel}|^{2}}}{\mathrm{d}{\ln k}}$ (62)
in both the relativistic and nonrelativistic limits. When $m_{A}(a)$ evolves slowly compared to $H$, the longitudinal mode equation of motion [Eq. 60c] simplifies to
00 0 $=\ddot{A}_{\parallel}+\frac{3k^{2}+a^{2}m_{A}(a)^{2}}{k^{2}+a^{2}m_{A}(a)^{2}}H\dot{A}_{\parallel}+\left(\frac{k^{2}}{a^{2}}+m_{A}(a)^{2}\right)A_{\parallel}.$ (63)
In the relativistic limit $k\gg am_{A}(a)$, $A_{\parallel}$’s dynamics are independent of $m_{A}(a)$, but once $A_{\parallel}$ is nonrelativistic its dynamics depend nontrivially on the evolution of its mass. If the dark photon’s mass evolves adiabatically (such that a Wentzel-Kramers-Brillouin [WKB] solution is valid) after it becomes nonrelativistic, its amplitude then scales $\propto 1/\sqrt{m_{A}(a)}$. This distinction between the relativistic and nonrelativistic regimes is crucial to determining the dynamics of $m_{A}(a)$ that alleviate defect formation bounds.
Equation 10 shows that the mode that dominates the abundance of dark photons produced gravitationally during inflation is that which enters the horizon at the same time it becomes nonrelativistic—that is, the mode with wave number $k_{\star}=a_{\star}H(a_{\star})=a_{\star}m_{A}(a_{\star})$. Assume that the peak remains at this mode, evaluated when $H(a)$ first drops below $m_{A}(a)$.101010 In principle the mass could subsequently decrease rapidly enough that this mode becomes temporarily relativistic and no longer dominates the energy spectrum. We neglect this possibility for simplicity, since it is avoided if $m_{A}$’s evolution either is sufficiently slow or begins sufficiently late. At $a>a_{\star}$ its energy density scales as
$\frac{\mathrm{d}{\rho_{\parallel}(k_{\star})}}{\mathrm{d}{\ln k}}$ $\sim\frac{m_{A}(a)^{2}}{2}\left(\frac{k_{\star}H_{I}}{2\pi a_{\star}{m_{A}}_{I}}\right)^{2}\left(\frac{a}{a_{\star}}\right)^{-3}\frac{m_{A}(a_{\star})}{m_{A}(a)}.$ (64)
These three factors respectively encode the longitudinal modes’ (current) mass; its initial conditions from inflation [50], during which its mass takes the value ${m_{A}}_{I}$; its redshifting upon re-entering the horizon (when it simultaneously becomes nonrelativistic); and the adiabatic evolution induced by its evolving mass. Assuming this mode enters the horizon during the radiation era and setting $k_{\star}=a_{\star}m_{A}(a_{\star})$,
$\frac{\mathrm{d}{\rho_{\parallel}(k_{\star})}}{\mathrm{d}{\ln k}}$ $\sim\frac{m_{A}(a)^{2}}{2}\left(\frac{m_{A}(a_{\star})}{{m_{A}}_{I}}\right)^{2}\left(\frac{H_{I}}{2\pi}\right)^{2}\left(\frac{H}{m_{A}(a_{\star})}\right)^{3/2}\frac{m_{A}(a_{\star})}{m_{A}(a)}.$ (65)
If $m_{A}$ only changes while the longitudinal mode is still relativistic, then the WKB factor in Eq. 65 reduces to unity and the abundance scales with the square of its late- and early-time masses:
$\frac{\mathrm{d}{\rho_{\parallel}(k_{\star})}}{\mathrm{d}{\ln k}}$ $\sim\frac{m_{A}^{2}}{2}\left(\frac{H_{I}}{2\pi}\right)^{2}\left(\frac{H}{m_{A}}\right)^{3/2}\cdot\left(\frac{m_{A}}{{m_{A}}_{I}}\right)^{2}.$ (66)
Compensating for the $(m_{A}/{m_{A}}_{I})^{2}$ suppression of the dark photon abundance (at a given present-day mass) requires increasing $H_{I}\propto{m_{A}}_{I}$, which nullifies any relaxation of defect constraints [Eq. 12], as noted by Ref. [48].111111 As mentioned in Sec. II.1.2, Ref. [48] (in Appendix C.2) considers a nonminimal coupling of the dark Higgs, $\xi R|\Phi|^{2}$, as a means to drive the Higgs VEV to larger values during inflation, which then returns to its vacuum value shortly afterward. The evolution of the mass scales in the theory thus necessarily terminates long before the dominant dark photon mode becomes nonrelativistic.
Now consider the scenario where $m_{A}$ instead remains at its inflationary value ${m_{A}}_{I}$ until after the dominant dark photon mode becomes nonrelativistic. In this case, the second factor in Eq. 65 (which encodes the superhorizon initial condition seeded during inflation) simplifies to unity. As a result, $m_{A}$’s evolution only affects the dark photon mode amplitude through adiabatic evolution and an earlier onset of subhorizon redshifting:
$\frac{\mathrm{d}{\rho_{\parallel}(k_{\star})}}{\mathrm{d}{\ln k}}$ $\sim\frac{m_{A}^{2}}{2}\left(\frac{H_{I}}{2\pi}\right)^{2}\left(\frac{H}{m_{A}}\right)^{3/2}\cdot\left(\frac{m_{A}}{{m_{A}}_{I}}\right)^{1/2}.$ (67)
Therefore, $H_{I}$ need only be $({m_{A}}_{I}/m_{A})^{1/4}$ times larger to maintain the correct relic abundance, while the defect bound on $H_{I}$ is weakened by ${m_{A}}_{I}/m_{A}$. Matching $\rho_{\parallel}$ to the dark matter relic density requires
$\frac{{m_{A}}_{I}}{m_{A}}$ $=\frac{m_{A}}{0.1{\mathrm{meV}}}\left(\frac{H_{I}}{5\times 10^{13}\leavevmode\nobreak\ {\mathrm{GeV}}}\right)^{4},$ (68)
while the defect bound limits
$g_{D}\leq\frac{{m_{A}}_{I}}{H_{I}}$ $\lesssim 2\times 10^{-19}\left(\frac{m_{A}}{{\mathrm{eV}}}\right)^{2}\left(\frac{H_{I}}{5\times 10^{13}\leavevmode\nobreak\ {\mathrm{GeV}}}\right)^{3}.$ (69)
The requisite early-late mass ratio is
$\frac{{m_{A}}_{I}}{m_{A}}$ $=10^{4}\left(\frac{g_{D}}{2\times 10^{-19}}\right)^{1/2}\left(\frac{H_{I}}{5\times 10^{13}\leavevmode\nobreak\ {\mathrm{GeV}}}\right)^{5/2}=10^{4}\left(\frac{g_{D}}{2\times 10^{-19}}\right)^{4/3}\left(\frac{m_{A}}{{\mathrm{eV}}}\right)^{-5/3}.$ (70)
Equation 69 is parametrically weaker than in the minimal scenario [Eq. 12], and is in fact weakest for the highest allowed inflationary scale because the relic abundance is more sensitive to $H_{I}$ than to the drop in the dark photon’s mass. Further, the bound scales with a higher power of $m_{A}$ than Eq. 69 because the relic abundance no longer requires a unique inflationary scale as a function of $m_{A}$; heavier (present-day) dark photons may therefore be accommodated at larger $H_{I}$ when simultaneously increasing ${m_{A}}_{I}/m_{A}$.
V.2.2 Damour-Polyakov dilaton
We now show that the model of the string dilaton proposed by Damour and Polyakov [53] can mechanize cosmological evolution of the dark photon mass that weakens defect bounds, as described phenomenologically in the previous section. We briefly review the mechanism of Ref. [53]. In general, the string dilaton $\phi$ is expected to couple to all Standard Model operators, essentially promoting the constants of nature to field-dependent quantities $\zeta\to\zeta(\varphi)$ (where $\varphi\equiv\phi/\sqrt{2}M_{\mathrm{pl}}$). If the “constant” has sufficiently large linear dependence on the dilaton, then it mediates a force between objects whose mass depends on the value of $\zeta$; such fifth forces are strongly constrained by tests of equivalence principle violation (unless $\phi$ is sufficiently heavy) [181, 182, 183]. If the dilaton only enters the action via the single function $\zeta(\phi)$ or by a collection of functions all with the same extrema (and has a canonical kinetic term), then one may shift its value such that it has no linear couplings, i.e., $\zeta(\varphi)=\zeta_{0}+\kappa\varphi^{2}/2+\mathcal{O}(\varphi^{3})$.
The now-leading quadratic coupling still mediates a fifth force, but one proportional to $\varphi$’s expectation value. Reference [53] identified a mechanism that efficiently drives $\phi$ to its minimum: if the mass $m_{X}$ of a particle species $X$ is generated by the confinement of a strongly coupled gauge group whose gauge coupling depends on $\phi$, then $m_{X}$ depends strongly on the dilaton:
$m_{X}(\phi)$ $=m_{X}\exp\left[\kappa\ln\left(\frac{e^{1/2}\hat{\Lambda}_{s}}{m_{X}}\right)\frac{\varphi^{2}}{2}+\mathcal{O}\left(\varphi^{3}\right)\right],$ (71)
where $m_{X}$ is the particle mass when $\phi$ is at the minimum of its effective potential, $\hat{\Lambda}_{s}$ is the string scale (which is about $3\times 10^{17}\leavevmode\nobreak\ {\mathrm{GeV}}$), and $\kappa$ is an $\mathcal{O}(1)$ coefficient of the Taylor expansion of the gauge coupling of the confining sector with respect to $\phi$. Because $\hat{\Lambda}_{s}\gg m_{X}$, the constant of proportionality
$\beta_{X}\equiv\kappa\ln(e^{1/2}\hat{\Lambda}_{s}/m_{X})\approx\kappa\left(40.75-\ln\frac{m_{X}}{{\mathrm{GeV}}}\right)$ (72)
can be quite large, making this attractor mechanism efficient (as required for the dilaton to evade fifth force bounds). The confining gauge group coupling changes by $\mathcal{O}(1)$ over a field range $\Delta\phi\sim M_{\mathrm{pl}}/\sqrt{\kappa}$, ensuring perturbative control even as $m_{X}(\phi)$ [and, for our purposes, $m_{A}(\phi)$] change by multiple orders of magnitude. If indeed the dilaton modulates the dark photon mass in this way, then per Sec. V.2.1 these dynamics are sufficient to alleviate inflationary defect formation bounds.
The dilaton’s cosmological dynamics are governed by its interactions with the SM plasma, which induce an effective potential that takes the form [184, 185, 155]
$V(\phi)$$\approx\sum_{\{X|m_{X}(\phi)\lesssim T\}}\frac{g_{X}}{24}T^{2}m_{X}(\phi)^{2}\cdot\begin{cases}1,&\text{boson},\\ 1/2,&\text{fermion}\end{cases}$ (73)
at one loop order, where $g_{X}$ is the number of degrees of freedom associated with species $X$ and the sum runs over all particles whose masses are below the plasma temperature and depend on $\phi$ via the confining gauge group. The homogeneous component of $\phi$ evolves according to
00 0 $=\ddot{\bar{\phi}}+\frac{3}{2t}\dot{\bar{\phi}}+\sum_{m_{X}(\bar{\phi})\lesssim T}\frac{g_{X}}{24}T^{2}m_{X}^{2}\beta_{X}e^{\beta_{X}\bar{\phi}^{2}/2M_{\mathrm{pl}}^{2}}\frac{\bar{\phi}}{M_{\mathrm{pl}}^{2}},$ (74)
which may be integrated directly in the slow-roll approximation, $\ddot{\bar{\phi}}\to 0$. The heaviest coupled species that hasn’t annihilated away dominates the effective potential, meaning the dilaton starts rolling from its initial condition around when the heaviest species begins to annihilate. The temperature at that time is
$T_{\mathrm{roll}}$ $\approx\sum_{m_{X}(\phi_{0})\lesssim T_{\mathrm{roll}}}\sqrt{\frac{g_{X}}{g_{\star}}}m_{X}(\phi_{0})\frac{\beta_{X}\phi_{0}}{M_{\mathrm{pl}}}.$ (75)
After this point, particle masses drop in proportion to the temperature,
$m_{X}(\phi(t))$ $\approx m_{X}(\phi_{0})\frac{T(t)}{T_{\mathrm{roll}}}.$ (76)
No particle falls out of thermal equilibrium as long as its mass continues to scale in proportion to its temperature, effectively pausing the thermal evolution of that species (or of the Universe if all masses are $\phi$ dependent) until the dilaton reaches the minimum of its effective potential,121212 The energy density of the dilaton during its subsequent oscillations about its minimum dilute faster than matter because its mass drops with $T$, ensuring it never contributes significantly to the dark matter density. at which point all masses take on their present-day values. These dynamics rely only on the slow-roll approximation and not the particular form of the $\phi$-dependence of $m_{X}$—only that $m_{X}$ is a sufficiently steep function of $\phi$ and that the scalar’s bare potential is negligible. The same phenomenon occurs in, for example, Brans-Dicke [186]–inspired models (see e.g., Ref. [187]).
To have any impact on the available parameter space of dark photon dark matter, both $m_{A}$ and at least one thermalized species must acquire a $\phi$-dependent mass in the manner described above (though in principle all masses could depend on $\phi$). Defect formation limits are thus alleviated by whatever amount the dark photon mass drops after $H={m_{A}}_{I}$, per Sec. V.2.1; however, the dilaton can in principle begin rolling earlier, which would not contribute any relaxation of defect limits but would still require a larger $H_{I}$ to set the correct relic abundance. In order for the dark photon mass to start shrinking only after the dominant mode has become nonrelativistic, the dilaton must start rolling at a Hubble rate $H_{\mathrm{roll}}\propto T_{\mathrm{roll}}^{2}/M_{\mathrm{pl}}$ below ${m_{A}}_{I}$. Using Eq. 75, the relic abundance constraint Eq. 68, and $m_{A,I}/m_{A}=\exp\left(\beta_{X}\phi_{0}^{2}/4M_{\mathrm{pl}}^{2}\right)$:
$\frac{H_{\mathrm{roll}}}{{m_{A}}_{I}}\approx 0.025\left(\frac{\beta_{X}}{36}\right)\left(\frac{m_{X}}{125{\mathrm{GeV}}}\right)^{2}\left(\frac{H_{I}}{5\times 10^{13}\leavevmode\nobreak\ {\mathrm{GeV}}}\right)^{4}\ln\left[\left(\frac{m_{A}}{0.1{\mathrm{meV}}}\right)\left(\frac{H_{I}}{5\times 10^{13}\leavevmode\nobreak\ {\mathrm{GeV}}}\right)^{4}\right].$ (77)
The expanded parameter space available to “dilaton-assisted” inflationary production is depicted in Fig. 7 and could be probed by LZ [114]. This accessible parameter space would require inflationary scales so large that primordial gravitational waves would be easily detected by future CMB observations [188, 189, 190] (but are not already excluded).
Figure 7: Kinetic mixing parameter space [Eq. 9] of dark photon dark matter produced during inflation. The red region depicts parameter space that is constrained by defect formation for dilaton-assisted inflationary production for the maximal inflationary scale presently allowed (corresponding to a tensor-to-scalar ratio $r=0.036$ [116]), and the upper limit for $r=10^{-3}$ (a value that can be probed by future CMB observations [188, 189, 190]) appears as labeled. The upper limit for standard inflationary production (without any mass evolution) is also plotted in red. Limits and projects are displayed as in Fig. 4.
However, there are currently no direct detection prospects at masses below a keV. Finally, as described in Ref. [53], this model predicts violations of the weak equivalence principle, which are stronger for $\kappa\lesssim 1$ and potentially observable in successors to Eöt-Wash [181, 182] or MICROSCOPE [191, 183].
V.3 Rapidly varying dark gauge coupling
Whether dark photons are produced gravitationally during inflation or via an oscillating axion afterward, the dark sector is populated at the end of inflation (if not earlier). What distinguishes the two scenarios is that gravitationally produced dark photons redshift even during inflation, whereas in the latter case the dark sector density remains frozen in the misaligned axion condensate until $H=m_{a}$. In this way, the axion effectively acts as a reservoir, allowing for a lower maximum dark photon density and reducing backreaction onto the dark Higgs once dark photons are produced. Kinematic efficiency ultimately limits how long dark photon production from axions may be delayed (Sec. II.3), requiring larger couplings or field excursions ($\beta\theta_{0}\gg 1$) and typically also a boost of the axion–dark-photon coupling by ad hoc means (but see Sec. IV.2 for a possible exception). In this section, we show that a scalar coupled to the kinetic term of the dark photon can realize further delayed dark photon production by eliminating the kinematic barrier (Sec. V.3.1). We present the dynamics in a concrete scenario (Sec. V.3.2), and we find that such scenarios do not require parametrically large dimensionless numbers, in contrast to axion production (Sec. V.3.3); however, we do not identify a concrete UV completion. We then discuss unique signatures of delayed dark photon production in Sec. V.3.4.
V.3.1 Resonant production with kinetic couplings
Whereas dilaton-assisted inflationary production (Sec. V.2) relied on couplings that modulate mass scales alone, the kinetic couplings $W$ and $X$ (to the dark photon and Higgs, respectively) modulate the dark photon mass without changing the backreaction threshold $\lambda v^{4}\propto Y$ [Eq. 59]. Thus, instead of raising the threshold for defect formation relative to the required dark photon energy density, kinetic couplings could suppress the dark photon mass, eliminating the kinematic barrier to resonance and delaying production to late times $H\sim m_{\phi}\ll m_{A}$. Moreover, kinetic couplings are derivative couplings, and the requisite evolution of $X$ or $Y$ by orders of magnitude (between their early-time and present-day values) can mechanize tachyonic resonance that could itself source the dark photon relic abundance.
The scalar-dependent dark photon mass $m_{A}(\bar{\phi})=m_{A}\sqrt{X(\bar{\phi})/W(\bar{\phi})}$ is small compared to its present-day value if either $X$ is initially very small or $W$ very large. (We normalize the values of $X$ and $Y$ today to unity without loss of generality). The massless limit of Eq. 60c is
00 0 $=\ddot{A}_{\parallel}+\partial_{t}\ln X(\bar{\phi})\dot{A}_{\parallel}+k^{2}A_{\parallel},$ (78)
neglecting cosmic expansion for illustration’s sake. Exponential growth requires a negative friction term, but the longitudinal mode’s friction term is positive when $X$ increases with time. The coupling $W$, on the other hand, does decrease in our hypothesized scenario. Setting $X=Y=1$, the canonical, transverse equation of motion Eq. 60b takes the simplified form
00 0 $=\ddot{\mathcal{A}}_{\pm}+H\dot{\mathcal{A}}_{\pm}+\omega_{\pm}(k)^{2}\mathcal{A}_{\pm},$ (79)
where the dark photon’s effective frequency squared is
$\omega_{\pm}(k)^{2}$ $=\frac{k^{2}}{a^{2}}+\frac{m_{A}^{2}}{\bar{W}}-\frac{H}{2}\frac{\dot{\bar{W}}}{\bar{W}}-\frac{\partial_{t}^{2}\sqrt{\bar{W}}}{\sqrt{\bar{W}}}.$ (80)
Transverse modes thus have a negative squared effective mass if $\dot{\bar{W}}^{2}<2\ddot{\bar{W}}\bar{W}$; functions that are both large in magnitude and accelerating yield the fastest growth. If the rate of change of $W$ is proportional to the scalar’s effective mass $m_{\phi}$, i.e., $\dot{\bar{W}}/\bar{W}=\dot{\bar{\phi}}\partial_{\phi}\ln W\sim m_{\phi}\partial\ln W/\partial\ln\phi$, the tachyonic growth condition $\omega_{\pm}^{2}\leq 0$ requires the kinetic function to be larger than the squared dark-photon–scalar mass ratio, $W(\phi_{0})\gtrsim(m_{A}/m_{\phi})^{2}$. This, however, is an underestimate: efficient production requires that $\omega_{\pm}^{2}$ is negative for a period long enough to build up all of the dark matter. Therefore, $W$ must decrease for a parametrically long time and so by many orders of magnitude. We next present a concrete realization of tachyonic production via a rapidly varying dark gauge coupling, i.e., a suitable choice of coupling function $W$ and scalar potential $V$ as first presented in Ref. [52].
V.3.2 Dynamics
Realizing functions $W$ that vary by orders of magnitude generally requires exponential dependence on the scalar and that the scalar rolls over a fairly large distance (in appropriate units). We therefore choose a potential that is compatible with large field excursions: the “runaway,” exponential potential
$V(\phi)$ $=M^{2}f^{2}e^{-\phi/f}\equiv m_{\phi}^{2}f^{2}e^{-\left(\phi-\phi_{0}\right)/f},$ (81)
which is motivated by certain constructions in string theory [192, 193, 194, 195] and whose self-similarity stabilizes it against quantum corrections from self-interactions [196]. The latter equality in Eq. 81 merely defines $m_{\phi}$ to coincide with the scalar’s effective mass when at its homogeneous initial condition $\phi=\phi_{0}$ (whereas $M$ is that when $\phi$ crosses zero). The evolution of a homogeneous scalar in Eq. 81 is approximately
$\bar{\phi}(t)$ $\approx\phi_{0}+f\ln\left[1+(m_{\phi}t)^{2}\right],$ (82)
up to oscillations that fall off with $(m_{\phi}t)^{-1/4}$.131313 These oscillations are logarithmic in time and change the resulting dark photon mode amplitude at the $\mathcal{O}(1)$ level. Since $\mathcal{O}(1)$ changes to the mode amplitude may be compensated by similar changes to parameter values, we neglect this subtlety for the sake of clarity.
Many possible coupling functions satisfy the tachyonic resonance condition $\partial_{t}^{2}\sqrt{W}/\sqrt{W}<0$, but an exponential does so even as $\phi$’s time derivatives become small [as for Eq. 82] and for all field values of $\phi$. However, were $W$ a pure exponential, nothing would prevent the dark photon’s mass from continuing to decrease to the present day. To stabilize the vector’s mass, one could augment the scalar potential with a minimum at some finite $\phi$ or modify the coupling function $W$ to asymptote to a constant at late times. We choose the latter option (for reasons explained at the end of this section), taking
$W(\phi)$ $=1+e^{-\eta\phi/f},$ (83)
normalized such that $W=1$ as $\phi$ rolls towards +$\infty$+ ∞ and with $\eta$ parametrizing the coupling strength. The initial condition $\phi_{0}$ must thus be negative so that $W(\phi_{0})\gg 1$.
Production begins when $H\approx m_{\phi}$ and continues until the latter term in Eq. 83 drops below unity, at which point the scalar and dark photon decouple, the dark photon’s mass reaches its present-day value, and production completes (when the Hubble rate is $H_{\star}$ by definition). Per the approximate solution Eq. 82, the scalar crosses zero (i.e., $W$ approaches unity) at $t_{\star}\approx e^{-\phi_{0}/2f}/m_{\phi}=1/M$. While $\phi<0$, we may take $W(\phi)\approx e^{-\eta\phi/f}$ and the time dependence of the effective frequency is approximately
$\frac{\omega_{\mathrm{eff}}(k)^{2}}{m_{\phi}^{2}}$ $\approx\frac{k^{2}}{a^{2}m_{\phi}^{2}}+\frac{m_{A}^{2}}{W(\bar{\phi})m_{\phi}^{2}}-\frac{\eta(2\eta+1)m_{\phi}^{2}t^{2}-3\eta}{2(1+m_{\phi}^{2}t^{2})^{2}}.$ (84)
While $W$ remains much larger than $(m_{A}/m_{\phi})^{2}$, the dark photon’s bare mass is negligible; modes with $k/a\lesssim\eta m_{\phi}(m_{\phi}t)^{-1}$ therefore have a negative $\omega_{\mathrm{eff}}^{2}$ and grow by a tachyonic resonance.
For modes with negligible wave numbers and at early enough times that the bare mass term can be neglected, the equation of motion Eq. 79 with effective frequencies given by Eq. 84 has the analytic solution (with $\dot{\mathcal{A}}_{\pm}$ initially zero)
$\mathcal{A}_{\pm}(t,0)$ $=\mathcal{A}_{\pm}(0,0)\left[1+(m_{\phi}t)^{2}\right]^{-\eta/2}.$ (85)
This result may be extended to small, nonzero $k$ with a perturbative expansion in the parameter $k/m_{\phi}$. Setting $a(t)=\sqrt{m_{\phi}t}$, we take the ansatz
$\mathcal{A}_{\pm}(t,k)$ $=\mathcal{A}_{\pm}(0,k)\left[1+(m_{\phi}t)^{2}\right]^{-\eta/2}\left[1+(k/m_{\phi})^{2}\delta\mathcal{A}(t)\right]$ (86)
and solve for the dimensionless perturbation $\delta\mathcal{A}(t)$. The equation of motion then reduces to one for $\delta\mathcal{A}$ of the form
00 0 $=\delta\ddot{\mathcal{A}}-\frac{(m_{\phi}t)^{2}(4\eta-1)-1}{2t\left[1+(m_{\phi}t)^{2}\right]}\delta\dot{\mathcal{A}}+\frac{m_{\phi}}{t}+\frac{k^{2}}{m_{\phi}t}\delta\mathcal{A}.$ (87)
The last term is negligible at leading order in $k/m_{\phi}$, and in this limit the equation is solved by an integral of a hypergeometric function:
$\delta\mathcal{A}(t)$ $=\int_{0}^{m_{\phi}t}\mathrm{d}x(1+x^{2})^{\eta}{}_{2}F_{1}(1/4,\eta,5/4;-x^{2})+\mathcal{O}(k^{2}/m_{\phi}^{2}).$ (88)
Eventually, either the gradient or bare mass term overtakes the $W$-derivative term as the dominant contribution to $\omega_{\mathrm{eff}}(k)^{2}$ and growth ceases. This behavior may be approximated by cutting off the integral Eq. 88 at the earlier of the two possibilities,
$t_{\mathrm{cut}}(k)$ $\approx\min\left[\frac{1}{M}\left(\sqrt{\eta^{2}+\eta/2}\frac{M}{m_{A}}\right)^{\frac{1}{\eta+1}},\frac{(\eta^{2}+\eta/2)m_{\phi}}{k^{2}}\right].$ (89)
As long as $m_{\phi}t_{\mathrm{cut}}(k)\gg 1$, the integral in Eq. 88 is dominated by $x\gg 1$, making it an excellent approximation to substitute $1+x^{2}\approx x^{2}$. The integral may then be evaluated analytically, yielding $\mathcal{A}_{\pm}(t,k)\propto(k/m_{\phi})^{2}(m_{\phi}t)^{\eta+1/2}$. After this point, the mode amplitude begins to decrease as $\mathcal{A}_{\pm}\propto W^{1/4}/a^{1/2}$ while the dark photon mass continues to increase toward its vacuum value, per the WKB approximation.
In summary, the mode amplitude evaluated at $t_{\star}$ is a product of factors encoding its initial condition $\mathcal{A}_{\pm}(0,k)\propto 1/\sqrt{2k/a}$, the $k=0$ solution [Eq. 85] and the leading-order correction $\delta\mathcal{A}$ for nonzero $k$ [Eq. 86], and factors encoding the adiabatic evolution between $t_{\mathrm{cut}}(k)$ and $t_{\star}$:
$\frac{\mathcal{A}_{\pm}(t_{\star},k)}{\mathcal{A}_{\pm}(0,k)}$ $=\frac{\mathcal{A}(t_{\mathrm{cut}},0)}{\mathcal{A}(0,0)}\left(\frac{k}{m_{\phi}}\right)^{2}\delta\mathcal{A}(t_{\mathrm{cut}})\sqrt[4]{\frac{W(\bar{\phi}(t_{\star}))}{W(\bar{\phi}(t_{\mathrm{cut}}))}}\sqrt{\frac{a(t_{\mathrm{cut}}(k))}{a(t_{\star})}}.$ (90)
Inserting the preceding results in the $m_{\phi}t\gg 1$ limit yields
$\frac{\mathcal{A}_{\pm}(t_{\star},k)}{\mathcal{A}_{\pm}(0,k)}$ $=\left[m_{\phi}t_{\mathrm{cut}}(k)\right]^{-\eta}\left(\frac{k}{m_{\phi}}\right)^{2}\left[m_{\phi}t_{\mathrm{cut}}(k)\right]^{2\eta+\frac{1}{2}}\frac{2\Gamma(\frac{5}{4})\Gamma(\eta-\frac{1}{4})}{\Gamma(\eta)(4\eta+1)}\left[2H_{\star}t_{\mathrm{cut}}(k)\right]^{\eta/2+1/4},$ (91)
where the Hubble rate at production
$H_{\star}\equiv 1/2t_{\star}=M/2$ (92)
is set by the time when $\phi$ crosses 00 and the scalar decouples.
The relic density of dark photons is obtained by a suitable integral of LABEL:{eqn:transverse-scalar-mode-amplitude} over wave number [Eq. 148]. The mode with peak power is that for which the mass and gradient terms are equally responsible for cutting off the resonance:
$k_{\star}$ $\approx\sqrt{(\eta^{2}+\eta/2)Mm_{\phi}}\left(\sqrt{\eta^{2}+\eta/2}\frac{M}{m_{A}}\right)^{-\frac{1}{2\eta+2}}.$ (93)
The mode amplitude is peaked at this wave number and falls off sharply at larger $k$, so the energy density is well approximated by integrating up to $k=k_{\star}$. The total dark photon energy density [Eq. 148] evaluated at $H=H_{\star}$ is thus approximately
$\frac{\rho_{\star}}{H_{\star}^{4}}$ $\approx\mathcal{N}_{\eta}\left(\frac{m_{\phi}}{M}\right)^{2\eta-1}\left(\frac{M}{m_{A}}\right)^{\frac{\eta-7/2}{\eta+1}},$ (94)
where the $\eta$-dependent coefficient is
$\mathcal{N}_{\eta}$ $=\frac{\left(\eta^{2}+\eta/2\right)^{\frac{9(2\eta+1)}{4(\eta+1)}}}{6\pi^{2}}\frac{\Gamma(5/4)^{2}\Gamma(\eta-1/4)^{2}}{(1+4\eta)^{2}\Gamma(\eta)^{2}}.$ (95)
In the large-$\eta$ limit, the power spectrum peaks at $k_{\star}\sim\eta\sqrt{Mm_{\phi}}\lesssim m_{A}$, so the dark photons are nonrelativistic when their mass reaches its final value and growth ends. Redshifting the energy density to matter-radiation equality, the dark photon relic abundance is
$\frac{\Omega_{A}}{\Omega_{\mathrm{DM}}}$ $\approx\frac{2}{3}\mathcal{N}_{\eta}\left(\frac{m_{\phi}}{2H_{\star}}\right)^{2\eta-1}\left(\frac{2H_{\star}}{m_{A}}\right)^{\frac{\eta-7/2}{\eta+1}}\frac{H_{\star}^{2}}{M_{\mathrm{pl}}^{2}}\sqrt{\frac{H_{\star}}{H_{\mathrm{eq}}}}.$ (96)
This result reproduces full numerical solutions within a couple of orders of magnitude for $-\phi_{0}/f\gtrsim 5$ [below which the errors are dominated by the approximation made in Eq. 82]—accuracy more than sufficient for our purposes.
For $\eta$ larger than a few, the relic abundance is most sensitive to $m_{\phi}/H_{\star}$, which effectively encodes how long production takes or equivalently the scalar’s initial displacement, $\phi_{0}=-2f\ln(m_{\phi}/H_{\star})$. The ratio $H_{\star}/m_{A}$ may then be chosen to delay production as long as desired. In addition, the initial dark photon mass must be far smaller than is necessitated by the kinematic condition $m_{A}(\phi_{0})\ll m_{\phi}$: neglecting the $\eta$-dependent prefactor in Eq. 96 and taking the large-$\eta$ limit (for which $H_{\star}\approx m_{\phi}$) with $\Omega_{A}=\Omega_{\mathrm{DM}}$ yields the estimate
$\frac{m_{A}^{2}}{W(\phi_{0})m_{\phi}^{2}}\sim\frac{H_{\star}^{4}}{m_{\phi}^{2}M_{\mathrm{pl}}^{2}}\frac{m_{A}}{m_{\phi}}\sqrt{\frac{H_{\star}}{H_{\mathrm{eq}}}}\ll 1.$ (97)
In other words, the initial condition for $\phi$ is determined primarily by requiring a sufficient dark photon abundance rather than the kinematic requirement that $W(\phi_{0})\gtrsim m_{A}^{2}/m_{\phi}^{2}$. The dynamics are thus effectively insensitive to the scalar–dark-photon mass ratio. This class of models thus certainly avoids backreaction onto the dark Higgs [Eq. 8] in the same parameter space as standard axion production (Sec. II.3), corresponding to $H_{\star}\sim m_{\phi}\gtrsim m_{A}$ [as depicted in Fig. 3]. But the suppression of the kinematic barrier also allows dark photon much heavier (today) than $m_{\phi}$. In fact, production may be delayed to at least as late as when $H_{\star}=10^{-22}\leavevmode\nobreak\ {\mathrm{eV}}$ (i.e., roughly the earliest time observations require the dark matter to exist and behave like matter)141414 Dark photon subcomponents of dark matter could be in principle be produced later with even larger kinetic mixing, but their direct-detection signals are also suppressed by their reduced abundance. Assessing precisely what fraction of dark matter could be produced at such late times, however, would require a dedicated cosmological analysis. and the kinetic coupling still eliminates the kinematic barrier for $m_{A}$ as large as $\sim(H_{\mathrm{eq}}/H_{\star})^{1/2}M_{\mathrm{pl}}^{2}/H_{\star}$.
The above calculations remain valid so long as the scalar’s dynamics are well approximated by Eq. 82, i.e., so long as backreaction of the dark photon onto the scalar and that of either field on the expansion history are negligible. The energy density of scalars in a runaway potentials tracks the critical density in both the radiation and matter dominated epochs: [197, 198, 199]
$\Omega_{\phi}(t)\equiv\frac{\bar{\rho}_{\phi}(t)}{3M_{\mathrm{pl}}^{2}H^{2}}$ $\approx\left(\frac{2f}{M_{\mathrm{pl}}}\right)^{2}.$ (98)
The expansion history thus only departs from the radiation solution near the onset of the matter era, as in standard cosmology. Comparing the scalar’s energy density to $\eta$ times the dark photon’s provides a decent proxy for whether backreaction onto the scalar is important. At production, the dark photon abundance (which redshifts more slowly than the radiation background by a power of $a$) is $\Omega_{A}\approx(H_{\mathrm{eq}}/H_{\star})^{1/2}/2$, so the scalar’s decay constant needs to be larger than
$f$ $\gtrsim 10^{-2}\sqrt{\eta}\left(\frac{H_{\star}}{10^{-22}\leavevmode\nobreak\ {\mathrm{eV}}}\right)^{-1/4}M_{\mathrm{pl}}$ (99)
to ensure backreaction never occurs.
The scalar’s tracking behavior, moreover, provides a natural means for the dark photon’s abundance to exceed the scalar’s in the late Universe: the scalar redshifts as radiation, faster than the nonrelativistic dark photons. In contrast, when dark photons are produced via a harmonically oscillating axion (Secs. II.3 and IV) or scalar, the parent field retains a sizeable energy density at late times, as discussed in Sec. II.3. For this reason, we implemented the termination of production by modifying $W$ from a pure exponential [Eq. 83] rather than augmenting the potential [Eq. 81] with a minimum at a finite field value. Barring some additional decay channel [54, 156, 159], production is terminated via nonlinear backreaction that can only be studied via $3+1$ dimensional simulations [54] and generically leads to two-component dark matter.
V.3.3 Quantum corrections
Unlike the axion, whose continuous shift symmetry protects its potential from perturbative quantum corrections, the scalar theories discussed in this section are susceptible to quantum corrections. Additional heavy degrees of freedom (that arise in a UV completion of the theory) may overwhelm any phenomenologically viable tree-level scalar potential as it rolls over large field ranges. Nonetheless, so long as the scalar only has self-interactions and a kinetic coupling to the dark photon, quantum corrections can remain small because of the smallness of $m_{A}$ at early times, only becoming important as $m_{A}$ reaches its vacuum value. We compute the one-loop scalar effective potential in scalar–Abelian-Higgs theory in Appendix B, from which we find that, for the particular choices of $W$ and $V$ above, the UV cutoff $\Lambda_{\mathrm{UV}}$ must be bounded above by
$\frac{\Lambda_{\mathrm{UV}}^{2}}{(4\pi f)^{2}}$ $<\frac{3H_{\star}^{2}}{2m_{A}^{2}}.$ (100)
The later production occurs, the lower the scalar’s potential and the smaller its quantum corrections must be to ensure the dynamics are not disrupted. Taking $\Lambda_{\mathrm{UV}}\sim 4\pi f$ as a criterion for naturalness sets $H_{\star}\gtrsim m_{A}$ for scalar–dark-photon production—that is, we expect that below the postinflationary line in Fig. 3 the scalar models of dark photon production we discuss are both detectable and not fine tuned in this sense. Larger kinetic mixing (without backreaction) is available to scalar-produced dark photons to the extent that one allows the models to be fine tuned, as illustrated by Fig. 8.
Figure 8: Parameter space in mass and kinetic mixing [Eq. 9] of dark photon dark matter produced via kinetically coupled scalars. Thin red lines depict the maximum allowed kinetic mixing for a given degree of tuning (as labeled) of either the UV cutoff $\Lambda_{\mathrm{UV}}/4\pi f$ for which one-loop quantum corrections are unimportant for runaway scalars (Sec. V.3) or the dark-photon–scalar mass tuning $\delta$ [Eq. 115] for oscillating scalars (Sec. V.4). The red shaded region depicts parameter space in which backreaction or defect formation occurs even if the dark photon is produced as late as allowed by the CMB (roughly $H_{\star}\approx 10^{-22}\leavevmode\nobreak\ {\mathrm{eV}}$). Gray and yellow regions respectively depict the current limits and projected reach of various probes, as labeled in Fig. 3.
The above discussion, taking only quadratically divergent contributions to the effective potential, derives from the one-loop computations in Appendix B which show that properly quantizing noncanonical field theories leads to a subtle cancellation of would-be quartic divergences (that one might intuitively expect to be present). Appendix B does not, however, attempt to determine whether such divergences are present at higher-loop order. Supposing that a quartic divergence was incurred at two loops by the kinetic coupling $W$, it would also be suppressed by $g_{D}^{2}$. Requiring $g_{D}^{2}\Lambda_{\mathrm{UV}}^{4}\lesssim H_{\star}^{2}f^{2}$ imposes that $g_{D}\lesssim H_{\star}/f\cdot(f/\Lambda_{\mathrm{UV}})^{2}$—nominally quite a constraining bound, given that it combines with the lower bound on $f$ [Eq. 99] to set $g_{D}\lesssim 1.3\times 10^{-20}\left(H_{\star}/\mathrm{eV}\right)^{5/4}/\sqrt{\eta}$ (taking $\Lambda_{\mathrm{UV}}\approx f$). This bound happens to coincide with that on standard inflationary production [Eq. 12] in scaling (with $H_{\star}$ in place of $m_{A}$) and nearly in magnitude and is therefore extremely constraining for delayed production. Suppose, however, that $m_{\phi}\gg m_{A}$ (so that production is instead early) and that $k_{\star}/a_{\star}\sim H_{\star}$. The generic backreaction bound Eq. 8 on transverse production then scales quite weakly with $H_{\star}$: $g_{D}\lesssim 10^{-14}\lambda^{1/4}\left(m_{A}/\mu\mathrm{eV}\right)^{5/8}\left(m_{A}/H_{\star}\right)^{1/8}$. The parameter space that is natural can be made larger than that which avoids backreaction simply by increasing $H_{\star}/m_{A}$ by, say, eight orders of magnitude, yielding a bound only one decade tighter than the fiducial case where $H_{\star}=m_{A}$. We therefore conclude that, even for the worst-case (and unconfirmed) UV divergences, the runaway scalar model allows for natural and backreaction-free parameter space in most of the fiducial postinflationary regime.
Since this class of production mechanism relies on the dark Abelian-Higgs sector being extremely weakly coupled at early times, one might worry that it runs afoul of weak gravity conjectures (WGCs) [200]. While it is plausible but uncertain whether the WGCs constrain the gauge couplings for displaced moduli [178], were this the case, they would limit the field range of the scalar field and possibly interfere with dark photon production. The dark photon mass and hence the gauge coupling must be very small at early times; their sizes are quantified by Eq. 97, which translates to an estimate for the initial gauge coupling of
$g_{D}(\phi_{0})$ $\sim g_{D}\frac{H_{\star}}{M_{\mathrm{pl}}}\sqrt{\frac{H_{\star}}{m_{A}}}\sqrt[4]{\frac{H_{\star}}{H_{\mathrm{eq}}}}$ (101)
in the large-$\eta$ limit. The tower WGC [201, 202, 203] implies a UV cutoff of $g_{D}(\phi_{0})^{1/3}M_{\mathrm{pl}}$. Using the upper limit on $g_{D}$ implied by the backreaction bound Eq. 8, we find that the UV cutoff associated with the WGC is
$g_{D}$ $\lesssim 3\times 10^{-18}\cdot\lambda^{7/22}\left(\frac{m_{A}}{10^{-15}\leavevmode\nobreak\ {\mathrm{eV}}}\right)^{25/22}\left(\frac{H_{\mathrm{max}}}{10^{-15}\leavevmode\nobreak\ {\mathrm{eV}}}\right)^{-9/22},$ (102)
where $H_{\mathrm{max}}$ is the maximum Hubble rate of the Universe and we’ve required $T_{\mathrm{max}}\sim\sqrt{M_{\mathrm{pl}}H_{\mathrm{max}}}<\Lambda_{\mathrm{UV}}$. Figure 9 depicts the range of parameter space in which this model would violate the WGC at early times for UV cutoffs of $1\leavevmode\nobreak\ \mathrm{TeV}$ and $10^{16}\leavevmode\nobreak\ \mathrm{GeV}$ (the latter corresponding to the energy scale of high-scale inflation).
V.3.4 Signatures
In the regime of extremely delayed production, where the scalar potential must be protected by a low effective UV cutoff, the signatures of production via a rapidly varying gauge coupling are particularly striking. Small-scale structure is significantly enhanced for sufficiently small $H_{\star}$ and is potentially observable. In principle, the since power below the peak (at which the power spectrum is order unity) scales as $(k/k_{\star})^{3}$, cosmological scales could also be enhanced above that of the baseline adiabatic fluctuations (whose power is $\sim 10^{-9}$) if $k_{\star}$ is within three decades of an observable scale. In our case, the peak wave number is $k_{\star}\approx 2\eta H_{\star}$ in the large-$\eta$ limit. As a rough guide, at matter-radiation equality $k_{\star}/a_{\mathrm{eq}}\sim 2\beta H_{\mathrm{eq}}\sqrt{H_{\star}/H_{\mathrm{eq}}}$. If production terminates long enough before equality (say, by three decades of expansion), then $\sqrt{H_{\star}/H_{\mathrm{eq}}}\gtrsim 10^{3}$; with $\beta\gtrsim 10$, enhanced power is important on scales at least an order of magnitude smaller than the horizon at equality. These scales are on the threshold of the sensitivity of Planck’s CMB observations and around those which the Lyman-$\alpha$ forest probe [204, 205, 206]. Further in the future, 21-cm intensity mapping could probe scales one to two order of magnitude smaller [207]. Ultimately, quantitative cosmological constraints require a dedicated analysis of dark matter with such power spectra (as would a bound on the latest time the dark matter could be produced).
Enhanced power on yet smaller scales would approach the critical overdensity $\delta_{c}=1.686$ during radiation domination [208], leading to the collapse of structures with radius $\sim\pi/(2k_{\star}/a_{\mathrm{eq}})$ around matter-radiation equality. (The peak of the density power spectrum is at roughly twice $k_{\star}$.) The density of these structures is approximately $178$ times that of the ambient matter at matter-radiation equality [209],
$\rho_{\mathrm{coll}}$ $\sim 178\cdot\frac{3}{2}M_{\mathrm{pl}}^{2}H_{\mathrm{eq}}^{2}\approx 2.7\times 10^{5}M_{\odot}/\mathrm{pc}^{3},$ (103)
and these collapsed structures contain a mass151515 Although gravitational interactions can change the subhalo’s mass and density over time [210], we calculate their initial values at collapse for a conservative representation of their late-time properties.
$M$ $\sim\frac{3}{2}M_{\mathrm{pl}}^{2}H_{\mathrm{eq}}^{2}\cdot\frac{4\pi}{3}\left(\frac{\pi}{2k_{\star}/a_{\mathrm{eq}}}\right)^{3}\approx 2.4\times 10^{5}M_{\odot}\left(\frac{\eta}{10}\right)^{-3}\left(\frac{H_{\star}}{10^{-22}\leavevmode\nobreak\ {\mathrm{eV}}}\right)^{-3/2}.$ (104)
Power on larger scales is also moderately enhanced, resulting in the collapse of larger, more diffuse structures at later times. The structures collapsing at $z=250$ are the largest that remain resilient against tidal stripping [211]; their density is again $178$ times the matter density at the time of collapse, $\rho_{\mathrm{coll}}\sim 9.2\times 10^{1}M_{\odot}/\mathrm{pc}^{3}$. We may estimate their mass by extrapolating the density power spectrum from its peak to small wave number, i.e. $\sim(2k/k_{\star})^{3}$, where the factor of two accounts for the broad peak of the density power spectrum. Since the density power spectrum is proportional the square of the fractional overdensity, it grows with $a^{2}$ during the matter era, so modes of ever longer wavelength $a/k\propto a^{5/3}$ collapse as the Universe continues to expand. On the other hand, the matter density dilutes as $a^{-3}$. In total, the mass of structures collapsing at a given redshift grows with $a^{2}$, and we estimate that structures collapsing at $z=250$ have mass
$M\sim\frac{3}{2}\frac{M_{\mathrm{pl}}^{2}H_{\mathrm{eq}}^{2}}{(a/a_{\mathrm{eq}})^{3}}\cdot\frac{4\pi}{3}\left(\frac{\pi}{k_{\mathrm{col}}/a_{\mathrm{eq}}}\right)^{3}$ $\sim 2.9\times 10^{9}M_{\odot}\left(\frac{\eta}{10}\right)^{-3}\left(\frac{H_{\star}}{10^{-22}\leavevmode\nobreak\ {\mathrm{eV}}}\right)^{-3/2}.$ (105)
Figure 9 depicts the expected substructure mass for maximally delayed dark photon production as a function of $m_{A}$ and $\varepsilon$, superimposed on the limits and prospective reach of direct detection experiments.
Figure 9: Minimum dark matter substructure sizes predicted by the kinetically coupled runaway scalar model in the direct detection parameter space. Blue lines display substructure masses $M_{\mathrm{densest}}$ for structures collapsing at the onset of matter-radiation equality and $M_{\mathrm{max}}$ for structures collapsing at redshift $z=250$ in scenarios where production is delayed until Hubble rates $H_{\star}$ as labeled on each line. Much of the expanded kinetic mixing parameter space—including that accessible to future direct detection programs—features substructures that are potentially accessible to upcoming astrometric ($M>10^{-3}M_{\odot}$[210] and photometric searches ($10^{-15}M_{\odot}<M<10^{-8}M_{\odot}$[212]. Dashed black lines represent upper bounds on the dark gauge coupling implied by the tower Weak Gravity Conjecture [202, 201, 203], should it apply to displaced moduli. Results are otherwise presented as in Figs. 3 and 8.
Note that production may well be delayed longer than strictly necessary to accommodate whatever kinetic mixing the dark photon happens to have; the substructure masses displayed in Fig. 9 are thus the minimum predicted by the model (rather than a unique value thereof).
Astrometric lensing, the correlated deflection in background stars induced by a dense, heavy halo passing through our line of sight, is a promising indirect signature of such substructure [213]. High-resolution telescopes like Gaia [214] and the Hubble Space Telescope [215], and future missions such as Theia [216], Roman [217], the Square Kilometre Array [218], and the Thirty Meter Telescope [219], could detect these effects by measuring the correlated proper motions and accelerations of stars. Gaia is sensitive to subhalos heavier than $10^{5}M_{\odot}$ for the densest subhalos that form around matter-radiation equality, while Theia and SKA will probe these dense substructures down to masses around $10^{-3}M_{\odot}$ and the more dilute structures forming at $z=250$ down to $10^{2}M_{\odot}$  [210]. Another signature is photometric microlensing, where changes in a star’s brightness suggest an intervening gravitational lens. This technique is particularly effective with highly magnified stars near critical lensing caustics of galaxy clusters. If dark matter subhalos are present, their virial motion introduces noise in the star’s position and brightness, depending on the subhalo mass spectrum [212]. However, these projections are subject to uncertainties from galactic evolution and tidal stripping of subhalos and require further simulation to confirm their viability [211].
The scalar coupled to the dark photon may yield its own signatures; their inherent model dependence could facilitate positively identifying the origin of dark photon dark matter. In the case of the runaway scalar considered in Sec. V.3.2, delaying production requires large enough decay constants $f$ to ensure resonant growth is not disrupted, as well as to supply the dark photon with sufficient energy to make up the dark matter. The model thus predicts some minimum cosmological abundance of the scalar in both the radiation and matter eras due to its tracking behavior [Eq. 98], even if the dark photon dominates the dark matter abundance. The runaway scalar would alter cosmological dynamics to a degree controlled by $f/M_{\mathrm{pl}}$. Recent work [220, 221] has quantitatively analyzed cosmology with tracking scalars as a means to resolve the Hubble tension, the mismatch between Hubble constants inferred from the early Universe (e.g., via CMB data [56]) and measured directly by the calibrated distance ladder [222]. In the scenarios studied in Refs. [220, 221], current data limit the fractional abundance of the scalar at early times at the few-percent level. Extrapolating this result to the runaway scalars we consider estimates that $f$ must be below about a tenth of the Planck scale.
Finally, though production must end sufficiently early that the dark matter exists well before matter-radiation equality, the dark photon’s mass need not necessarily stop evolving at this point. The coupling function [Eq. 83], rather than asymptoting to unity, could instead transition to a slower evolution with time. While not necessary for the scenario we identify, a concrete realization thereof may well predict that the dark photon mass continues to evolve to some extent after production, a possibility that can also be tested by cosmological data [223, 224].
V.4 Narrow resonance with an oscillating scalar
The final possibility we consider is, rather than a runaway scalar, a scalar oscillating in a quadratic potential
$V(\phi)$ $=\frac{1}{2}m_{\phi}^{2}\phi^{2}.$ (106)
Such a scenario was studied in Ref. [51], which showed that a kinetic coupling $W(\phi)=e^{\eta\phi/f}$ leads not just to broad tachyonic resonance familiar from analogous models of gauge preheating [144, 225, 147, 148] but also to a unique regime of narrow resonance when $m_{A}/m_{\phi}=1/2$ that can in principle persist for arbitrarily small scalar field amplitudes (and so an arbitrarily long time). As a result, conversion of the scalar into dark photons eventually becomes fast compared to the Hubble rate without requiring large coupling [51]—indeed, only relying on the leading-order expansion $W(\phi)\approx 1+\eta\phi/f$, which would be expected to dominate the behavior from the perspective of bottom-up effective field theory.
Moreover, smaller couplings delay the time that dark photons are efficiently produced, which Ref. [51] posited as a means to alleviate defect formation bounds. Before examining this possibility in detail, we briefly review the mechanics of this scenario, focusing on the novel regime of persistent resonance at small scalar field amplitudes.161616 The broad resonance scenario (i.e., $\eta\phi_{0}/f>1$ for any mass $m_{A}\lesssim\eta m_{\phi}$) is certainly a viable production mechanism, but its phenomenology (in terms of allowed kinetic mixing, at least) is identical to axion production. We also show that the structure of the equations of motion that gives rise to the small-amplitude regime is not unique to the kinetic coupling considered in Ref. [51].
The solution to the homogeneous equation of motion Eq. 60a for a massive, misaligned scalar that begins oscillating in the radiation era [when $a(t)\propto\sqrt{t}$] is
$\bar{\phi}(t)$ $=\Gamma(5/4)\phi_{0}\frac{J_{1/4}(m_{\phi}t)}{\sqrt[4]{m_{\phi}t/2}}$ (107)
from initial conditions $\bar{\phi}=\phi_{0}$ and $\dot{\bar{\phi}}=0$ at times $m_{\phi}t\to 0$. Define $a_{\mathrm{osc}}=a(t_{\mathrm{osc}})$ where $H(t_{\mathrm{osc}})=m_{\phi}$, such that $H(t)=m_{\phi}[a(t)/a_{\mathrm{osc}}]^{-2}$. At late times $m_{\phi}t\gg 1$,
$\bar{\phi}(t)$ $\approx\frac{2^{3/2}\Gamma(5/4)\phi_{0}}{\sqrt{\pi}[a(t)/a_{\mathrm{osc}}]^{3/2}}\sin\left(m_{\phi}t+\pi/8\right)\equiv\Phi(t)\sin\left(m_{\phi}t+\pi/8\right),$ (108)
where the second equality defines the time-dependent oscillation amplitude $\Phi(t)$.
Expanding coupling functions in Eq. 60b to leading order in small $\bar{\phi}$ and neglecting Hubble friction (as appropriate at the late times when resonance becomes efficient),
00 0 $=\ddot{\mathcal{A}}_{\pm}+\left[\frac{k^{2}}{a^{2}}+m_{A}^{2}+m_{A}^{2}\bar{\phi}(t)\left(\frac{X^{\prime}(\bar{\phi})}{X(\bar{\phi})}-\frac{W^{\prime}(\bar{\phi})}{W(\bar{\phi})}\right)+\ddot{\bar{\phi}}(t)\frac{W^{\prime}(\bar{\phi})}{2W(\bar{\phi})}\right]\mathcal{A}_{\pm}.$ (109)
Substituting Eq. 108 and taking the time coordinate $z=m_{\phi}t/2-3\pi/16$, Eq. 109 takes the form of the Mathieu equation,
00 0 $=\frac{\mathrm{d}^{2}\mathcal{A}_{\pm}}{\mathrm{d}z^{2}}+\left[p-2q\cos(2z)\right]\mathcal{A}_{\pm},$ (110)
with $m_{\phi}^{2}p/4=k^{2}/a^{2}+m_{A}^{2}$ and
$q$ $=\Phi(t)\left[\frac{2m_{A}^{2}}{m_{\phi}^{2}}\left(\frac{X^{\prime}(\bar{\phi})}{X(\bar{\phi})}-\frac{W^{\prime}(\bar{\phi})}{W(\bar{\phi})}\right)-\frac{W^{\prime}(\bar{\phi})}{W(\bar{\phi})}\right].$ (111)
If the mass ratio $m_{A}/m_{\phi}$ is precisely $1/2$ then a narrow resonance for $k\ll m_{\phi}$ persists to arbitrarily small $q$ (i.e., field amplitudes) [51]. (Otherwise—as is typical of such scenarios—the resonance terminates once the scalar redshifts below some nonzero $\bar{\phi}$.)
Evidently, Eq. 111 is nonzero regardless of whether the coupling is via the kinetic term ($W$) or the mass term ($X$) or both (i.e., a universal coupling); only coupling functions contrived to make Eq. 111 identically zero would exhibit no such instability. Furthermore, longitudinal modes experience the same instability: as discussed below Eq. 60, in the $k\ll m_{A}(\bar{\phi})$ limit where the small-amplitude resonance would exist if present, the longitudinal mode’s dynamics [Eq. 60c] are identical to the transverse modes’.
The narrow resonance scenario allows for dark photon production parametrically later than the standard expectation that $H_{\star}\gtrsim m_{A}$. Modes in the narrow resonance band have exponential growth rate $\Gamma=\eta m_{\phi}\Phi(t)/8f$, which decays more slowly than the Hubble rate in the radiation era. Production therefore becomes efficient at a scale factor [51]
$\frac{a_{\star}}{a_{\mathrm{osc}}}$ $=\left(\frac{3\eta\phi_{0}}{4f}\right)^{-2}\ln\left(\frac{m_{\phi}^{2}}{m_{A}^{2}}\frac{\phi_{0}^{2}}{m_{\phi}^{2}}\right)^{2},$ (112)
where $\phi_{0}$ is the initial misalignment of the scalar. The corresponding Hubble rate at production is
$\frac{H_{\star}}{m_{A}}$ $=2\left(\frac{3\eta\phi_{0}}{4f}\right)^{4}\ln\left(\frac{m_{\phi}^{2}}{m_{A}^{2}}\frac{\phi_{0}^{2}}{m_{\phi}^{2}}\right)^{-4}.$ (113)
Since nonrelativistic dark photons are produced, the energy density in the sector always redshifts as $a^{-3}$ (after $a_{\mathrm{osc}}$); the relic abundance is therefore simply that inherited from the scalar’s misalignment:
$\frac{\Omega_{A}}{\Omega_{\mathrm{DM}}}$ $\approx\left(\frac{m_{A}}{2\times 10^{-18}\leavevmode\nobreak\ {\mathrm{eV}}}\right)^{1/2}\left(\frac{\phi_{0}}{10^{16}\leavevmode\nobreak\ \mathrm{GeV}}\right)^{2}.$ (114)
As suggested by Ref. [51], such delayed production could allow for larger $g_{D}$ while avoiding defect formation, akin to the mechanisms discussed above. Moreover, since $W(\phi)$ need not be particularly large (or small) for efficient production, the effect of the scalar on the parameters of the theory [Eq. 59] is negligible. However, the persistence of narrow resonance to small amplitudes $\Phi(t)$ (and therefore at late times) depends how precisely the mass ratio $m_{A}/m_{\phi}$ is tuned to $1/2$. Concretely, the instability persists until $a_{\star}$ only if the tuning parameter [51]
$\delta$ $\equiv\left(m_{A}/m_{\phi}\right)^{2}-1/4$ (115)
satisfies
${\left|\delta\right|}<\frac{3\eta\phi_{0}}{16f(a_{\star}/a_{\mathrm{osc}})^{3/2}}.$ (116)
Inserting Eq. 113 for $\eta\phi_{0}/f$, production thus may only be delayed to a Hubble rate
$\frac{H_{\star}}{m_{A}}$ $>4{\left|\delta\right|}\ln\left(\frac{m_{\phi}^{2}}{m_{A}^{2}}\frac{\phi_{0}^{2}}{m_{\phi}^{2}}\right)^{-1}\gtrsim 0.02{\left|\delta\right|},$ (117)
using the relic abundance Eq. 114 evaluated with a fiducial scalar mass $m_{\phi}=4\times 10^{-18}\leavevmode\nobreak\ {\mathrm{eV}}$. Using Eq. 8, the maximum dark gauge coupling that evades defect formation is then
$g_{D}$ $\lesssim 4\times 10^{-14}\,\lambda^{1/4}{\left|\delta\right|}^{-3/8}\left(\frac{m_{A}}{\mu{\mathrm{eV}}}\right)^{5/8}\left(\frac{\Omega_{A}}{\Omega_{\mathrm{DM}}}\right)^{-1/4}.$ (118)
The allowed kinetic mixing parameter space as a function of $\delta$ is depicted in Fig. 8.
V.4.1 Thermal masses and tuning
Even assuming the existence of a UV completion that guarantees such a mass tuning that is robust to quantum corrections, in-medium effects inevitably source a time-dependent mass for the dark photon via its kinetic mixing with the SM photon. To assess the impact of the SM plasma on this mechanism, we extend the scalar–Abelian-Higgs theory to explicitly include kinetic mixing as well as a current term representing nonrelativistic electrons:
$\begin{split}\mathcal{L}&=\mathcal{L}_{\mathrm{SAH}}-\frac{\varepsilon}{2}F_{\mu\nu}F_{\mathrm{SM}}^{\mu\nu}-\frac{1}{4}F^{\mathrm{SM}}_{\mu\nu}F_{\mathrm{SM}}^{\mu\nu}-eA^{\mathrm{SM}}_{\mu}J^{\mu},\end{split}$ (119)
where $J^{\mu}=n_{e}u_{e}^{\mu}$ and $n_{e}$ is the electron number density and $u_{e}^{\mu}$ is their velocity. We remove the explicit kinetic mixing by redefining the SM photon field to
$\mathsf{A}^{\mu}$ $=A_{\mathrm{SM}}^{\mu}+\varepsilon A^{\mu}$ (120)
and define the electromagnetic field strength tensor $\mathsf{F}_{\mu\nu}\equiv\partial_{\mu}\mathsf{A}_{\nu}-\partial_{\nu}\mathsf{A}_{\mu}$. In this basis, the Lagrangian takes the form
$\begin{split}\mathcal{L}&=\frac{1}{2}\partial_{\mu}\phi\partial^{\mu}\phi-V(\phi)-\frac{W(\phi)-\varepsilon^{2}}{4}F_{\mu\nu}F^{\mu\nu}+\frac{X(\phi)}{2}m_{A}^{2}A_{\mu}A^{\mu}-\frac{1}{4}\mathsf{F}_{\mu\nu}\mathsf{F}^{\mu\nu}-eJ_{\mu}\left(\mathsf{A}^{\mu}-\varepsilon A^{\mu}\right).\end{split}$ (121)
The formalism developed in the absence of mixing must only be augmented by including the SM current in the equations of motion and replacing $W(\phi)$ with $W(\phi)-\varepsilon^{2}$.
For simplicity we consider only transverse modes; the equations of motion for $\mathsf{A}_{\pm}$ and $\mathcal{A}_{\pm}$ for the Lagrangian Eq. 121 are
$eJ_{\pm}$ $=\ddot{\mathsf{A}}_{\pm}+H\dot{\mathsf{A}}_{\pm}+\frac{k^{2}}{a^{2}}\mathsf{A}_{\pm}$ (122a)
$-e\varepsilon(\bar{\phi})J_{\pm}$ $=\ddot{\mathcal{A}}_{\pm}+H\dot{\mathcal{A}}_{\pm}+\left[\frac{k^{2}}{a^{2}}+m_{A}(\bar{\phi})^{2}-\frac{H}{2}\frac{\partial_{t}W(\bar{\phi})}{W(\bar{\phi})-\varepsilon^{2}}-\frac{\partial_{t}^{2}\sqrt{W(\bar{\phi})-\varepsilon^{2}}}{\sqrt{W(\bar{\phi})-\varepsilon^{2}}}\right]\mathcal{A}_{\pm}.$ (122b)
In Appendix E, we show that the transverse modes of the current evolve approximately by
$\dot{J}_{\pm}+\nu(t)J_{\pm}$ $=\frac{\omega_{p}(t)^{2}}{e}\left[\dot{\mathsf{A}}_{\pm}-\varepsilon(\bar{\phi})\dot{\mathcal{A}}_{\pm}\right],$ (123)
where the damping coefficient $\nu(t)$ is defined in Eq. 247a, and its value relative to the plasma frequency $\omega_{p}\equiv\sqrt{e^{2}n_{e}/m_{e}}$ is given by Eq. 250. For simplicity, we take $W=1$ so that $\varepsilon$ is constant. In this case, the factor of $1-\varepsilon^{2}$ multiplying the kinetic term in Eq. 121 may be absorbed via a redefinition of $g_{D}$.
We reduce the system to a set of third-order equations of motion for $\mathsf{A}_{\pm}$ and $\mathcal{A}_{\pm}$ alone by differentiating Eq. 122 with respect to time, inserting Eq. 123 for $\dot{J}_{\pm}$, and then substituting Eq. 122a for $J_{\pm}$. The resulting equations may then be written in the form
$\begin{split}0&=\frac{\mathrm{d}{}}{\mathrm{d}{t}}\left(\ddot{\mathsf{A}}_{\pm}+\left[H+\nu\right]\dot{\mathsf{A}}_{\pm}+\left[\frac{k^{2}}{a^{2}}+\omega_{p}^{2}-\dot{\nu}+H\nu\right]\mathsf{A}_{\pm}-\varepsilon\omega_{p}^{2}\mathcal{A}_{\pm}\right)\\ &\hphantom{{}={}}+\left(\ddot{\nu}-H\dot{\nu}+\nu\frac{k^{2}}{a^{2}}-\nu\dot{H}-2\omega_{p}\dot{\omega}_{p}\right)\mathsf{A}_{\pm}+2\omega_{p}\dot{\omega}_{p}\mathcal{A}_{\pm}\end{split}$ (124a)
$\begin{split}0&=\frac{\mathrm{d}{}}{\mathrm{d}{t}}\left(\ddot{\mathcal{A}}_{\pm}+H\dot{\mathcal{A}}_{\pm}+\left[\frac{k^{2}}{a^{2}}+m_{A}(\bar{\phi})^{2}+\varepsilon^{2}\omega_{p}^{2}\right]\mathcal{A}_{\pm}-\varepsilon\dot{\nu}\dot{\mathsf{A}}_{\pm}+\left[\dot{\nu}-H\nu-\omega_{p}^{2}\right]\varepsilon\mathsf{A}_{\pm}\right)\\ &\hphantom{{}={}}+\varepsilon\left(2\omega_{p}\dot{\omega}_{p}+\left[\dot{H}-\frac{k^{2}}{a^{2}}+H\dot{\nu}-\ddot{\nu}\right]\right)\mathsf{A}_{\pm}-2\varepsilon^{2}\omega_{p}\dot{\omega}_{p}\mathcal{A}_{\pm}.\end{split}$ (124b)
If the second lines of each equality in Eq. 124 are negligible, then the system reduces to second order in time with additional damping and mass contributions from the plasma. Since $\dot{\omega}_{p}/\omega_{p}\propto H$ and $\dot{\nu}/\nu\propto H$, this condition is met by taking both $H$ and $k/a$ small compared to $\omega_{p}$ and $m_{A}(\bar{\phi})$. Equation 124 then simplifies to
00 0 $=\ddot{\mathsf{A}}_{\pm}+\nu\dot{\mathsf{A}}_{\pm}+\omega_{p}^{2}\mathsf{A}_{\pm}-\varepsilon\omega_{p}^{2}\mathcal{A}_{\pm}$ (125a)
00 0 $=\ddot{\mathcal{A}}_{\pm}+\left[m_{A}(\bar{\phi})^{2}+\varepsilon^{2}\omega_{p}^{2}\right]\mathcal{A}_{\pm}-\varepsilon\nu\dot{\mathsf{A}}_{\pm}-\varepsilon\omega_{p}^{2}\mathsf{A}_{\pm}.$ (125b)
Equation 250 (and the surrounding discussion) shows that the relevant regime for $T\ll m_{e}$ is that of negligible damping, $\nu\ll\omega_{p}$.171717 As a matter of interest, the system decouples in the limit of infinite damping ($\nu\gg\omega_{p}$) because the SM photon slowly rolls in this case. That is, $\dot{\mathsf{A}}_{\pm}\approx-(\omega_{p}^{2}/\nu)\left({\mathsf{A}}_{\pm}-\varepsilon{\mathcal{A}}_{\pm}\right)$ solves Eq. 125a when $\ddot{\mathsf{A}}_{\pm}$ is neglected at leading order in $\omega_{p}/\nu$. As a result, $\ddot{\mathcal{A}}_{\pm}\approx-m_{A}(\bar{\phi})^{2}\mathcal{A}_{\pm}$. In other words, if the plasma is extremely resistive to currents, it decouples from the dark photon. In this case, were $m_{A}(\bar{\phi})^{2}+\varepsilon^{2}\omega_{p}^{2}$ constant we could simply diagonalize the equations and solve for the mode functions. Though $m_{A}(\bar{\phi})$ is explicitly not constant, in the narrow resonance band its relative variations are small. Writing this small perturbation in the form $m_{A}(\bar{\phi})^{2}=m_{A}^{2}+\delta m_{A}(\bar{\phi})^{2}$, we may meaningfully define new fields
$\begin{pmatrix}a_{\pm}\\ b_{\pm}\end{pmatrix}$ $=\begin{pmatrix}\cos\varphi&\sin\varphi\\ -\sin\varphi&\cos\varphi\end{pmatrix}\begin{pmatrix}\mathsf{A}_{\pm}\\ \mathcal{A}_{\pm}\end{pmatrix}\approx\begin{pmatrix}1&\varepsilon/[1-m_{A}^{2}/\omega_{p}^{2}]\\ -\varepsilon/[1-m_{A}^{2}/\omega_{p}^{2}]&1\end{pmatrix}\begin{pmatrix}\mathsf{A}_{\pm}\\ \mathcal{A}_{\pm}\end{pmatrix},$ (126)
where
$\frac{1}{2}\tan 2\varphi$ $=\frac{\varepsilon}{1-\varepsilon^{2}-m_{A}^{2}/\omega_{p}^{2}}$ (127)
and the second equality in Eq. 126 shows the result at leading order in $\varepsilon$. Then to quadratic order in $\varepsilon$ and $\delta m_{A}(\bar{\phi})^{2}/m_{A}^{2}$, these fields evolve according to
00 0 $=\ddot{a}_{\pm}+\omega_{p}^{2}\left(1+\frac{\varepsilon^{2}}{1-m_{A}^{2}/\omega_{p}^{2}}\right)a_{\pm}-\varepsilon\frac{\delta m_{A}(\bar{\phi})^{2}}{1-m_{A}^{2}/\omega_{p}^{2}}b_{\pm}$ (128a)
00 0 $=\ddot{b}_{\pm}+m_{A}(\bar{\phi})^{2}\left(1-\frac{\varepsilon^{2}}{1-m_{A}^{2}/\omega_{p}^{2}}\right)b_{\pm}-\varepsilon\frac{\delta m_{A}(\bar{\phi})^{2}}{1-m_{A}^{2}/\omega_{p}^{2}}a_{\pm}.$ (128b)
If $\ddot{a}_{\pm}\sim\omega_{p}a_{\pm}$, then $a_{\pm}\sim\varepsilon\delta m_{A}(\bar{\phi})^{2}/(\omega_{p}^{2}-m_{A}^{2})\cdot b_{\pm}$. So long as $\delta m_{A}(\bar{\phi})^{2}$ is small compared to either $\omega_{p}^{2}$ or $m_{A}^{2}$ (the latter of which is certainly the case), the contribution of $a_{\pm}$ to the equation of motion for $b_{\pm}$ is negligible and the two equations decouple. Plasma effects therefore do not disrupt resonance when their contribution to the effective mass of $b_{\pm}$ is smaller than the required tuning [defined in Eq. 115]:
$\delta$ $\gtrsim\varepsilon^{2}\frac{\min\left(\omega_{p}^{2},m_{A}^{2}\right)}{4\max\left(\omega_{p}^{2},m_{A}^{2}\right)}.$ (129)
Note that this analysis breaks down if $m_{A}$ crosses $\omega_{p}$.
Since the thermal mass decreases monotonically with time, Eq. 129 need only hold at the latest possible production time allowed by the degree of bare mass tuning. Inserting Eq. 249 for $\omega_{p}$ and then evaluating at $H_{\star}$ according to Eq. 117 yields
$\varepsilon$ $\lesssim 1.76\times 10^{-8}\begin{dcases}\left(\frac{m_{A}}{6.28\times 10^{-5}\leavevmode\nobreak\ {\mathrm{eV}}}\right)^{-3/2}\left(\frac{H_{\star}}{10^{-22}\leavevmode\nobreak\ {\mathrm{eV}}}\right)^{5/4}\sqrt{\frac{\eta_{B}}{6\times 10^{-10}}},&m_{A}\ll\omega_{p},\\ \left(\frac{m_{A}}{6.28\times 10^{-5}\leavevmode\nobreak\ {\mathrm{eV}}}\right)^{1/2}\left(\frac{H_{\star}}{10^{-22}\leavevmode\nobreak\ {\mathrm{eV}}}\right)^{-1/4}\sqrt{\frac{6\times 10^{-10}}{\eta_{B}}},&m_{A}\gg\omega_{p},\end{dcases}$ (130)
where $\eta_{B}$ is the baryon-to-photon ratio. [The plasma frequency is $6.28\times 10^{-5}\leavevmode\nobreak\ {\mathrm{eV}}$ when $H_{\star}=10^{-22}\leavevmode\nobreak\ {\mathrm{eV}}$, so Eq. 130 is written relative to the most constrained dark photon mass.] Notably, the kinetic mixing can be much larger than the maximum that would otherwise be allowed by backreaction [Eq. 8]. The detuning due to plasma effects redshifts faster than the tuning required for efficient resonance, and the more tuned the bare masses are, the later production may viably occur (i.e., by reducing $\eta$ as much as possible). Moreover, at a fixed bare mass tuning, Eq. 130 is a weaker constraint than Eq. 118. Note also that we assumed the plasma is nonrelativistic because our interest is in maximally delayed production; production at earlier times when electrons are relativistic is certainly allowed, but a different analysis would be necessary.
V.4.2 Matter power spectrum
Although the narrow resonance is centered on $k=0$, with time the resonance band only grows in width over comoving wave number since [51]
$\left|\left(\frac{k}{am_{\phi}}\right)^{2}+\delta\right|$ $<\frac{\eta\phi_{0}}{f}\left(\frac{H}{m_{\phi}}\right)^{3/4}.$ (131)
The dark photon’s power spectrum thus peaks at the largest wave number that is unstable when the scalar starts oscillating (at $H\approx m_{\phi}$). As long as the detuning $\delta$ is negligible, which is certainly the case during the early phases of the resonance, the peak wave number is approximately $k_{\star}/a_{\mathrm{osc}}\sim 2m_{A}\sqrt{\eta\phi_{0}/f}$. The density power spectrum is thus peaked as in Sec. V.3.4, with scales $\sim a_{\mathrm{eq}}/2k_{\star}$ collapsing at matter-radiation equality. The resulting substructures have a characteristic mass of
$M$ $\sim\frac{3}{2}M_{\mathrm{pl}}^{2}H_{\mathrm{eq}}^{2}\frac{4\pi}{3}\left(\frac{\pi}{2k_{\star}/a_{\mathrm{eq}}}\right)^{3}\sim 1.5\times 10^{-6}M_{\odot}\left(\frac{m_{A}}{10^{-12}\leavevmode\nobreak\ {\mathrm{eV}}}\right)^{-9/8}\left(\frac{H_{\star}}{10^{-22}\leavevmode\nobreak\ {\mathrm{eV}}}\right)^{-3/8},$ (132)
where we used Eq. 113 to solve for $\eta\phi_{0}/f$ and then used the relic abundance Eq. 114 to set a fiducial value for $\phi_{0}/m_{\phi}$ in terms of $m_{A}$ and $H_{\star}$ at the benchmark parameters. (The above result thus neglects logarithmic dependence on $m_{A}$.) While these structures are smaller than those produced by the kinetic coupling in the previous section, they may be interesting targets for photometric microlensing [212].
VI Discussion
In this paper, we mapped the parameter space of kinetically mixed dark photon dark matter, charting regions that require varying degrees of assumptions and model complexity to mechanize cosmologically consistent dark photon production. Building off the results of Ref. [48], in Sec. II we reviewed how early-Universe backreaction onto the dark Higgs impacts the viability of established dark photon production mechanisms. A fundamental assumption is that the dark photon’s mass is UV-completed by a dark Higgs mechanism, for which the dark gauge coupling and kinetic mixing would generally be of the same order. This minimal scenario faces stringent constraints due to the excitation of new degrees of freedom—namely, the dark Higgs, which disrupts production, and the formation of magnetic strings, whose presence rules out cold dark photon dark matter [48]. Our atlas of the Higgsed dark photon parameter space is summarized by Fig. 10; in the remainder of this section we review our main results.
Figure 10: Summary of various limits on viable kinetic mixing of dark photon dark matter. From right to left, blue lines correspond to the upper bounds for the various production scenarios outlined in the text: gravitational production during inflation (Sec. II.1); inflationary production assisted by a dilaton mechanizing an evolving dark photon mass (Sec. V.2); postinflationary production taking place at a fiducial Hubble rate $H_{\star}\approx m_{A}$, as exemplified by tachyonic production via axion oscillations (Sec. II.3); and maximally delayed postinflationary production, as realized by multiple models with rolling scalars (Secs. IV, V.3, and V.4). The red region above this last line cannot be realized by modified dark photon production, as the Proca theory breaks down even while modes observed in the CMB enter the horizon ($H_{\star}\lesssim 10^{-22}\leavevmode\nobreak\ {\mathrm{eV}}$); this parameter space instead requires a hierarchy in $\varepsilon$ and $g_{D}$ (see Sec. III). Above the red lines, defects nucleate due solely to the Standard Model plasma’s mixing with the dark photon, independent of whatever other mechanisms produce dark photons (see Sec. II.2); solid and dashed lines depict bounds assuming the minimum and maximum possible reheating temperatures within currently viable inflationary cosmologies. Signatures of the various nonminimal production scenarios become more prominent towards the upper boundaries of their allowed parameter space. For dilaton-assisted inflationary production, larger kinetic mixing requires larger inflationary scales and therefore could be corroborated by a detection of primordial gravitational waves with future CMB observations; this scenario also predicts equivalence principle violation. Dark matter substructure probed via astrometric lensing (darker blue) or photometric microlensing (green) could corroborate delayed production via, e.g., a runaway scalar (Fig. 9) with kinetic mixing above the fiducial postinflationary production line; the enhancement of small-scale structure could also be probed by future cosmological observations. Current exclusions from astrophysical [58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68] and cosmological [69, 70, 71, 72, 73] probes are depicted in dark gray and those from haloscope and other laboratory searches [76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 94, 95, 96, 97, 98, 99, 100, 101, 102, 103, 104, 105, 106] in light gray.
VI.1 Inflationary production
The minimal scenario of gravitational dark photon production during inflation [50] is the most strongly constrained [48]—no planned experiments come close to detecting viable kinetic mixing in this case. Sec. II.1 discussed the dependence of this result on the postinflationary expansion history, showing that a prolonged period of postinflationary kination (rather than instantaneous reheating) allows for kinetic mixings that are just marginally accessible to LZ [114] and XENON [83, 90, 94, 95] (see Fig. 4). A perhaps more plausible early-Universe cosmology—a prolonged matter era after inflation—leads instead to even more stringent upper limits.
In Sec. V we explored a class of models extending the Abelian-Higgs sector with a singlet scalar, whose couplings induce field-dependent modulation of the theory’s parameters. In such models, the scalar’s dynamics could directly alter the backreaction bound Eq. 6 throughout cosmological history—for instance, between inflation and the present day. In Sec. V.2 we showed that a universal coupling, for which the mass scales in the theory depend on the scalar, can expand the parameter space available to inflationary production if the dark photon’s mass drops after it becomes nonrelativistic. Further, we demonstrated that the dilaton model of Damour and Polyakov [53] provides a realization of such dynamics: if the dilaton is also universally coupled to Standard Model particles, it only starts rolling at late times when Standard Model particles begin to annihilate. Optimally weakening the backreaction bounds requires the inflationary scale to approach the upper limit permitted by CMB constraints, $H_{I}\sim 5\times 10^{13}\leavevmode\nobreak\ {\mathrm{GeV}}$ [56]. This “dilaton assist” model therefore predicts that if LZ detects a dark photon, future CMB experiments will detect primordial gravitational waves from inflation (see Figs. 10 and 7). Furthermore, the dilaton’s coupling to Standard Model matter potentially violates the equivalence principle and could be probed in future space-based experiments [226, 183].
VI.2 Postinflationary and delayed production
While gravitational production during inflation is inevitable, inflation need not occur at an energy scale high enough for inflationary dark photon production to fully account for the dark matter abundance. Postinflationary mechanisms, while less minimal, allow for dark photon dark matter not only over a much broader range of masses [54, 156, 159, 137] but also, as explained in Sec. II, a much broader range of kinetic mixing. The general bound of Eq. 8 demonstrates that backreaction on the dark Higgs and the formation of topological defects are evaded at larger couplings the later dark photon production occurs. Dark matter must exist in time for the standard $\Lambda$CDM model to describe the observed CMB; a reasonable benchmark is that production occurs no later than a Hubble rate of $H_{\star}\sim 10^{-22}\leavevmode\nobreak\ {\mathrm{eV}}$ (or redshifts around $10^{6}$). This bound is saturated along the “delayed production” contour in Fig. 10, and a signal near this boundary would indicate extremely delayed production, potentially leaving a signatures in the CMB at small angular scales.
In practice, however, mechanizing such delayed production is nontrivial, as demonstrated in the case of resonant production of dark photons from axion oscillations [54, 156, 157, 158] discussed in Sec. II.3. Kinematic requirements (for efficient parametric resonance) generally restrict production to early epochs when the Hubble rate exceeded the dark photon mass, setting the fiducial postinflationary benchmark with $H_{\star}\approx m_{A}$ in Fig. 10. A signal from experiments such as ALPHA [110], BREAD [111], Dark E-Field Radio [106], DM Radio [107], or MADMAX [110] could therefore point to production delayed beyond this minimal benchmark. In contrast, a signal from LAMPOST [112] or SuperCDMS [113] would be consistent with these standard postinflationary production scenarios.
In Sec. IV, we showed that a number of underappreciated regimes of axion production (see Fig. 5) are indeed viable in broader parameter space. The standard regime of broad tachyonic resonance requires dimensionless axion–dark-photon coupling $\beta\gtrsim 10$, which is parametrically larger than the typical expectation that $\beta\sim\alpha_{D}$. If axion oscillations are delayed by some means (such that $H_{\star}\ll m_{a}$), resonance becomes more efficient even at $\beta$ much smaller than unity (see Sec. IV.1 and also Ref. [158]). This regime of narrow resonance thus both increases the efficiency of production and decreases the degree of backreaction on the dark Higgs. Conversely, in Sec. IV.2 we describe a slow-roll regime at very large coupling ($\beta\gg 50$), which could delay production until $H_{\star}\sim m_{A}/\beta$ independent of the axion mass. Although these dynamics ultimately must be corroborated with numerical simulations (see Sec. D.2), it would offer a straightforward and predictive production mechanism for dark photons at parametrically late times. We also raised the possibility that if the axion couples to the dark photon in the magnetic basis, then the axion–dark-photon coupling may naturally be $\beta\sim\alpha_{D}^{-1}\gg 1$ [174, 227, 165], correlating the timing of production with the coupling strength: $H_{\star}\sim m_{A}\alpha_{D}$ (see Fig. 10). Although this scenario is rather speculative due to the theoretical difficulties outlined in Sec. IV.2, its potential simplicity and near-term detectability make it a compelling avenue for further study.
Finally, a scalar field responsible for the cosmological variation of the Abelian-Higgs parameters could directly produce dark photons. We considered two explicit scenarios: a rapidly varying dark gauge coupling driven by a runaway scalar [52] and narrow parametric resonance induced by an oscillating scalar [51]. Despite their distinct dynamics, both mechanisms can produce dark photons not just at early times when $H_{\star}\gtrsim m_{A}$ but also—at the cost of fine tuning (see Fig. 8)—at much later times when $H_{\star}\ll m_{A}$. In the case of a rapidly varying dark gauge coupling (Sec. V.3), the scalar not only sources a tachyonic resonance for the dark photon but also exponentially suppresses its bare mass, removing the kinematic barrier. The narrow resonance case (Sec. V.4), in contrast, relies on a unique instability that persists to arbitrarily late times provided that the dark-photon–scalar mass ratio is finely tuned to $1/2$.
While these scalar production scenarios must be tuned to provide an exception to the fiducial “postinflationary” line in Fig. 10, they both predict additional signatures that could test the models. Delayed production in both cases is accompanied by enhanced dark matter density perturbations on larger length scales, which could be probed by high-resolution CMB observations, the Lyman-$\alpha$ forest, line intensity mapping, and searches for dark matter substructure. Should vector dark matter be discovered at smaller masses and larger kinetic mixing, these complementary signatures could corroborate scalar couplings as the origin of the dark photon. The converse, however, does not hold: detecting small-scale structures or primordial gravitational waves (for dilaton-assisted inflationary production, discussed above) of a given size does not require the kinetic mixing to be as large as the corresponding defect-formation bounds permit.
VI.3 Model-angostic considerations
A model-independent bound on the kinetic mixing derives from the possibility that the hot Standard Model plasma restores symmetry to the dark Higgs at sufficiently high temperatures via the kinetic mixing portal (Sec. II.2). If symmetry is restored, strings later form through the Kibble mechanism, precluding stable, cold dark photon dark matter at late times and leading to the bound of Eq. 18. Although this bound is much weaker than the postinflationary production bound [Eq. 8] for the lowest possible reheating temperatures, it does imply that sufficiently light and strongly kinetically mixed dark photons (extending beyond the parameter space depicted in Fig. 10) are generically accompanied by a cosmic string network—a potential signature of an Abelian-Higgs dark sector that does not constitute the dark matter. Signals from dark matter detected by successors to SuperMAG [68], SNIPE [228], and AMAILS [229], which probe $\epsilon\gtrsim 10^{-8}$ at masses below $10^{-12}\leavevmode\nobreak\ {\mathrm{eV}}$, would require some nondynamical explanation—specifically, a kinetic mixing that is parametrically larger than the dark gauge coupling.
One possible resolution to this issue lies in the mechanism that generates kinetic mixing. Sec. II assumed that kinetic mixing arises from the exchange of heavy fermions charged under both electromagnetism and the dark gauge group, as originally proposed in Ref. [57]. Modifications to this scenario that allow the kinetic mixing to be parametrically larger than the dark gauge coupling could, in principle, circumvent the bounds discussed in Sec. II. Abelian clockwork is a well-established candidate mechanism to generate such a hierarchy [162]. In Sec. III, we assessed whether the additional degrees of freedom (i.e., Higgses) introduced by clockwork would impose additional constraints on the grounds of backreaction and defect formation [48]. Our analysis indicates that clockwork indeed provides a consistent means to boost kinetic mixing: as shown in Sec. III.1, the dark photon’s mixing with the added clockwork fields is suppressed by the already-small dark gauge coupling, so it couples no more strongly to any clockwork Higgs. Sec. III.2 also indicates that any production of the clockwork gauge fields induced by production of the dark photon itself imposes no stronger backreaction bounds.
While the clockwork mechanism does in principle provide an explanation for kinetic mixing larger than expected on backreaction grounds, such a proposal is rather ad hoc since the added field content plays no role in the origin of the dark matter other than to make it arbitrarily detectable. Dark photon production from axions is an exception: clockwork can simultaneously enhance the axion–dark-photon coupling and the kinetic mixing, and indeed is likely required to ensure the former is large enough for production to occur at all.
Finally, the dark photon’s mass could originate from the Stückelberg mechanism rather than the Higgs mechanism we focus on. Like a Higgs mass, a Stückelberg mass in known examples from string theory entails both radial and angular scalar degrees of freedom [203]. The angular field is absorbed by the dark photon, while the radial mode remains as a physical degree of freedom with a mass constrained similarly to the dark Higgs, i.e., smaller than about $4\pi m_{A}/g_{D}$. The phase of the Stückelberg field can also wind and form strings; in contrast to the Higgs case, the symmetric point where the dark photon mass vanishes lies infinitely far away in field space [203]. Even if defects do not form, backreaction onto radial Stückelberg modes presumably occurs at the same parametric threshold as in the Higgs case, suggesting that the low-energy Proca theory breaks down in similar parameter space. We leave a dedicated study of dark photon dark matter with Stückelberg masses to future work.
VII Conclusion
The direct detection parameter space for kinetically mixed dark photons has long been treated with a tacit assumption that there exists a consistent cosmology for dark photons of any mass and kinetic mixing—one that breaks down catastrophically for dark photons that acquire their mass through a (dark) Higgs mechanism. In such minimal scenarios, the magnitude of the kinetic mixing is determined by the dark gauge coupling, which also sets the regime of validity for the low-energy effective theory of a massive vector field. Breaching this threshold nucleates a string network that precludes cold dark photon dark matter [48]. In exploring the implications of these constraints, one perspective is to identify the most well-motivated regions of parameter space by examining concrete, minimal models of dark photon production. While the model that requires the fewest new degrees of freedom—gravitational production during inflation—is entirely undetectable, we showed that simple, UV-inspired physics can extend its parameter space to the reach of future electron recoil and noble gas detectors. Additional degrees of freedom, like rolling scalars, can also resonantly produce dark photons after inflation in parameter space that is broader but still inaccessible to most future haloscope experiments. We proposed a number of extensions thereof that, at the cost of some form of tuning, realize consistent dark photon production in all of the parameter space in reach of these searches.
An alternative approach studies the general implications of direct detection by any particular upcoming experiment, whose search space is already determined by its design. Many of these dark photon searches are simply byproducts of experiments specifically designed to search for axions; much kinetic mixing parameter space will therefore be probed at little added cost, regardless of theoretical motivation. Our analysis demonstrates that the full dark photon parameter space can indeed be made viable, either through dynamical mechanisms or more ad hoc model building such as clockwork. The reach of modified production dynamics, however, is limited by the requirement that dark photons behave like cold dark matter at observed cosmological epochs, imposing a model-independent threshold beyond which mechanisms like clockwork are required. Notably, this regime encompasses almost all existing cosmological constraints, exemplifying the importance of theoretical input to infer the nature of the dark sector from direct detection searches.
One motif of the various nonminimal models of dark photon production we studied is a correlation between laboratory measurements and cosmological and astrophysical signatures. Namely, the dynamics that allow for experimentally detectable kinetic mixing also predict unique phenomenology that could be probed by, for instance, future observations of the cosmic microwave background or searches for dark matter substructure. Ultralight dark photon dark matter scenarios also place strong requirements on cosmology more broadly. For the dilaton assist model discussed in Sec. V.2, the dark photon’s kinetic mixing is only detectable if the energy scale of inflation is large enough to be observed via gravitational waves in the CMB. On the other hand, even if dark photon dark matter originates by some other mechanism—as it must if lighter than an meV, regardless of its kinetic mixing—avoiding the nucleation of defects during inflation places a strong bound on its energy scale [Eq. 12]. Detectable, postinflationary production scenarios therefore generally require the energy scale of inflation be too low to be observed by future CMB experiments. By contrapositive, a detection thereof would place stringent limits on viable dark photon dark matter scenarios. The reheating temperature also cannot be too high, lest the Standard Model plasma restore the dark U(1) symmetry, seeding a string network via the Kibble mechanism.
Among the nonminimal models we studied, the scenarios with the most promising detection prospects frequently require some degree of parameter tuning. Nevertheless, the dynamical mechanisms we exploit—and our more model-agnostic analyses—provide a blueprint for pursuing viable constructions that may be more elegant or motivated. The potential for additional unique signatures, moreover, offers an avenue to falsify models or identify the physics underlying the dark sector. Then again, another reasonable response to the challenges in constructing dark photon models with large kinetic mixing is to pursue purely gravitational signatures of vector dark matter as a means of detection [230, 231, 232, 233, 234, 235, 236, 237, 238, 239, 240, 241]. A deeper, synergistic understanding of the interplay between particle models and cosmological dynamics stands to further advance the theoretical status of detectable dark photon dark matter.
Acknowledgements.
We are grateful to Peter Adshead, Isabel Garcia Garcia, Anson Hook, Justin Kaidi, Andrew Long, Matthew Reece, Katelin Schutz, Olivier Simon, Lorenzo Sorbo, and especially Masha Baryakhtar and Junwu Huang for discussions that have been invaluable in the preparation of this manuscript. D.C. is supported through the Department of Physics and College of Arts and Science at the University of Washington and by the U.S. Department of Energy Office of Science under Award Number DE-SC0024375. Z.J.W. is supported in part by the Department of Physics and College of Arts and Science at the University of Washington and the Dr. Ann Nelson Endowed Professorship. D.C. is grateful for the hospitality of Perimeter Institute where part of this work was carried out. Research at Perimeter Institute is supported in part by the Government of Canada through the Department of Innovation, Science and Economic Development Canada and by the Province of Ontario through the Ministry of Colleges and Universities. This material is partially supported by a grant from the Simons Foundation and the hospitality of the Aspen Center for Physics. This research was supported in part by grant NSF PHY-2309135 to the Kavli Institute for Theoretical Physics (KITP). Dark photon parameter space limits and projections are compiled in Ref. [47, 242].
Appendix A Scalar–Abelian-Higgs model
We first review the standard formulation of the spontaneously broken Abelian-Higgs model, extended to include arbitrary couplings to a scalar field. The classical Lagrangian describing a scalar–Abelian-Higgs theory is Eq. 55, quoted here:
$\mathcal{L}_{\mathrm{SAH}}$ $=\frac{1}{2}\partial_{\mu}\phi\partial^{\mu}\phi-V(\phi)-\frac{W(\phi)}{4}F_{\mu\nu}F^{\mu\nu}+\frac{X(\phi)}{2}D_{\mu}\Phi\left(D^{\mu}\Phi\right)^{\ast}-\frac{\lambda Y(\phi)}{4}\left({\left|\Phi\right|}^{2}-v^{2}\right)^{2}.$ (133)
The Higgs $\Phi$ is a complex scalar that we decompose into a phase and radial displacement from its vacuum expectation value, $\Phi=\left(v+h\right)e^{\mathrm{i}\Pi/v}$. The covariant derivative is $D_{\mu}\Phi=\partial_{\mu}\Phi-\mathrm{i}q_{\Phi}g_{D}A_{\mu}\Phi$, with $g_{D}$ the dark gauge coupling and $q_{\Phi}$ the (integer-valued) Higgs charge. In the broken phase and written in terms of radial fluctuations $h$ about the VEV and the Goldstone boson $\Pi$, Eq. 133 expands to
$\begin{split}\mathcal{L}_{\mathrm{SAH}}&=\frac{1}{2}\partial_{\mu}\phi\partial^{\mu}\phi-V(\phi)-\frac{W(\phi)}{4}F_{\mu\nu}F^{\mu\nu}+\frac{X(\phi)}{2}m_{A}^{2}A_{\mu}A^{\mu}\\ &\hphantom{{}={}}+\frac{X(\phi)}{2}\partial_{\mu}h\partial^{\mu}h-Y(\phi)\left(\frac{\lambda}{4}h^{4}+\lambda vh^{3}+\frac{1}{2}m_{h}^{2}h^{2}\right)\\ &\hphantom{{}={}}+\frac{X(\phi)}{2}\partial_{\mu}\Pi\partial^{\mu}\Pi+\frac{X(\phi)}{2}\left(\frac{h^{2}}{v^{2}}+\frac{2h}{v}\right)\left(\partial_{\mu}\Pi-m_{A}A_{\mu}\right)\left(\partial^{\mu}\Pi-m_{A}A^{\mu}\right),\end{split}$ (134)
where $m_{A}=q_{\Phi}g_{D}v$ and $m_{h}=\sqrt{2\lambda}v$.
We fix unitary gauge, in which $\Pi=0$. In a general spacetime, the Euler-Lagrange equation for the vector is
$-\nabla_{\mu}\left[W(\phi)F^{\mu\nu}\right]$ $=X(\phi)(1+h/v)^{2}m_{A}^{2}A^{\nu}\equiv\mathcal{X}(\phi,h)m_{A}^{2}A^{\nu}.$ (135)
The covariant divergence of thereof sets
00 0 $=\nabla^{\alpha}\left[\mathcal{X}(\phi,h)A_{\alpha}\right],$ (136)
which for $\mathcal{X}(\phi,h)=1$ reduces to the Lorenz “gauge” condition $\nabla^{\alpha}A_{\alpha}=0$ as applicable for massive vectors (i.e., Proca fields). The equations of motion are therefore
$\begin{split}0&=\nabla_{\mu}\nabla^{\mu}\phi+V^{\prime}(\phi)+\frac{W^{\prime}(\phi)}{4}F_{\mu\nu}F^{\mu\nu}-\frac{X^{\prime}(\phi)}{2}m_{A}^{2}A_{\mu}A^{\mu}\\ &\hphantom{{}={}}-\frac{X^{\prime}(\phi)}{2}\nabla_{\mu}h\nabla^{\mu}h+\frac{\lambda Y^{\prime}(\phi)}{4}\left(h^{2}+2hv\right)^{2}\end{split}$ (137a)
00 0 $=\nabla_{\mu}\nabla^{\mu}h+\frac{X^{\prime}(\phi)}{X(\phi)}\nabla_{\mu}\phi\nabla^{\mu}h+\frac{Y(\phi)}{X(\phi)}\left[\lambda h^{3}+3\lambda vh^{2}+m_{h}^{2}h\right]+\left(1+h/v\right)m_{A}^{2}A_{\mu}A^{\mu}/v$ (137b)
00 0 $=\nabla^{\alpha}\nabla_{\alpha}A_{\beta}-R_{\sigma\beta}A^{\sigma}+\nabla_{\nu}\left[A_{\mu}\nabla^{\mu}\ln\mathcal{X}(\phi,h)\right]+\frac{\nabla^{\alpha}W(\phi)}{W(\phi)}F_{\alpha\beta}+\frac{\mathcal{X}(\phi,h)}{W(\phi)}m_{A}^{2}A_{\beta},$ (137c)
where $R_{\sigma\beta}$ is the Ricci tensor and the Higgs mass is $m_{h}=\sqrt{2\lambda}v$.
As in the main text, we restrict our discussion to the FLRW spacetime with metric
$\mathrm{d}s^{2}$ $=\mathrm{d}t^{2}-a(t)^{2}\delta_{ij}\mathrm{d}x^{i}\mathrm{d}x^{j}$ (138)
and take the scalar to be homogeneous: $\phi(t,\mathbf{x})=\bar{\phi}(t)$. The form of Eq. 137 motivates absorbing the $\bar{\phi}$ dependence of the theory into its fundamental parameters as in Eq. 59. We expand the Higgs and vector components in Fourier space as
$h(t,\mathbf{x})$ $=\bar{h}(t)+\int\frac{\mathrm{d}^{3}k}{(2\pi)^{3}}\delta h(t,\mathbf{k})e^{\mathrm{i}\mathbf{k}\cdot\mathbf{x}}$ (139a)
$A_{0}(t,\mathbf{x})$ $=\int\frac{\mathrm{d}^{3}k}{(2\pi)^{3}}A_{0}(t,\mathbf{k})e^{\mathrm{i}\mathbf{k}\cdot\mathbf{x}}$ (139b)
$A_{i}(t,\mathbf{x})$ $=\sum_{\lambda}^{\{\pm,\parallel\}}\int\frac{\mathrm{d}^{3}k}{(2\pi)^{3}}A_{\lambda}(t,\mathbf{k})\varepsilon^{\lambda}_{i}(\mathbf{k})e^{\mathrm{i}\mathbf{k}\cdot\mathbf{x}},$ (139c)
where the polarization vectors $\bm{\varepsilon}^{\pm}(\mathbf{k})$ and $\bm{\varepsilon}^{\parallel}(\mathbf{k})$ form an orthogonal basis of polarizations [i.e., $\varepsilon^{\lambda}_{m}(\mathbf{k})\varepsilon^{\lambda^{\prime}}_{m}(\mathbf{k})^{\ast}=\delta^{\lambda\lambda^{\prime}}$] that are respectively transverse [$k_{m}\varepsilon^{\pm}_{m}(\mathbf{k})$] and longitudinal [$\mathrm{i}k_{m}\varepsilon^{\parallel}_{m}(\mathbf{k})=k$] to the wave number $\mathbf{k}$. They are Hermitian [$\varepsilon^{\lambda}_{m}(-\mathbf{k})=\varepsilon^{\lambda}_{m}(\mathbf{k})^{\ast}$] and the circular polarization vectors satisfy $\varepsilon^{\pm}_{m}(\mathbf{k})^{\ast}=\varepsilon^{\mp}_{m}(\mathbf{k})$. The linearized equations of motion decomposed onto this basis (and written in terms of the $\bar{\phi}$-dependent parameters) evaluate to
00 0 $=\ddot{\bar{\phi}}+3H\dot{\bar{\phi}}+V^{\prime}(\bar{\phi})$ (140a)
00 0 $=\ddot{\bar{h}}+\left[3H+\frac{\partial_{t}X(\bar{\phi})}{X(\bar{\phi})}\right]\dot{\bar{h}}+\left[X(\bar{\phi})\lambda(\bar{\phi})\bar{h}^{2}+3\sqrt{X(\bar{\phi})}\lambda(\bar{\phi})v(\bar{\phi})\bar{h}+m_{h}(\bar{\phi})^{2}\right]h$ (140b)
00 0 $=\delta\ddot{h}+\left[3H+\frac{\partial_{t}X(\bar{\phi})}{X(\bar{\phi})}\right]\delta\dot{h}+\left[\frac{k^{2}}{a^{2}}+3X(\bar{\phi})\lambda(\bar{\phi})\bar{h}^{2}+6\sqrt{X(\bar{\phi})}\lambda(\bar{\phi})v(\bar{\phi})\bar{h}+m_{h}(\bar{\phi})^{2}\right]\delta h$ (140c)
00 0 $=\ddot{A}_{\pm}+\left[H+\frac{\partial_{t}W(\bar{\phi})}{W(\bar{\phi})}\right]\dot{A}_{\pm}+\left[\frac{k^{2}}{a^{2}}+m_{A}(\bar{\phi},\bar{h})^{2}\right]A_{\pm}$ (140d)
$\begin{split}0&=\ddot{A}_{\parallel}+\left[H+\frac{\partial_{t}W(\bar{\phi})}{W(\bar{\phi})}\right]\dot{A}_{\parallel}+\left[\frac{k^{2}}{a^{2}}+m_{A}(\bar{\phi},\bar{h})^{2}\right]A_{\parallel}-\left[2H+\frac{\partial_{t}\mathcal{X}(\bar{\phi},\bar{h})}{\mathcal{X}(\bar{\phi},\bar{h})}-\frac{\partial_{t}W(\bar{\phi})}{W(\bar{\phi})}\right]kA_{0}.\end{split}$ (140e)
Here we use the shorthand $m_{A}(\phi,h)=m_{A}(\phi)(1+h/v)^{2}$. In addition, Eq. 136 reduces to
00 0 $=-\dot{A}_{0}-\left(3H+\frac{\partial_{t}\mathcal{X}(\bar{\phi},\bar{h})}{\mathcal{X}(\bar{\phi},\bar{h})}\right)A_{0}+\frac{\partial_{i}A_{i}}{a^{2}}$ (141)
and the $\nu=0$ component of Eq. 135 (i.e., Gauss’s law) reads
00 0 $=k\dot{A}_{\parallel}+\left[k^{2}+a^{2}m_{A}(\bar{\phi},\bar{h})\right]A_{0};$ (142)
solving Eq. 142 for $A_{0}$ and substituting into Eq. 140e yields
00 0 $=\ddot{A}_{\parallel}+\frac{\left[3H+\partial_{t}\ln\mathcal{X}(\bar{\phi},\bar{h})\right]k^{2}+\left[H+\partial_{t}\ln W(\bar{\phi})\right]a^{2}m_{A}(\bar{\phi},\bar{h})^{2}}{k^{2}+a^{2}m_{A}(\bar{\phi},\bar{h})^{2}}\dot{A}_{\parallel}+\left[\frac{k^{2}}{a^{2}}+m_{A}(\bar{\phi},\bar{h})^{2}\right]A_{\parallel}.$ (143)
For the rescaled fields $\mathfrak{h}=\sqrt{X(\bar{\phi})}h$ and $\mathcal{A}_{\pm}=\sqrt{W(\bar{\phi})}A_{\pm}$, the equations of motion reduce to
00 0 $=\ddot{\bar{\mathfrak{h}}}+3H\dot{\bar{\mathfrak{h}}}+\left[\lambda(\bar{\phi})\bar{\mathfrak{h}}^{2}+3\lambda(\bar{\phi})v(\bar{\phi})\bar{\mathfrak{h}}+m_{h}(\bar{\phi})^{2}-\frac{3H}{2}\frac{\partial_{t}X(\bar{\phi})}{X(\bar{\phi})}-\frac{\partial_{t}^{2}\sqrt{X(\bar{\phi})}}{\sqrt{X(\bar{\phi})}}\right]\bar{\mathfrak{h}}$ (144)
00 0 $=\delta\ddot{\mathfrak{h}}+3H\delta\dot{\mathfrak{h}}+\left[\frac{k^{2}}{a^{2}}+3\lambda(\bar{\phi})\bar{\mathfrak{h}}^{2}+6\lambda(\bar{\phi})v(\bar{\phi})\bar{\mathfrak{h}}+m_{h}(\bar{\phi})^{2}-\frac{3H}{2}\frac{\partial_{t}X(\bar{\phi})}{X(\bar{\phi})}-\frac{\partial_{t}^{2}\sqrt{X(\bar{\phi})}}{\sqrt{X(\bar{\phi})}}\right]\delta\mathfrak{h}$ (145)
00 0 $=\ddot{\mathcal{A}}_{\pm}+H\dot{\mathcal{A}}_{\pm}+\left[\frac{k^{2}}{a^{2}}+m_{A}(\bar{\phi},\bar{h})^{2}-\frac{H}{2}\frac{\partial_{t}W(\bar{\phi})}{W(\bar{\phi})}-\frac{\partial_{t}^{2}\sqrt{W(\bar{\phi})}}{\sqrt{W(\bar{\phi})}}\right]\mathcal{A}_{\pm}.$ (146)
Observe that the only couplings to $\bar{\phi}$ not embedded in the parameters defined in Eq. 59 are derivative couplings.
The stress-energy tensor for Eq. 134 (in unitary gauge) is
$\begin{split}T^{\mu}_{\hphantom{\mu}\nu}&=W(\phi)\left[F^{\mu\alpha}F_{\nu\alpha}-m_{A}(\phi,h)^{2}A^{\mu}A_{\nu}+\delta^{\mu}_{\hphantom{\mu}\nu}\left(-\frac{1}{4}F_{\alpha\beta}F^{\alpha\beta}+\frac{1}{2}m_{A}(\phi,h)^{2}A_{\alpha}A^{\alpha}\right)\right]\\ &\hphantom{{}={}}+X(\phi)\left[-\partial^{\mu}h\partial_{\nu}h+\delta^{\mu}_{\hphantom{\mu}\nu}\left(\frac{1}{2}\partial_{\alpha}h\partial^{\alpha}h-\frac{Y(\phi)}{X(\phi)}\left[\frac{\lambda}{4}h^{4}+\lambda vh^{3}+\frac{1}{2}m_{h}^{2}h^{2}\right]\right)\right]\\ &\hphantom{{}={}}-\partial^{\mu}\phi\partial_{\nu}\phi+\delta^{\mu}_{\hphantom{\mu}\nu}\left(\frac{1}{2}\partial^{\alpha}\phi\partial_{\alpha}\phi-V(\phi)\right).\end{split}$ (147)
Using Eq. 142 to solve for $A_{0}$, we may express the dark photon’s average energy density $\bar{\rho}_{A}$ and pressure $\bar{P}_{A}$ in terms of the polarization decomposition Eq. 139c. To the same level of approximation as Eq. 140 (and assuming isotropic initial conditions), the contributions from the transverse and longitudinal components to the energy density $\bar{\rho}_{A}$ are respectively
$\bar{\rho}_{\pm}(t)$ $=\frac{W(\bar{\phi})}{4\pi^{2}a^{2}}\int\mathrm{d}k\,k^{2}\left[{\left|\dot{A}_{\pm}\right|}^{2}+\frac{k^{2}}{a^{2}}{\left|A_{\pm}\right|}^{2}+m_{A}(\bar{\phi},\bar{h})^{2}{\left|A_{\pm}\right|}^{2}\right]$ (148a)
and
$\bar{\rho}_{\parallel}(t)$ $=\frac{W(\bar{\phi})m_{A}(\bar{\phi},\bar{h})^{2}}{4\pi^{2}a^{2}}\int\mathrm{d}k\,k^{2}\left[\frac{1}{k^{2}+a^{2}m_{A}(\bar{\phi},\bar{h})^{2}}{\left|a\dot{A}_{\parallel}\right|}^{2}+{\left|A_{\parallel}\right|}^{2}\right],$ (148b)
and those to the pressure $P_{A}$ are
$\bar{P}_{\pm}(t)$ $=\frac{W(\bar{\phi})}{12\pi^{2}a^{2}}\int\mathrm{d}k\,k^{2}\left[{\left|\dot{A}_{\pm}(\tau,k)\right|}^{2}+\frac{k^{2}}{a^{2}}{\left|A_{\pm}(\tau,k)\right|}^{2}-m_{A}(\bar{\phi},\bar{h})^{2}{\left|A_{\pm}(\tau,k)\right|}^{2}\right]$ (149a)
and
$\bar{P}_{\parallel}(t)$ $=\frac{W(\bar{\phi})m_{A}(\bar{\phi},\bar{h})^{2}}{12\pi^{2}a^{2}}\int\mathrm{d}k\,k^{2}\left[\frac{3k^{2}+a^{2}m_{A}(\bar{\phi},\bar{h})^{2}}{\left[k^{2}+a^{2}m_{A}(\bar{\phi},\bar{h})^{2}\right]^{2}}{\left|a\dot{A}_{\parallel}(\tau,k)\right|}^{2}-{\left|A_{\parallel}(\tau,k)\right|}^{2}\right].$ (149b)
Appendix B Quantum corrections
We seek to compute the one-loop quantum corrections to the theory of a scalar coupled to the standard Abelian-Higgs model, Eq. 55. The quantization of theories with noncanonical kinetic terms entails subtleties whose neglect leads to incorrect scattering amplitudes and can, in our case, severely alter the ultraviolet (UV) divergences of radiative corrections. We first review the appropriate formulation before proceeding to compute one-loop effects in the scalar–Abelian-Higgs model.
B.1 Quantization of noncanonical field theories
To compute quantum corrections, we require the Hamiltonian $\mathcal{H}$ (specifically as a functional of fields and their four-gradients and not the conjugate momenta as well) corresponding to Eq. 134, which generates the time evolution and interactions of the theory. In classical field theory, the Hamiltonian and Lagrangian are related by a Legendre transform, and the only complication in theories with noncanonical kinetic terms is that fields’ conjugate momenta are not trivially related to their time derivatives. In quantum field theory, the appropriate Hamiltonian coincides with the Legendre transform of the Lagrangian only if the kinetic terms are canonical.
Deriving the effect of kinetic couplings is most straightforward in the path-integral formalism (see Ref. [243] section 9.3, whose notation we partly adopt below), in which the generating functional for a general multiplet of bosons $\phi_{a}$ with conjugate momenta $\pi_{a}$ is
$\mathcal{Z}$ $=\int\prod_{a}\left[\mathcal{D}\phi_{a}\frac{\mathcal{D}\pi_{a}}{2\pi}\right]\exp\left\{\frac{\mathrm{i}}{\hbar}\int\mathrm{d}^{4}x\,\left[\sum_{a}\dot{\phi}_{a}(x)\pi_{a}(x)-\mathcal{H}[\phi_{a}(x),\pi_{a}(x)]\right]\right\}.$ (150)
For ease of illustration and generality, parametrize the Hamiltonian as
$\mathcal{H}[\phi_{a},\pi_{a}]$ $=\frac{1}{2}\sum_{a,b}\pi_{a}A_{ab}[\phi]\pi_{b}+\sum_{a}B_{a}[\phi]\pi_{a}+C[\phi].$ (151)
In a canonical theory, $A_{ab}[\phi]=\delta_{ab}$. The functional $C[\phi]$ includes (spatial) gradient and potential terms for a scalar theory and mass, current-coupling, and magnetic-field terms for a gauge theory. The salient effects we aim to illustrate depend only on $A_{ab}[\phi]$; we retain a general parametrization for succinctness and provide an explicit mapping of Eq. 151 to the scalar–Abelian-Higgs theory later on.
Because the kinetic term in Eq. 151 is still quadratic in the conjugate momenta, the path integral over $\pi_{a}$ may be computed exactly. The argument of the exponential in the path integral indeed coincides with the (inverse) Legendre transform of the Hamiltonian Eq. 151, as a consequence of the integration over $\pi_{a}$ instead of Hamilton’s equations [243]:
$\mathcal{Z}$ $=\int\prod_{a}\left[\mathcal{D}\phi_{a}\right]\det\left(2\pi\mathrm{i}A[\phi]/\hbar\right)^{-1/2}\exp\left\{\frac{\mathrm{i}}{\hbar}\int\mathrm{d}^{4}x\,\mathcal{L}[\phi,\dot{\phi}]\right\},$ (152)
where
$\mathcal{L}[\phi_{a},\dot{\phi}_{a}]$ $=\sum_{a,b}\left[\frac{1}{2}\dot{\phi}_{a}A[\phi]^{-1}_{ab}\dot{\phi}_{b}+\frac{1}{2}B_{a}[\phi]A[\phi]^{-1}_{ab}B_{b}[\phi]-B_{a}[\phi]A[\phi]^{-1}_{ab}\dot{\phi}_{b}\right]-C[\phi].$ (153)
While Eq. 153 is superficially identical to the corresponding classical result, any field dependence in the determinant factor of Eq. 152 (via the kinetic term in $A[\phi]$) does modify the path integral, effectively introducing to the Lagrangian an additional term
$\Delta\mathcal{L}$ $=\frac{\mathrm{i}\hbar}{2}\int\frac{\mathrm{d}^{4}k}{(2\pi)^{4}}\operatorname{tr}\ln A[\phi].$ (154)
Instead of performing the path integral over $\pi_{a}$, we could shift the conjugate momentum to
$\pi_{a}^{\prime}$ $=\pi_{a}-A[\phi]^{-1}_{ab}\left(B_{b}[\phi]-\dot{\phi}_{b}\right)$ (155)
and retain it as an auxiliary degree of freedom in the Lagrangian that interacts with $\phi$:
$\mathcal{L}[\phi_{a},\dot{\phi}_{a},\pi_{a}^{\prime}]$ $=-\frac{1}{2}\sum_{a,b}A_{ab}[\phi]\pi_{a}^{\prime}\pi_{b}^{\prime}+\mathcal{L}[\phi_{a},\dot{\phi}_{a}].$ (156)
Though $\pi_{a}^{\prime}$ is nonpropagating, it modifies correlation functions of $\phi_{a}$ through loops. That the leading contributions from the new terms in the Lagrangian should be counted at one-loop level is less obvious in the form Eq. 154, but is affirmed by the relative factor of $\hbar$ acquired by $\Delta\mathcal{L}$ when the determinant factor is reabsorbed into the Lagrangian.181818 Related observations were made in previous studies of noncanonical field theories: the first exposition on the additional interactions arising in noncanonical field theories [244] found in chiral perturbation theory that their contribution to the pion propagator appears at next-to-leading order in the pion decay constant, and when considering loop corrections during inflation their contribution is suppressed in the slow-roll expansion [245, 246].
We now map the preceding general results to a massive vector field kinetically coupled to a scalar field. We replace $\phi_{a}\to A_{i}$ and $\pi_{a}\to E_{i}$. Recovering the covariant Lagrangian
$\mathcal{L}$ $=-\frac{W}{4}F_{\mu\nu}F^{\mu\nu}+\frac{1}{2}m_{A}^{2}A_{\mu}A^{\mu}+J_{\mu}A^{\mu}$ (157)
(with $J_{\mu}$ standing in for the Higgs interactions) requires choosing
$A_{ij}$ $=W^{-1}\delta_{ij}$ (158a)
$B_{i}$ $=\partial_{i}A_{0}$ (158b)
$C$ $=\frac{1}{2}m_{A}^{2}\left(A_{i}\right)^{2}+\frac{W}{2}\left(\epsilon^{ijk}\partial_{j}A_{k}\right)^{2}+J_{i}A_{i}$ (158c)
in Eq. 151 and, just as for a canonical vector field, introducing $A_{0}$ as an auxiliary field (also integrated over in $\mathcal{Z}$) and adding to the Hamiltonian
$\Delta\mathcal{H}$ $=-\frac{1}{2}m_{A}^{2}\left(A_{0}-\frac{\partial_{i}E_{i}-J_{0}}{m_{A}^{2}}\right)^{2}.$ (159)
The Higgs and the Goldstone, whose kinetic terms in the Lagrangian are multiplied by $X$, each contribute a copy of Eq. 154 with $A$ set to $X^{-1}$. The effective Lagrangian therefore includes the contribution
$\Delta\mathcal{L}$ $=-\frac{3\mathrm{i}\hbar}{2}\int\frac{\mathrm{d}^{4}k}{(2\pi)^{4}}\ln W-\mathrm{i}\hbar\int\frac{\mathrm{d}^{4}k}{(2\pi)^{4}}\ln X$ (160)
in addition to the classical Lagrangian.
B.2 Background-field expansion
With the full quantum theory in tow, we may compute radiative corrections. We regulate with a finite momentum cutoff $\Lambda$ under the usual interpretation that it corresponds to the scale of new physics in a given UV completion of the effective theory described by Eq. 134. We seek to assess what limitations on $\Lambda$ are imposed to ensure consistency of the scalar production scenario—principally, avoiding large corrections to the scalar’s classical potential. Since we consider scalar field configurations with classical homogeneous components, we perform a standard background-field expansion to compute the effective potential due to quantum fluctuations.
We first expand $\phi\to f\varphi_{b}+\phi$ in Eq. 134, where $\varphi_{b}=\phi_{b}/f$ is treated as a spacetime constant. We then drop constant terms and tadpoles (i.e., terms with fewer than two powers of the fluctuating fields). We gauge fix according to the standard Faddeev-Popov procedure, choosing the gauge fixing functional
$G(A_{\mu},\Pi)$ $=\sqrt{X(\phi)}\left(\partial_{\mu}A^{\mu}-\xi m_{A}\Pi\right);$ (161)
the factor of $\sqrt{X(\phi)}$ introduced relative to the textbook choice ensures that terms that mix $A_{\mu}$ and $\Pi$ in the quadratic action still cancel when $-G(A_{\mu},\Pi)^{2}/2\xi$ is added to the action. The Faddeev-Popov determinant also picks up this factor of $\sqrt{X(\phi)}$, which we account for by including it in the action for the ghosts $c$ and $\bar{c}$. Gauge fixing then adds to the Lagrangian
$\mathcal{L}_{\mathrm{FP}}$ $=-\frac{X(\phi)}{2\xi}\left(\partial_{\mu}A^{\mu}-\xi m_{A}\Pi\right)^{2}+\sqrt{X(\phi)}\left(\partial_{\mu}\bar{c}\partial^{\mu}c-\xi m_{A}^{2}c\bar{c}\right).$ (162)
Finally, we canonically normalize191919 “Canonically normalize” in the sense that propagators take their standard form—no interactions are altered by doing so, only the point at which we rearrange factors of $W(\varphi_{b})$ and $X(\varphi_{b})$ to write results in canonical form. fields by replacing $A_{\mu}$ with $A_{\mu}/\sqrt{W(\varphi_{b})}$, $h$ with $h/\sqrt{X(\varphi_{b})}$, $\Pi$ with $\Pi/\sqrt{X(\varphi_{b})}$, and $c$ with $c/\sqrt[4]{X(\varphi_{b})}$. To compute the effective potential we require only the quadratic action (including background-field dependence in effective masses). The part of $\mathcal{L}_{\mathrm{SAH}}+\mathcal{L}_{\mathrm{FP}}$ quadratic in fields is
$\mathcal{L}_{(2)}$ $=-\frac{1}{4}F_{\mu\nu}F^{\mu\nu}-\frac{1}{2\xi(\varphi_{b})}\left(\partial_{\mu}A^{\mu}\right)^{2}+\frac{1}{2}m_{A}(\varphi_{b})^{2}A_{\mu}A^{\mu}+\frac{1}{2}\partial_{\mu}\phi\partial^{\mu}\phi-\frac{1}{2}m(\varphi_{b})^{2}\phi^{2}$ (163)
[…]
where $m_{A}(\varphi_{b})$ and $m_{h}(\varphi_{b})$ are defined in analogy to Eq. 59 and
$m(\varphi_{b})^{2}$ $\equiv V^{\prime\prime}(f\varphi_{b})/f^{2}$ (165)
$\xi(\varphi_{b})$ $\equiv\xi W(\varphi_{b})/X(\varphi_{b}).$ (166)
For completeness, we also report the interaction Lagrangian (in unitary gauge, $\xi\to\infty$, so that we can neglect the Goldstone boson and ghosts), as would be needed to compute radiative corrections:
$\begin{split}\mathcal{L}_{\mathrm{int}}&=-\frac{1}{4}F_{\mu\nu}F^{\mu\nu}\sum_{n=1}^{\infty}\frac{1}{n!}\frac{W^{(n)}(\varphi_{b})}{W(\varphi_{b})}\frac{\phi^{n}}{f^{n}}+\left[\frac{1}{2}\partial_{\mu}h\partial^{\mu}h+\frac{1}{2}m_{A}(\varphi_{b})^{2}A_{\mu}A^{\mu}\right]\sum_{n=1}^{\infty}\frac{1}{n!}\frac{X^{(n)}(\varphi_{b})}{X(\varphi_{b})}\frac{\phi^{n}}{f^{n}}\\ &\hphantom{{}={}}+m_{A}(\varphi_{b})^{2}\left(\frac{h}{v(\varphi_{b})}+\frac{h^{2}}{2v(\varphi_{b})^{2}}\right)A_{\mu}A^{\mu}\sum_{n=0}^{\infty}\frac{1}{n!}\frac{X^{(n)}(\varphi_{b})}{X(\varphi_{b})}\frac{\phi^{n}}{f^{n}}-\sum_{n=3}^{\infty}\frac{\tilde{V}^{(n)}(\varphi_{b})}{n!f^{n}}\phi^{n}\\ &\hphantom{{}={}}-\frac{1}{2}m_{h}(\varphi_{b})^{2}h^{2}\sum_{n=1}^{\infty}\frac{1}{n!}\frac{Y^{(n)}(\varphi_{b})}{Y(\varphi_{b})}\frac{\phi^{n}}{f^{n}}-\left[\frac{\lambda(\varphi_{b})}{4}h^{4}+\lambda(\varphi_{b})v(\varphi_{b})h^{3}\right]\sum_{n=0}^{\infty}\frac{1}{n!}\frac{Y^{(n)}(\varphi_{b})}{Y(\varphi_{b})}\frac{\phi^{n}}{f^{n}}.\end{split}$ (167)
The additional interactions incurred in the quantum theory [Eq. 160] expand to
$\Delta\mathcal{L}$ $=-\frac{\mathrm{i}}{2}\int\frac{\mathrm{d}^{4}k}{(2\pi)^{4}}\sum_{n=0}^{\infty}\frac{1}{n!}\left(\frac{\phi}{f}\right)^{n}\left[3\frac{\partial^{n}\ln W(\varphi)}{\partial\varphi^{n}}+2\frac{\partial^{n}\ln X(\varphi)}{\partial\varphi^{n}}\right]_{\varphi=\varphi_{b}}.$ (168)
Both the $\phi$-independent term and the quadratic term are required explicitly for our purposes; they are
$\Delta\mathcal{L}_{(0)}$ $=-\frac{\mathrm{i}}{2}\int\frac{\mathrm{d}^{4}k}{(2\pi)^{4}}\left[3\ln W(\varphi_{b})+2\ln X(\varphi_{b})\right]$ (169)
$\begin{split}\Delta\mathcal{L}_{(2)}&=-\frac{3\mathrm{i}}{4f^{2}}\int\frac{\mathrm{d}^{4}k}{(2\pi)^{4}}\left[\frac{W^{(2)}(\varphi_{b})}{W(\varphi_{b})}-\left(\frac{W^{(1)}(\varphi_{b})}{W(\varphi_{b})}\right)^{2}\right]\phi^{2}.\\ &\hphantom{{}={}}-\frac{\mathrm{i}}{2f^{2}}\int\frac{\mathrm{d}^{4}k}{(2\pi)^{4}}\left[\frac{X^{(2)}(\varphi_{b})}{X(\varphi_{b})}-\left(\frac{X^{(1)}(\varphi_{b})}{X(\varphi_{b})}\right)^{2}\right]\phi^{2}.\end{split}$ (170)
The coefficient of $-\phi^{2}/2$ is an effective one-loop contribution to the scalar’s squared mass. In practice, Eqs. 169 and 170 act like counterterms to photon and Higgs loops and cancel leading divergences.
B.3 Effective potential
The one loop effective potential may be computed by integrating the quadratic action over the fields, Eq. 164 [247]. The contribution from scalar fluctuations is
$e^{\mathrm{i}\Delta\Gamma[\phi_{b}]^{(1),\phi}}$ $=\int\mathcal{D}\phi\exp\left\{\mathrm{i}\int\mathrm{d}^{4}x\,\left[\frac{1}{2}\partial_{\mu}\phi\partial^{\mu}\phi-\frac{1}{2}m(\varphi_{b})^{2}\phi^{2}\right]\right\},$ (171)
which, integrated over $\phi$, adds to the effective action
$\mathrm{i}\Delta\Gamma[\phi_{b}]^{(1),\phi}$ $=-\frac{1}{2}\int\mathrm{d}^{4}x\int\frac{\mathrm{d}^{4}k}{(2\pi)^{4}}\ln\left(1-\frac{m(\varphi_{b})^{2}}{k^{2}}\right).$ (172)
At next-to-leading order in $\Lambda\gg m(\varphi_{b})$, the resulting effective potential is
$\Delta V[\phi_{b}]^{(1),\phi}$ $=\frac{\Lambda^{2}m(\varphi_{b})^{2}}{2(4\pi)^{2}}-\frac{m(\varphi_{b})^{4}}{8(4\pi)^{2}}\left[1+2\ln\left(1+\frac{\Lambda^{2}}{m(\varphi_{b})^{2}}\right)\right],$ (173)
which is subdominant to the bare potential when $\Lambda^{2}/2(4\pi f)^{2}<1$. In the special case that $V(\varphi_{b})$ is a pure exponential function, the quantum corrections to the scalar effective potential arising from scalar loops organize into an expansion in the number of vertices [196]. At one-vertex level, the corrections are directly proportional to $V(\varphi_{b})$ and can therefore be absorbed into the overall normalization of the potential.
The Higgs contribution is
$e^{\mathrm{i}\Delta\Gamma[\phi_{b}]^{(1),h}}$ $=\int\frac{\mathcal{D}h}{\sqrt{X(\varphi_{b})}}\exp\left\{\mathrm{i}\int\mathrm{d}^{4}x\,\left[\frac{1}{2}\partial_{\mu}h\partial^{\mu}h-\frac{1}{2}m_{h}(\varphi_{b})^{2}h^{2}\right]\right\}$ (174)
Observe that the measure is rescaled in accordance with the field redefinition made previously. The resulting effective potential contribution is
$\Delta V[\phi_{b}]^{(1),h}$ $=-\frac{\mathrm{i}}{2}\int\frac{\mathrm{d}^{4}k}{(2\pi)^{4}}\left[\ln X(\varphi_{b})+\ln\left(1-\frac{m_{h}(\varphi_{b})^{2}}{k^{2}}\right)\right].$ (175)
The calculation for the Goldstone proceeds identically:
$\Delta V[\phi_{b}]^{(1),\Pi}$ $=-\frac{\mathrm{i}}{2}\int\frac{\mathrm{d}^{4}k}{(2\pi)^{4}}\left[\ln X(\varphi_{b})+\ln\left(1-\frac{\xi m_{A}^{2}}{k^{2}}\right)\right].$ (176)
Recall that the ghosts were rescaled by a factor of $4]{X(\varphi_{b})}$; however, as Grassman numbers their measure changes by the inverse of the applied scaling transformation. Therefore,
$e^{\mathrm{i}\Delta\Gamma[\phi_{b}]^{(1),c}}$ $=\int\mathcal{D}c\mathcal{D}\bar{c}\,\sqrt{X(\varphi_{b})}\exp\left\{\mathrm{i}\int\mathrm{d}^{4}x\,\left(\partial_{\mu}\bar{c}\partial^{\mu}c-\xi m_{A}^{2}c\bar{c}\right)\right\},$ (177)
so
$\Delta V[\phi_{b}]^{(1),c}$ $=\frac{\mathrm{i}}{2}\int\frac{\mathrm{d}^{4}k}{(2\pi)^{4}}\left[\ln X(\varphi_{b})+2\ln\left(1-\frac{\xi m_{A}^{2}}{k^{2}}\right)\right].$ (178)
Finally, the quadratic action for the photon may be written as
$e^{\mathrm{i}\Delta\Gamma[\phi_{b}]^{(1),A}}$ $=\int\frac{\mathcal{D}A_{\mu}}{W(\varphi_{b})^{2}}\exp\left\{-\frac{\mathrm{i}}{2}\int\mathrm{d}^{4}x\,A^{\mu}\Delta^{-1}_{A,\mu\nu}A^{\nu}\right\}=W(\varphi_{b})^{-2}\det\left[\Delta_{A}^{-1}\right]^{-1/2}$ (179)
where
$\Delta^{-1}_{A,\mu\nu}$ $\equiv-g^{\mu\nu}\left[\partial_{\alpha}\partial^{\alpha}+m_{A}(\varphi_{b})^{2}\right]+\left(1-\frac{1}{\xi(\varphi_{b})}\right)\partial^{\mu}\partial^{\nu}.$ (180)
The functional determinant may be computed in Euclidean time ($\tau=t/\mathrm{i}$) as an integral over its (Fourier-space) eigenvalues,
$\ln\det\Delta^{-1}_{A,\mu\nu}$ $=\mathrm{i}\int\frac{\mathrm{d}^{4}k_{E}}{(2\pi)^{4}}\operatorname{tr}\ln\left\{\left[-k_{E}^{2}-m_{A}(\varphi_{b})^{2}\right]\left[-\delta^{\mu\nu}+\left(1-\frac{1}{\xi(\varphi_{b})}\right)\frac{k_{E}^{\mu}k_{E}^{\nu}}{k_{E}^{2}+m_{A}(\varphi_{b})^{2}}\right]\right\}.$ (181)
Series expanding the matrix logarithm, simplifying each term, resumming, and shuffling constant factors (which are dropped from the effective potential) yields
$\ln\det\Delta^{-1}_{A,\mu\nu}$ $=\mathrm{i}\int\frac{\mathrm{d}^{4}k_{E}}{(2\pi)^{4}}\left\{3\ln\left(1+\frac{m_{A}(\varphi_{b})^{2}}{k_{E}^{2}}\right)+\ln\left(1+\frac{\xi m_{A}^{2}}{k_{E}^{2}}\right)-\ln\xi(\varphi_{b})\right\}.$ (182)
The first term is the contribution from the three physical dark photon polarizations. Substituting Eq. 166 to compute $\Delta V[\phi_{b}]^{(1),A}$ via Eq. 179 and summing with Eqs. 175, 176, and 178 yields a net contribution from the Abelian-Higgs sector of
$\Delta V[\phi_{b}]^{(1)}$ $=\frac{1}{2}\int\frac{\mathrm{d}^{4}k_{E}}{(2\pi)^{4}}\left[3\ln\left(1+\frac{m_{A}(\varphi_{b})^{2}}{k_{E}^{2}}\right)+\ln\left(1+\frac{m_{h}(\varphi_{b})^{2}}{k_{E}^{2}}\right)+3\ln W(\varphi_{b})+2\ln X(\varphi_{b})\right].$ (183)
Observe that the latter two terms, which diverge quartically, precisely cancel the contributions incurred from quantizing the theory, $\Delta\mathcal{L}_{(0)}$ [Eq. 169], once Eq. 183 is subtracted from the Lagrangian. (In addition, the $\xi$-dependent contributions from the Goldstone, ghosts, and photon all cancel, though they are independent of the background field anyway.) Absorbing $\Delta\mathcal{L}_{(0)}$ into the effective potential, at next-to-leading order in large $\Lambda$ we have
$\begin{split}\Delta V[\phi_{b}]^{(1)}&=\frac{3\Lambda^{2}m_{A}(\varphi_{b})^{2}}{2(4\pi)^{2}}-\frac{3m_{A}(\varphi_{b})^{4}}{8(4\pi)^{2}}\left[1+2\ln\left(1+\frac{\Lambda^{2}}{m_{A}(\varphi_{b})^{2}}\right)\right]\\ &\hphantom{{}={}}+\frac{\Lambda^{2}m_{h}(\varphi_{b})^{2}}{2(4\pi)^{2}}-\frac{m_{h}(\varphi_{b})^{4}}{8(4\pi)^{2}}\left[1+2\ln\left(1+\frac{\Lambda^{2}}{m_{h}(\varphi_{b})^{2}}\right)\right].\end{split}$ (184)
Since we expect $\Lambda$ to be much larger than any mass scale in the problem, the one-loop contributions to the effective potential are subdominant to the classical potential when
$\frac{\Lambda^{2}}{(4\pi f)^{2}}$ $\lesssim\frac{m(\varphi_{b})^{2}}{\max\left[m(\varphi_{b})^{2},m_{A}(\varphi_{b})^{2},m_{h}(\varphi_{b})^{2}\right]}.$ (185)
If coupling functions are chosen such that any of the fields’ masses are not dependent on $\varphi_{b}$, then they do not contribute to the one loop effective potential and are not considered in Eq. 185.
At two-loop order, it is unclear whether quartic divergences cancel in a manner similar to those that arise at one loop [via Eq. 170]. Though addressing this question is important to assess the naturalness of coupled scalar theories in general, the two-loop calculations are substantially more involved and we leave them to future work.
Appendix C Axion to dark photon conversion at large coupling
In this section, we detail the dynamics of dark photon production from axions in the large-$\beta$ regime up until perturbation theory breaks down. We show that the breakdown of perturbation theory coincides with the time when friction sourced by the dark photon becomes important. As a result, the full study of this production mechanism ultimately requires nonlinear simulations.
The equations of motion that we consider are Eq. 21b for the dark photon and
00 0 $=\ddot{\phi}+3H\dot{\phi}+\left(\frac{k^{2}}{a^{2}}+m_{a}^{2}\right)\phi-\frac{\beta}{f_{a}}\left[\mathbf{E}\cdot\mathbf{B}\right](k)$ (186)
for the axion, neglecting possible higher-order self-interactions in the axion potential. The backreaction of the dark photon onto the axion’s equation of motion is given by (see, e.g., Ref. [248])
$\langle\mathbf{E}\cdot\mathbf{B}\rangle$ $=-\frac{1}{2a^{3}}\sum_{\lambda=\pm}\lambda\int\frac{\mathrm{d}^{3}k}{(2\pi)^{3}}\frac{\mathrm{d}{}}{\mathrm{d}{t}}\left|A_{\lambda}\right|^{2}$ (187)
at the background level. We also require
[…]
where the polarization vectors are defined below Eq. 139, $\xi_{1}$ and $\xi_{2}$ denote two different times, and we keep only the $+$ polarization in the four-point function for simplicity. Note the identity
$\bm{\varepsilon}^{\lambda}(\mathbf{p})^{\ast}\cdot\bm{\varepsilon}^{\lambda^{\prime}}(\mathbf{p}+\mathbf{k})$ $=\frac{1}{2}+\frac{\lambda\lambda^{\prime}}{2}\frac{\left|\mathbf{p}\cdot(\mathbf{p}+\mathbf{k})\right|}{\left|\mathbf{p}\right|\left|\mathbf{p}+\mathbf{k}\right|}.$ (189)
The initial condition of the dark photon is set by the Bunch-Davies vacuum:
$A_{\lambda}(\mathbf{k})$ $=\frac{1}{\sqrt{2}\sqrt[4]{k^{2}+a^{2}m_{A}^{2}}}$ (190a)
$\dot{A}_{\lambda}(\mathbf{k})$ $=\mathrm{i}\frac{k}{a}A_{\lambda}(\mathbf{k}).$ (190b)
We decompose the axion into a homogeneous part $\bar{\phi}(t)$ and a perturbation $\delta\phi(t,\mathbf{k})$ which satisfy
$\frac{\beta}{f_{a}}\langle\mathbf{E}\cdot\mathbf{B}\rangle$ $=\ddot{\bar{\phi}}+3H\dot{\bar{\phi}}+m_{a}^{2}\bar{\phi},$ (191)
$\frac{\beta}{f_{a}}\left(\mathbf{E}\cdot\mathbf{B}-\langle\mathbf{E}\cdot\mathbf{B}\rangle\right)$ $=\delta\ddot{\phi}+3H\delta\dot{\phi}+\left(\frac{k^{2}}{a^{2}}+m_{a}^{2}\right)\delta\phi.$ (192)
We take a homogeneous initial condition for the axion,
$\delta\phi=\delta\dot{\phi}=0.$ (193)
Finally, we enumerate a number of useful identities. If
$A(t)=A_{0}\bar{A}(t)\exp\left\{-\frac{1}{2}\left(\frac{k-k_{0}(t)}{a(t)\epsilon(t)}\right)^{2}\right\},$ (194)
then
$A(t)A(t^{\prime})\xrightarrow[\epsilon\to 0]{}A_{0}^{2}\bar{A}^{2}(t)\pi^{3/4}2^{1/4}[a(t)\epsilon(t)]^{3/2}\delta(k-k_{0}(t))\sqrt{\delta(k_{0}(t)-k_{0}(t^{\prime}))}.$ (195)
Note that we’ve represented the temporal delta function inside a square root, anticipating the appearance of another factor of the delta function coming from the other pair of $A$s. One can solve Eq. 192 for the variance of the axion field in terms of the dark photon as
[…]
If we assume that time derivatives can be replaced a factor of some scale $\Gamma$ and that $A_{\mathbf{k}}(\xi)$ is of the form Eq. 195 with $A_{0}=1/\sqrt{k}$, then
$\langle\delta\phi^{2}\rangle$ $=\frac{1}{2^{7/2}\pi^{5/2}}\frac{\beta^{2}m_{A}^{4}}{f_{a}^{2}}\int_{0}^{\tau}\frac{\mathrm{d}\xi}{\xi^{3/2}}{\left|\frac{d\kappa_{0}}{d\xi}\right|}^{-1}\kappa_{0}^{2}\Gamma^{2}\bar{A}^{4}\epsilon^{3}\int_{0}^{2k_{0}}\kappa\mathrm{d}\kappa G_{\kappa}^{2}(\xi,\tau)\left(1-\left(\frac{\kappa}{2\kappa_{0}}\right)^{2}\right)^{2}$ (197)
with $t=\tau/m_{A}$, $a=a_{0}\sqrt{\tau}$ and $\kappa=a_{0}m_{A}\tau$, with all other quantities implicitly rewritten in units of $m_{A}$.
Having enumerated the important formulae, we now discuss the dynamics. The dark photon is not produced so long as $\dot{\bar{\phi}}$ remains below the production threshold velocity $\dot{\bar{\phi}}_{\mathrm{P}}$, during which time the axion evolves as an entirely free field:
$\bar{\phi}=\bar{\phi}(0)\left(1-\frac{1}{5}(m_{a}t)^{2}\right)$ (198)
and $\delta\phi=0$. The threshold velocity $\dot{\bar{\phi}}_{\mathrm{P}}$ is the velocity beyond which at least one dark photon mode becomes tachyonic, i.e.,
$\frac{k_{\mathrm{P}}^{2}}{a^{2}}+\frac{\beta}{f_{a}}\dot{\bar{\phi}}_{\mathrm{P}}\frac{k_{\mathrm{P}}}{a}+m_{A}^{2}=0,$ (199)
which picks out the velocity
$\dot{\bar{\phi}}_{\mathrm{P}}=-\frac{2m_{A}f_{a}}{\beta}$ (200)
and wave number $k_{\mathrm{P}}/a=m_{A}$. Once the velocity of the axion exceeds this threshold one of the dark photon polarizations (in our convention, the $+$ polarization) undergoes a tachyonic resonance. The characteristic dark photon wave number produced is roughly $k/a\sim\beta\dot{\Phi}/f_{a}$, so very shortly after resonance starts we may neglect the dark photon’s mass. In order to simplify our discussion, we introduce the following set of variables
[…]
and the dimensionless initial axion displacement is $\theta_{0}\equiv\bar{\phi}(0)/f_{a}$. In terms of these dimensionless variables, the dark photon equation of motion becomes
00 0 $=\frac{1}{\tau^{1/4}}\frac{\mathrm{d}^{2}{}}{\mathrm{d}{\tau}^{2}}(\tau^{1/4}A)+\left(\frac{\kappa^{2}}{\tau}+1+\frac{3}{16\tau^{2}}-\frac{2}{5}B\kappa\tau^{1/2}\right)A.$ (202)
We may solve this equation in the WKB approximation to find
$A^{+}\approx A^{+}(0)\left[\frac{\tau_{\mathrm{P}}}{\tau}\right]^{1/4}\exp\left[\int_{\tau_{\mathrm{P}}}^{\tau}\mathrm{d}\tau^{\prime}\sqrt{\frac{2}{5}B\kappa\tau^{\prime 1/2}-\frac{\kappa^{2}}{\tau^{\prime}}-1-\frac{3}{16\tau^{\prime 2}}}\right],$ (203)
where $\tau_{\mathrm{P}}$ is the dimensionless time at which the exponent becomes real. For modes that start growing some time after $\tau_{\mathrm{P}}$, it is a decent approximation to drop both the third and fourth terms (bare mass and Hubble respectively). In this approximation, we have
$\tau_{\mathrm{P}}=\left(\frac{5\kappa}{2B}\right)^{2/3},$ (204)
and the integral reduces to
$A^{+}\approx A_{0}\left[\frac{\tau_{\mathrm{P}}}{\tau}\right]^{1/4}\exp\left[\frac{\kappa^{4/3}}{B^{1/3}}\int_{(5/2)^{2/3}}^{\tau(B/\kappa)^{2/3}}\mathrm{d}x\sqrt{\frac{2}{5}x^{1/2}-\frac{1}{x}}\right],$ (205)
which may be expressed in terms of a Gaussian hypergeometric function. However, we simply approximate $A^{+}$ by a Gaussian centered on its local maximum—and excellent approximation because the dark photon transfer function is sharply peaked.
The local maximum of the argument of the exponent in Eq. 205 is located at
$\kappa=\kappa_{0}\equiv C_{0}B\tau^{3/2}$ (206)
with $C_{0}=0.077082$ and is locally approximated by a parabola
$C_{1}B\tau^{2}\left(1-C_{2}\left(\frac{\kappa}{\kappa_{0}}-1\right)^{2}\right)$ (207)
with $C_{1}=0.0788847$ and $C_{2}=0.253765$. With this approximation, it is straightforward to calculate
$\langle\mathbf{E}\cdot\mathbf{B}\rangle$ $=-\frac{m_{A}^{5}a_{0}}{2\tau^{3/2}}\int\frac{\mathrm{d}^{3}\kappa}{(2\pi)^{3}}\kappa\frac{\mathrm{d}{}}{\mathrm{d}{\tau}}\left|A\right|^{2}$ (208)
$\approx-\frac{m_{A}^{4}}{(2\pi)^{2}}\sqrt{\frac{2\pi C_{1}B}{C_{2}}}e^{2C_{1}B\tau^{2}}\left(\frac{5\kappa_{0}}{2B}\right)^{1/3}\frac{\kappa_{0}^{3}}{\tau^{2}},$ (209)
where we have approximated the Gaussian by a $\delta$-function
$e^{-2C_{1}B\tau^{2}C_{2}\left[\frac{\kappa}{\kappa_{0}}-1\right]^{2}}\approx\sqrt{\frac{\pi}{2C_{1}C_{2}B\tau^{2}}}\delta\left(\frac{\kappa}{\kappa_{0}}-1\right).$ (210)
Using this estimate for $\langle\mathbf{E}\cdot\mathbf{B}\rangle$, we can compute the backreaction onto the homogeneous mode $\bar{\phi}$ using the Green’s function for Eq. 191 with appropriate boundary conditions:
$G(t,\xi)=\frac{\pi}{2^{1/2}}\frac{\xi^{5/4}}{t^{1/4}}\left(J_{-\frac{1}{4}}(m_{a}\xi)J_{\frac{1}{4}}(m_{a}t)-J_{\frac{1}{4}}(m_{a}\xi)J_{-\frac{1}{4}}(m_{a}t)\right).$ (211)
We note that for large $\beta$, backreaction kicks in well before $H=m_{a}$, and therefore the axion may be approximated as massless [or equivalently, we may take the small $t,\xi$ limit of Eq. 211], and so we may approximate
$G(t,\xi)\approx 2\xi\left(1-\sqrt{\frac{\xi}{t}}\right).$ (212)
Thus, we find
$\bar{\phi}$ $=\bar{\phi}(0)\left(1-\frac{1}{5}(m_{a}t)^{2}\right)+\frac{\beta}{f_{a}}\int_{0}^{t}\mathrm{d}\xi G(t,\xi)\langle\mathbf{E}\cdot\mathbf{B}\rangle$ (213)
$\approx f_{a}\theta_{0}\left(1-\frac{1}{5}(m_{a}t)^{2}\right)+\frac{\beta}{f_{a}}N_{\mathrm{FR,0}}m_{a}^{2}(\beta\theta_{0})^{3/2}(m_{a}t)e^{2C_{1}\beta\theta_{0}(m_{a}t)^{2}},$ (214)
where the numerical factor is
$N_{\mathrm{FR,0}}=\sqrt{\frac{C_{1}}{(2\pi)^{3}C_{2}}}\frac{5^{1/3}}{2^{1/3}}\frac{C_{0}^{10/3}}{16C_{1}^{2}}=9.4058\times 10^{-5}.$ (215)
The time of backreaction is roughly determined by $\dot{\bar{\phi}}=0$, which occurs at
$t_{\mathrm{BR}}=\frac{1}{m_{a}}\sqrt{\frac{1}{4C_{1}\beta\theta_{0}}W\left[\frac{f_{a}^{4}}{25C_{1}N_{\mathrm{FR,0}}^{2}m_{a}^{4}\beta^{6}\theta_{0}^{2}}\right]},$ (216)
where $W$ is the Lambert $W$ function evaluated on its principal branch. At this time, the energy density stored in the dark photon is approximately $\rho_{A}\sim\rho_{a}/\beta$.
Now that we have determined the evolution of the homogeneous mode, we compute the perturbations to the axion to determine whether they become large prior to $t_{\mathrm{BR}}$. We apply Eq. 196, where $G_{k}$ is the Green’s function of the Klein-Gordon equation with appropriate boundary conditions [Eq. 193],
$G_{k}(t,\xi)$ $=2\xi e^{-\mathrm{i}m_{a}(\xi+t)}\left[{{}_{1}F_{1}}\left(\frac{\mathrm{i}k^{2}t}{2m_{a}a^{2}}+\frac{1}{4};\frac{1}{2};2\mathrm{i}m_{a}\xi\right){{}_{1}F_{1}}\left(\frac{\mathrm{i}k^{2}t}{2m_{a}a^{2}}+\frac{3}{4};\frac{3}{2};2\mathrm{i}m_{a}t\right)\right.$ (217)
$\hskip 56.9055pt\left.-\sqrt{\frac{\xi}{t}}{{}_{1}F_{1}}\left(\frac{\mathrm{i}k^{2}t}{2m_{a}a^{2}}+\frac{3}{4};\frac{3}{2};2\mathrm{i}m_{a}\xi\right){{}_{1}F_{1}}\left(\frac{\mathrm{i}k^{2}t}{2m_{a}a^{2}}+\frac{1}{4};\frac{1}{2};2\mathrm{i}m_{a}t\right)\right].$ (218)
In order to evaluate the integral in Eq. 196, we make use of the $\delta$-function approximation for $A$. In this case, we are dealing with the product of two $A$s evaluated at different momenta. Since each $A$ on its own is essentially a $\delta$-function, the product of four $A$s of different arguments imposes an additional condition: not only must $k$ and $k+p$ match $m_{A}a_{0}\kappa_{0}$, but we must also have $\xi_{1}=\xi_{2}$ because $\kappa_{0}$ is a function of $\xi_{1}$ or $\xi_{2}$. To see this explicitly, we complete the square as in Eq. 195, with
$A=A_{0}\left[\frac{\tau_{\mathrm{P}}}{\tau_{\xi}}\right]^{1/4}e^{C_{1}B\tau_{\xi}^{2}}\exp\left\{-\frac{1}{2}\left(\frac{\kappa-\kappa_{0}}{\kappa_{0}/\sqrt{2C_{1}C_{2}B\tau_{\xi}^{2}}}\right)^{2}\right\}$ (219)
where $\tau_{\xi}$ is $\tau$ in Eq. 201 evaluated at $t=\xi$. Upon multiplying $A(\tau_{\xi_{1}})A(\tau_{\xi_{2}})$, we find a pair of Gaussians, one of which imposes $\kappa=\kappa_{0}$, and the other which imposes $\tau_{\xi_{1}}=\tau_{\xi_{2}}$. This is the simplification that allows us to reduce Eq. 196 to Eq. 197.
To further simplify the integral, we can approximate the Green’s function Eq. 217 in precisely the same way as we approximated it for the homogeneous mode i.e. by dropping the axion’s bare mass, leading to
$G_{k}(t,\xi)\approx\frac{\tau_{\xi}}{m_{A}\kappa\sqrt{\tau}}\sin\left[2\kappa\left(\sqrt{\tau}-\sqrt{\tau_{\xi}}\right)\right].$ (220)
Further, it turns out that the dominant contribution to the integral comes from $\kappa$ and $\tau_{\xi}$ such that the argument of $\sin$ is small, so Eq. 212 is an excellent approximation to Eq. 220 when evaluating Eq. 197. Finally, we replace time derivatives of $A$ with $\Gamma=2C_{1}B\tau$, which is valid once modes start growing rapidly, i.e. $C_{1}B\tau^{2}\gg 1$. With these approximations, evaluating $\langle\delta\phi^{2}\rangle$ becomes trivial. We apply Eq. 197 with
$\begin{split}\kappa_{0}=C_{0}B\tau^{3/2},\hskip 14.22636pt\epsilon=\frac{\kappa_{0}}{\sqrt{2C_{1}C_{2}B\tau^{3}}},\hskip 14.22636pt\Gamma=2C_{1}B\tau,\hskip 14.22636pt{\left|\frac{d\kappa_{0}}{d\tau}\right|}^{-1}=\frac{2}{3C_{0}B\tau^{1/2}},\\ G_{\kappa}=2\xi\left(1-\sqrt{\frac{\xi}{\tau}}\right),\hskip 14.22636pt\bar{A}^{2}=\left(\frac{\tau_{\mathrm{P}}}{\tau}\right)^{1/2}e^{2C_{1}B\tau^{2}},\hskip 14.22636pt\text{and }\tau_{\mathrm{P}}=\left(\frac{5\kappa_{0}}{2B}\right)^{2/3},\end{split}$ (221)
yielding
$\langle\delta\phi^{2}\rangle$ $\approx N_{\mathrm{FR,1}}\frac{m_{a}^{4}\beta^{2}}{f_{a}^{2}}(\beta\theta_{0})^{7/2}(m_{a}t)^{3}e^{4C_{1}\beta\theta_{0}(m_{a}t)^{2}},$ (222)
where
$N_{\mathrm{FR,1}}=\frac{5^{2/3}C_{0}^{20/3}}{2^{2/3}\pi^{5/2}4608C_{1}^{5/2}C_{2}^{3/2}}=3.88565\times 10^{-9}.$ (223)
Having computed the perturbations, we may now compare them to the backreaction onto the homogeneous mode
$\frac{\langle\delta\phi^{2}\rangle}{\left[\bar{\phi}(t)-\bar{\phi}(0)\left(1-\frac{1}{5}(m_{a}t)^{2}\right)\right]^{2}}\approx 0.43921\sqrt{\beta\theta_{0}}m_{a}t.$ (224)
Evaluating this ratio at $t=t_{\mathrm{BR}}$, we find that the perturbations become order 1 at roughly the same time backreaction becomes important. This coincidence could have been anticipated, given that it is a common process $\mathbf{E}\cdot\mathbf{B}$ that drives both perturbations and backreaction.
Appendix D Delayed axion oscillations
D.1 Efficient growth condition
As discussed in IV.1, dark photon production from axions can be efficient with $\beta\theta_{0}\ll 1$ if oscillations are substantially delayed relative to the standard time, $H\sim m_{a}$. In this appendix, we derive the modified condition for efficient resonance, Eq. 48.
The growth rate of the transverse modes may be calculated from the equation of motion Eq. 21b by mapping it onto an equivalent Mathieu-type equation,
00 0 $=A^{\prime\prime}(\tau)+(a+2q\cos 2\tau)A(\tau),$ (225)
with $2\tau=m_{a}t$, primes representing $\tau$ derivatives, and
$\tilde{a}$ $=4\left(\frac{k^{2}}{a^{2}m_{a}^{2}}+\frac{m_{A}^{2}}{m_{a}^{2}}\right)$ (226)
$\tilde{q}$ $=\frac{2k\beta\theta_{0}}{am_{a}}\left(2H_{\mathrm{osc}}t\right)^{-3/4}.$ (227)
Eq. 225 is not technically a Mathieu-type equation since the coefficients $\tilde{a}$ and $\tilde{q}$ depend on $t$. However, so long as $m_{a}t\gg 1$, this time dependence is slow and can be treated adiabatically. The instantaneous growth rate for the transverse dark photon modes is therefore
$\Gamma$ =±$=\pm\frac{m_{a}}{4}\sqrt{\tilde{q}^{2}-(1-\tilde{a})^{2}}.$ (228)
The total growth of a given mode is then
$A(t)/A(0)\approx\exp\left\{\int_{1/2H_{\mathrm{osc}}}^{t}\mathrm{d}t^{\prime}\,\Re\left[\Gamma\right]\right\}.$ (229)
Efficient production requires growth rates exceeding $H$. Near threshold, the resonance takes place almost entirely at a single wave number; as long as this single wave number remains inside the resonance band (i.e., has not redshifted out of it), the energy density growth is approximated by202020 Once this mode has redshifted outside of the resonance band, the energy density growth is linear instead of exponential.
$\frac{\dot{\rho}_{A}}{\rho_{A}}$ $\approx\max_{k}2\Gamma(k)=\frac{1}{2}\beta\theta_{0}m_{a}\sqrt{\left(1-\frac{4m_{A}^{2}}{m_{a}^{2}}\right)\left(\frac{H}{H_{\mathrm{osc}}}\right)^{3/2}+\frac{1}{4}\beta^{2}\theta_{0}^{2}\left(\frac{H}{H_{\mathrm{osc}}}\right)^{3}}.$ (230)
There are two limits of this equation to consider. First, if $m_{A}\leq 2m_{A}$, then the first term inside the root is negative, and must be compensated for by the second term, requiring $\beta\theta_{0}\gg 1$. In this case, resonance starts almost immediately and quickly becomes nonperturbative (corresponding to the case discussed in Appendix C). On the other hand, if $\beta\theta_{0}$ is small, then resonance requires $m_{a}\geq 2m_{A}$. In this limit, modes redshift from one edge of the resonance band to the other in a time interval
$\Delta t=\frac{\beta\theta_{0}}{H_{\mathrm{osc}}},$ (231)
where we’ve taken the $2m_{A}\ll m_{a}$, $t\sim 1/2H_{\mathrm{osc}}$, and $\beta\theta_{0}\ll 1$ limits. The resulting total growth is
$\rho_{A}\left([2H_{\mathrm{osc}}]^{-1}+\Delta t\right)\sim\rho_{A}\left([2H_{\mathrm{osc}}]^{-1}\right)\exp\left\{\frac{m_{a}}{2H_{\mathrm{osc}}}\beta^{2}\theta_{0}^{2}\right\}.$ (232)
Thus, we arrive at the efficient resonance condition Eq. 48,
$\beta\theta_{0}\sqrt{\frac{m_{a}}{2H_{\mathrm{osc}}}}\gg 1.$ (233)
D.2 Axion fragmentation time
As discussed in Sec. IV.1, attractive self-interactions can fragment the axion before it transfers all of its energy to the dark photon. For a general axion potential approximated as
$V(\phi_{a})=m_{a}^{2}f_{a}^{2}\left[\frac{1}{2}\left(\frac{\phi_{a}}{f_{a}}\right)^{2}+\frac{1}{3!}A\left(\frac{\phi_{a}}{f_{a}}\right)^{3}+\frac{1}{4!}B\left(\frac{\phi_{a}}{f_{a}}\right)^{4}\right],$ (234)
the growth rate of axion perturbations (derived in Appendix C of Ref. [169]) is
$\Gamma=\left|\delta\omega\right|\Re\sqrt{1-\left(1+\frac{k^{2}}{2m_{a}\delta\omega a^{2}}\right)^{2}},$ (235)
where $\delta\omega$ is the change in the oscillation frequency of the axion relative to its bare mass. At leading order in self-interactions,
$\delta\omega=m_{a}\frac{3B-5A^{2}}{24}\langle\Theta^{2}\rangle.$ (236)
We compare this to the expression Eq. 230 in the limit $m_{A}\to 0$, $t\sim 1/2H_{\mathrm{osc}}$, and $\beta\theta_{0}\ll 1$: the dark photon mode growth rate is $\Gamma=m_{a}\beta\theta_{0}/4$. Thus, the axion fragments before the dark photon absorbs $\mathcal{O}(1)$ of the axion energy density unless
$\left|m_{a}\beta\theta_{0}/4\right|>-\delta\omega.$ (237)
Implicit in this inequality is the statement that if axion self-interactions are repulsive then the axion perturbations do not grow. To achieve efficient resonance at the natural expectation $\beta\theta_{0}\sim\alpha_{D}$ without backreaction, attractive self-interactions are practically excluded.
In summary, though delayed axion oscillations can significantly reduce the requisite dimensionless axion–dark-photon coupling, even very weak attractive self-interactions cause the axion to fragment before all the dark photon dark matter is produced. Viable dark photon production essentially requires repulsive self-interactions. As the axion is a periodic field, oscillations of large enough amplitude are guaranteed to induce attractive self-interactions. Mechanizing repulsive self-interactions require quite a bit of machinery (see the extensive discussion in Ref. [171]), and even then only extends over a limited field range.
“Axion friendship” [169, 170] is a phenomenon whereby pairs of axions with nearby masses exchange energy through nonlinear autoresonance [249, 250, 251, 252, 253]. If the axions in the friendly pair have hierarchical decay constants $f_{j+1}/f_{j}=\mathcal{F}\gg 1$, the axion with the smaller decay constant can siphon nearly all the energy from its friend, parametrically delaying the onset of decaying oscillations to $H_{\mathrm{osc}}\sim m_{a}\mathcal{F}^{4/3}$. Although the requirement of nearby masses constitutes a coincidence, such pairs may be common in axiverse-type scenarios [43]. In fact, it has been found that in realistic string compactifications the friendship parameter $\mu\equiv m_{j}/m_{j+1}$, which measures the ratio of neighboring axion masses, lies in the requisite range $\mu\in[0.75,1]$ for a growing fraction of pairs as the number of axions increases [172]. [^21] The range $\mu\in[0.75,1]$ is for a cosine potential which has attractive self-interactions. In the case of repulsive self-interactions the range switches to $\mu\in[1,1.25]$ or so.
Appendix E Plasma dynamics
Between electron-positron annihilation and recombination, the SM plasma comprises photons, nonrelativistic electrons, and nonrelativistic nuclei. The Universe is charge neutral, so the electron and proton number densities are both
$n_{e}=\eta_{B}\frac{2\zeta(3)}{\pi^{2}}T^{3},$ (238)
where $\eta_{B}\approx 6\times 10^{-10}$ is the baryon-to-photon ratio and $T$ the plasma temperature. The rate of Thomson scattering between photons and electrons and of Coulomb scattering between electrons and protons are respectively $\dot{\kappa}_{T}=n_{e}\sigma_{T}$ and $\dot{\kappa}_{C}=n_{e}\sigma_{C}$, where the corresponding cross sections are [61]
$\sigma_{T}$ $=\frac{8\pi\alpha^{2}}{3m_{e}^{2}}$ (239)
$\sigma_{C}$ $=\frac{4\sqrt{2\pi}\alpha^{2}}{3\sqrt{m_{e}T^{3}}}\ln\left(\frac{4\pi T^{3}}{\alpha^{3}n_{e}}\right),$ (240)
In terms of the proton-to-electron mass ratio $\mu_{p}\equiv m_{p}/m_{e}$, the Thomson cross section for protons is $\mu_{p}^{-2}\sigma_{T}$. In the epoch of interest, the Thomson and Coulomb rates are quite rapid relative to the Hubble rate:
$\frac{\dot{\kappa}_{T}}{H}$ $\approx 10^{3}\frac{T}{{\mathrm{eV}}}\frac{\eta_{B}}{6\times 10^{-10}}$ (241a)
$\frac{\dot{\kappa}_{C}}{H}$ $\approx 5.8\times 10^{12}\left(\frac{T}{{\mathrm{eV}}}\right)^{1/2}\frac{\eta_{B}}{6\times 10^{-10}}$ (241b)
taking $H^{2}=g_{\star}\pi^{2}T^{4}/90M_{\mathrm{pl}}^{2}$ where $g_{\star}\approx 3.38$. Treating the plasma as three coupled, perfect fluids that jointly satisfy the energy-momentum conservation equations is therefore an excellent approximation. For simplicity of exposition, we ignore the fact that about a quarter of the protons are bound in helium nuclei, considering a single fluid of protons. Strictly speaking, this description is only applicable after electron-positron annihilation completes, which occurs around a temperature of $T\approx 20\leavevmode\nobreak\ \mathrm{keV}$ [254]; at higher temperatures the positrons and electrons are equally abundant, and $e^{+}e^{-}$ scattering processes are also important.
Large-scale photons (i.e., with wave number much smaller than the plasma temperature) couple to the bulk motion of the fluid via Lorentz forces. Longitudinal modes (or scalar degrees of freedom in cosmological perturbation theory) couple to density perturbations and gravitational potentials, making for a large, complex set of coupled differential equations. Transverse photon modes, on the other hand, induce vorticity in the plasma. Vector perturbations of the metric, however, do not appear explicitly in the momentum conservation equations, and in fact do not couple directly to a fluid’s velocity if that fluid supports no anisotropic stress [255]. As already argued, the photons, electrons, and protons interact rapidly enough to suppress any anisotropic stress in the epoch of interest. The equations of motion for the transverse (vortical) modes [decomposed onto the same basis as Eq. 139c] in the plasma are
$\dot{u}_{\gamma,\pm}$ $=\dot{\kappa}_{T}\left(u_{e,\pm}-u_{\gamma,\pm}\right)+\mu_{p}^{-2}\dot{\kappa}_{T}\left(u_{p,\pm}-u_{\gamma,\pm}\right)$ (242a)
$\dot{u}_{e,\pm}$ $=-Hu_{e,\pm}+R_{e}\dot{\kappa}_{T}\left(u_{\gamma,\pm}-u_{e,\pm}\right)+\dot{\kappa}_{C}\left(u_{p,\pm}-u_{e,\pm}\right)-\frac{e}{m_{e}}\dot{A}_{\pm}$ (242b)
$\dot{u}_{p,\pm}$ $=-Hu_{p,\pm}+R_{e}\mu_{p}^{-3}\dot{\kappa}_{T}\left(u_{\gamma,\pm}-u_{p,\pm}\right)+\mu_{p}^{-1}\dot{\kappa}_{C}\left(u_{e,\pm}-u_{p,\pm}\right)+\frac{e}{m_{p}}\dot{A}_{\pm}.$ (242c)
Here $u_{a,\pm}$ are the transverse polarizations of the fluid velocity for the photons, electrons, and protons ($a=\gamma$, $e$, and $p$, respectively). The photon-to-electron density ratio
$R_{e}$ $\equiv\frac{4\bar{\rho}_{\gamma}}{3\bar{\rho}_{e}}=1.175\times 10^{4}\,\frac{T}{{\mathrm{eV}}}\left(\frac{\eta_{B}}{6\times 10^{-10}}\right)^{-1}$ (243)
encodes the relative rate of momentum transfer for electrons and photons due to Thomson scattering.
The photon couples to the SM 3-current, which is given terms of the proton-electron slip by
$J_{\pm}$ $=n_{e}\left(u_{p,\pm}-u_{e,\pm}\right)$ (244)
and, combining Eqs. 242b and 242c, evolves according to
$\begin{split}\dot{J}_{\pm}+\left[4H+\left(1+\mu_{p}^{-1}\right)\dot{\kappa}_{C}\right]J_{\pm}&=-R_{e}\dot{\kappa}_{T}n_{e}\left[u_{\gamma,\pm}-u_{e,\pm}+\mu_{p}^{-3}\left(u_{\gamma,\pm}-u_{p,\pm}\right)\right]\\ &\hphantom{{}={}}+n_{e}\frac{e}{m_{e}}\left(1+\mu_{p}^{-1}\right)\dot{A}_{\pm}.\end{split}$ (245)
From here on out we drop the term arising from Thomson scattering of protons, $\mu_{p}^{-3}\left(u_{\gamma,\pm}-u_{p,\pm}\right)$, and also neglect Hubble friction when it is directly added with $\dot{\kappa}_{T}$ or $\dot{\kappa}_{C}$. Taking a time derivative of Eq. 245 and simplifying with Eq. 242b, the current equation of motion may be written in the form
$\ddot{J}_{\pm}+\nu(t)\dot{J}_{\pm}+\omega(t)^{2}J_{\pm}$ $=-HR_{e}\dot{\kappa}_{T}n_{e}u_{e,\pm}+n_{e}\frac{e}{m_{e}}\ddot{A}_{\pm}+\gamma(t)n_{e}\frac{e}{m_{e}}\dot{A}_{\pm}$ (246)
where
$\nu(t)$ $=\dot{\kappa}_{C}+\left(1+R_{e}\right)\dot{\kappa}_{T}$ (247a)
$\omega(t)^{2}$ $=\left(1+\mu_{p}^{-1}R_{e}\right)\dot{\kappa}_{C}\dot{\kappa}_{T}$ (247b)
$\gamma(t)$ $=\left(1+\mu_{p}^{-1}R_{e}\right)\dot{\kappa}_{T}.$ (247c)
In Eqs. 246 and 247 we took $1+\mu_{p}^{-1}\approx 1$, but $\mu_{p}^{-1}R_{e}$ is not necessarily small. However, consultation with Eqs. 241 and 243 shows that $\omega(t)$ and $\gamma(t)$ are always negligible relative to $\nu(t)$. On the same basis, the term proportional to the electron velocity in Eq. 246 is also negligible unless the electron velocity is drastically larger than $u_{p,\pm}-u_{e,\pm}$. With these approximations, Eq. 246 depends only $\dot{J}_{\pm}$, $\ddot{J}_{\pm}$, and $\ddot{A}_{\pm}$. Since all time-dependent coefficients vary only on the Hubble time scale, which we have already taken to be negligible, we may approximate
$\dot{J}_{\pm}+\nu(t)J_{\pm}$ $=\frac{\omega_{p}(t)^{2}}{e}\dot{A}_{\pm}$ (248)
with the plasma damping rate $\nu(t)$ defined in Eq. 247a and the plasma frequency defined by
$\omega_{p}^{2}$ $=\frac{e^{2}n_{e}}{m_{e}}=\frac{2\zeta(3)e^{2}\eta_{B}}{\pi^{2}}\frac{T^{3}}{m_{e}}=\left[5.121\times 10^{-9}\leavevmode\nobreak\ {\mathrm{eV}}\sqrt{\frac{\eta_{B}}{6\times 10^{-10}}}\left(\frac{T}{{\mathrm{eV}}}\right)^{3/2}\right]^{2}.$ (249)
Relative to Ref. [61], which considered the coupling of kinetically mixed dark photons to the electron-ion plasma in the interstellar medium, the Thomson scattering of photons with electrons relevant at early times provides an additional resistance of the plasma to bulk charge separation. In fact, momentum transfer due to Thomson scattering becomes more important than that due to Coulomb scattering at temperatures above $T\sim 190\leavevmode\nobreak\ {\mathrm{eV}}$, coincident with the time period of interest (redshifts $z\sim 10^{6}$).
Finally, the Thomson scattering and momentum exchange rates and the Coulomb rate relative to the plasma frequency are
$\frac{R_{e}\dot{\kappa}_{T}}{\omega_{p}}$ $=\frac{8\alpha^{2}\sqrt{2\zeta(3)\eta_{B}}}{3e}\frac{2\pi^{4}}{45\zeta(3)\eta_{B}}\left(\frac{T}{m_{e}}\right)^{5/2}\approx 106.9\left(\frac{T}{m_{e}}\right)^{5/2}$ (250a)
$\frac{\dot{\kappa}_{C}}{\omega_{p}}$ $=\frac{4\sqrt{2\pi}\alpha^{2}\sqrt{m_{e}n_{e}}}{3e\sqrt{m_{e}T_{e}^{3}}}\ln\Lambda_{C}=\frac{8\alpha^{2}\sqrt{\zeta(3)\eta_{B}/\pi}}{3e}\ln\Lambda_{C}\approx 2.84\times 10^{-7}.$ (250b)
At temperatures $T\lesssim 10^{-2}m_{e}$ (but still before recombination), the resistivity of the plasma is therefore negligible compared to the plasma frequency.